### v2

# Jiehan Liu

Abstract
This Element offers a practical guide to estimating conditional marginal effects—how treatment effects vary with a moderating variable—using modern statistical methods. Commonly used approaches, such as linear interaction models, often suffer from unclarified estimands, limited overlap, and restrictive functional forms. This guide begins by clearly defining the estimand and presenting the main identification results. It then reviews and improves upon existing solutions, such as the semiparametric kernel estimator, and introduces robust estimation strategies, including augmented inverse propensity score weighting with Lasso selection (AIPW-Lasso) and double machine learning (DML) with modern algorithms. Each method is evaluated through simulations and empirical examples, with practical recommendations tailored to sample size and research context. All tools are implemented in the accompanying interflex package for R.
Keywords: causal inference, conditional marginal effects, double machine learning
Table 1: List of Acronyms and Their First Appearance
Acronym Meaning First Appearance
AIPW Augmented Inverse Propensity Weighting Ch1
ATE Average Treatment Effect Ch1
CATE Conditional Average Treatment Effect Ch1
CME Conditional Marginal Effect Ch1
DML Double Machine Learning Ch1
DGP Data Generating Process Ch1
SUTVA Stable Unit Treatment Value Assumption Ch1
IPW Inverse Propensity Weighting Ch1
LSCV Least-Squares Cross-Validation Ch2
MSPE Mean-Squared Prediction Error Ch2
OLS Ordinary Least Squares Ch2
SVCM Smooth Varying-Coefficient Model Ch2
PLRM Partially Linear Regression Model Ch3
PO-Lasso Partialing-Out Lasso Ch3
IRM Interactive Regression Model Ch4
Chapter 1. Introduction
Understanding conditional relationships is central to social science research. The impact of a treatment—such as an experimental intervention, a policy, or an institution—on social, political, and economic outcomes often varies systematically across subgroups or contexts. Researchers are particularly interested in how the effect of a treatment $D$ on an outcome $Y$ changes with the value of a pre-determined covariate $X$, known as the moderator, which is unaffected by $D$.
Methods for analyzing conditional relationships have evolved over time. Traditionally, researchers have relied on linear interaction models to probe these relationships. BCG2006 (hereafter, BCG 2006) introduced best practices for estimating and interpreting such models, including the widely used “marginal effect plot” to visualize these relationships when the moderator $X$ is continuous. Figure 1, reproduced from huddy2015expressive, provides an example. The x-axis represents the moderator $X$ (partisan identity), while the y-axis represents the effect of the treatment $D$ (experimental partisan threat) on the outcome $Y$ (anger). The gray ribbon indicates the 95% (pointwise) confidence intervals.
Figure 1: Adapted from huddy2015expressive, Figure 2A
Researchers typically draw several inferences from figures like this. First, they assess whether the effect of $D$ on $Y$ differs from zero by checking whether the horizontal zero line intersects the gray ribbon at a specific value of the moderator. Second, they identify regions where the gray ribbon lies entirely above or below zero (e.g., when partisan identity is larger than $0.2$), interpreting these as conditions under which the theoretical effect is supported. Third, they examine whether the effect of $D$ on $Y$ changes monotonically with $X$, inferring whether the moderator amplifies or weakens the treatment effect, either causally or descriptively.
More recently, hainmueller2019much (hereafter, HMX 2019) highlighted the often-violated assumptions underlying these models, including the assumption of overlap and the reliance on linear interaction effects. To address these limitations, they introduced semiparametric kernel estimators, which relax functional form assumptions. Despite these advances, existing methods have several limitations. First, they are primarily motivated by an outcome-modeling perspective and often lack clear connections between estimation strategies and the underlying estimands of interest, leading to potential misinterpretation of estimated coefficients. Second, these approaches frequently rely on strong parametric assumptions. While the kernel estimator relaxes many of the functional form constraints of the linear estimator, it still imposes some structural constraints that may be violated in practice—for example, assuming that the higher order terms of the covariates, such as interactions, do not enter the outcome equation. Moreover, the overlap assumption is typically not critically evaluated. Third, existing methods have not systematically addressed important applications, such as those involving discrete outcomes.
This Element addresses these limitations through three key contributions. First, we provide precise definitions of the estimands of interest within the modern causal inference framework, clarifying the targets of identification and aligning estimation strategies—including parametric, semiparametric, and doubly robust machine-learning methods—with theoretical objectives. Second, we leverage recent advances in double/debiased machine learning (DML) estimators to present a flexible framework that accommodates a wide range of data-generating processes (DGPs) and use cases, including discrete and continuous treatments or outcomes, as well as potentially high-dimensional covariates, while ensuring valid inference for the targeted parameters. Finally, we offer practical recommendations for researchers based on evidence from simulations and empirical examples.
1.1 Define the Problem
We begin by defining our problem using the notations of the Neyman-Rubin potential outcomes framework (neyman1923application; rubin1974estimating), which provides the foundation for the rest of the Element. Consider a study with $n$ units, indexed by $i=1,2,\dots,n$. These units are independently and identically distributed (i.i.d.) from an infinite population denoted by $\mathcal{P}$. To simplify the discussion of uncertainty measures, we assume the existence of a super-population. This assumption is also standard in the DML literature.
For each unit $i$, we observe the quadruple $(Y_{i},D_{i},X_{i},Z_{i})$, where $Y_{i}$ denotes the outcome of interest, $D_{i}$ is the treatment indicator, and $V_{i}=(X_{i},Z_{i})$ represents a set of covariates not causally affected by the treatment and comprises two components: the moderator of interest $X_{i}\in\mathcal{X}$, which can be either discrete or continuous, and the remaining covariates used for identification $Z_{i}\in\mathcal{Z}$, which can be potentially high-dimensional. Within the scope of this Element, researchers are committed to studying the effect of $D$ on $Y$ conditional on $X$ (and only $X$), meaning that they do not select variables for analysis after the data have been collected. We formalize the data structure and the sampling procedure as follows:
Assumption 1
Random Sampling.
$\{(Y_{i},D_{i},V_{i}),i=1,2,\cdots,n\}$ are i.i.d. sampled from the population $\mathcal{P}$.
Next, let’s invoke the stable unit treatment value assumption (SUTVA) to define the potential outcomes for unit $i$.
Assumption 2
SUTVA.
$Y_{i}(d_{1},d_{2},\cdots,d_{n})=Y_{i}(d_{i})$
SUTVA states that unit $i$’s potential outcomes depend only on the treatment that unit receives, not on what others receive, and that each treatment has a single version (rubin1986comment; imbens2015causal).
As a starting point, let us consider the case in which the treatment is binary, where $D_{i}=1$ indicates that unit $i$ receives the treatment condition, and $D_{i}=0$ indicates the control condition. Under SUTVA and a binary treatment, unit $i$ has exactly two potential outcomes, denoted by $\{Y_{i}(d_{i}):d_{i}=0,1\}$. Therefore, we can link the observed outcome to the potential outcomes as follows:
$Y_{i}=D_{i}Y_{i}(1)+(1-D_{i})Y_{i}(0).$ (1)
Equation (1) is sometimes referred to as “consistency” in the causal inference literature; we avoid using this term to prevent confusion with the statistical concept of an estimator converging to the true value of a target parameter.
Conditional Marginal Effect
With a binary treatment, unit $i$’s individual treatment effect is defined as $\tau_{i}=Y_{i}(1)-Y_{i}(0)$, and the average treatment effect (ATE) of the population as $\tau_{ATE}=\mathbb{E}[Y_{i}(1)-Y_{i}(0)]$, where the expectation is taken over the population $\mathcal{P}$. Building on the ATE, the conditional average treatment effect (CATE) captures the heterogeneous treatment effect given the value of covariates $V_{i}$.
$\tau(v)=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid V_{i}=v],$ (2)
CATE measures the average treatment effect for specific groups defined by the value of $V_{i}$. In practice, estimating the CATE helps researchers better understand how the treatment effect varies across different population segments.
The main purpose of this Element is to examine how the treatment effect of $D$ on $Y$ changes given the value of $X$, a single moderator, and remaining covariates $Z$ used for identification. Therefore, we introduce the conditional marginal effect (CME) as a special case of the CATE:
Definition 1 (CME w/ Binary Treatments)
$\theta(x)=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid X_{i}=x].$
Note that $\theta(x)$ is essentially an aggregated version of $\tau(v)$, where we average out the effects of $Z_{i}$ while keeping $X_{i}=x$. Formally, this means:
$\theta(x)=\mathbb{E}[\tau(x,Z_{i})\mid X_{i}=x]=\int\tau(x,z)dF_{Z\mid X=x}(z).$
in which the expectation integrates out the influence of $Z_{i}$, focusing solely on the variation introduced by $X_{i}$, hence the term “marginal.”
Sometimes researchers are interested in comparing $\theta(x)$ at two (or more) different values of $x$. This estimand is referred to as effect modification in the statistics literature (tyler2009; bansak2020estimating):
$\Delta\theta(x_{1},x_{2})=\theta(x_{1})-\theta(x_{2}).$
Importantly, effect modification is an associative estimand that describes treatment heterogeneity and should not be conflated with causal moderation, which refers to the causal effect of $X$ on the effect of $D$ on $Y$ (such as in a factorial experiment). In this Element, we do not discuss causal moderation because the moderator $X$ is usually not quasi-random or even manipulable.
A Toy Example
To clarify the definition of $\theta(x)$, we provide a toy example with eight units, each assigned a survey weight, shown in Table 2. Assume that, after weighting, these units represent the population of interest. In this example, $(X_{i},Z_{i})$ comprises two binary covariates, $D_{i}$ is a binary treatment indicator, and $Y_{i}(1)$ and $Y_{i}(0)$ denote the potential outcomes under treatment and control, respectively. The observed outcome $Y_{i}$ corresponds to the treatment actually received by each unit. $\tau_{i}$ represents the individual treatment effect for unit $i$. We arrange the rows of the table according to the values of $V_{i}$. Our objective is to define the CATE and CME.
Table 2: Toy Example
$i$ $X_{i}$ $Z_{i}$ $Y_{i}(0)$ $Y_{i}(1)$ $\tau_{i}$ $D_{i}$ $Y_{i}$ weight
1 0 0 2 3 1 1 3 1/2
2 0 0 2 0 -2 0 2 1/2
3 0 1 3 7 4 1 7 1/2
4 0 1 5 3 -2 0 5 1/2
5 1 0 10 8 -2 1 8 2/3
6 1 0 4 1 -3 0 4 2/3
7 1 1 9 9 0 1 9 1/3
8 2 1 1 0 -1 0 1 1
Note: Numbers in shaded cells represent counterfactuals and are not observed in a real dataset.
Example 1
(CATE) Calculate $\tau(x,z)$ for $v=(0,0)$:
$\tau(v)=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid V_{i}=(0,0)]=\frac{1+(-2)}{2}=-0.5$
Example 2
(CME) Calculate $\theta(x)$ for $x=0$:
$v=(0,0):\quad\tau(v)=-0.5$ $\Pr(Z_{i}=0\mid X_{i}=0)=\frac{2}{4}=0.5$
$v=(0,1):\quad\tau(v)=1$ $\Pr(Z_{i}=1\mid X_{i}=0)=\frac{2}{4}=0.5$
$\theta(0)$ $=\tau(0,0)\cdot\Pr(Z_{i}=0\mid X_{i}=0)+\tau(0,1)\cdot\Pr(Z_{i}=1\mid X_{i}=0)$
$=(-0.5)\times 0.5+(1)\times 0.5=0.25$
Example 3
(Effect Modification) Calculate $\Delta\theta(x_{1},x_{2})$ for $x_{1}=0$, $x_{2}=2$:
$\theta(0)$ $=0.25$
$\theta(2)$ $=\tau((2,1))=-1$
$\Delta\theta(0,2)$ $=0.25-(-1)=1.25$
The fundamental problem of causal inference states that for each unit, only one potential outcome can be observed at a time (holland1986statistics). For example, those numbers in shaded cells will never be observed. Therefore, given a sample, we cannot directly calculate the CATE or CME as demonstrated above. In the remainder of this Element, we will explore various estimation strategies that allow us to approximate these estimands under conditions of large sample sizes and when the necessary identifying assumptions are satisfied.
1.2 Identification
In this section, we focus on the identification and estimation of CME in the simplest scenario, where both the treatment $D$ and the moderator $X$ are discrete. While this setting is less common in the empirical literature, it provides a clear foundation for understanding the key concepts and results that will be applied throughout the Element.
We begin by presenting the identification results under two key assumptions: unconfoundedness and strict overlap. The goal of identification is to connect a statistical estimand, which can be estimated using observable data, to a meaningful causal estimand, such as the CME, which involves counterfactuals that are inherently unobservable. This connection is established through identifying assumptions. To simplify the discussion further, we assume that the treatment is binary.
Assumption 3 (Unconfoundedness)
$\{Y_{i}(0),Y_{i}(1)\}\perp\!\!\!\perp D_{i}\mid V_{i}=v,\text{ for all }v\in\mathcal{V}.$
The unconfoundedness assumption—also known as ignorability or selection on observables—is a key identifying assumption, though it is typically untestable. In randomized controlled trials, where the treatment assignment $D_{i}$ is randomized, this assumption is satisfied because the potential outcomes, $\{Y_{i}(0),Y_{i}(1)\}$, are independent of the randomly assigned $D_{i}$. Furthermore, in an randomized controlled trial, the probability of receiving the treatment for each participant is known, ensuring that this assumption holds by design.
In observational studies, however, unconfoundedness assumes that treatment assignment $D_{i}$ can be considered as good as randomized only after conditioning on a set of covariates $V_{i}=(X_{i},Z_{i})$, which capture all relevant confounders. This is equivalent to stating that within each cell defined by the vector $v$, the treatment is as-if randomly assigned. However, similar to a stratified randomized experiment, the probability of receiving the treatment may vary across cells. The key difference is that, in observational studies, researchers do not know what that probability is.
Assumption 4 (Strict overlap)
 $\text{ for all }v\in\mathcal{V},\text{ }\exists\eta>0,\ \eta\leq\Pr(D_{i}=1\mid V_{i}=v)\leq 1-\eta,\ \text{with prob. }1.$
In addition to unconfoundedness, ensuring the identifiability of the estimand across all values of the covariates requires the overlap assumption, also known as positivity or common support assumption. This assumption states that every unit must have a non-zero probability of being assigned to each treatment condition, i.e., $0<\Pr(D_{i}=1\mid V_{i}=v)<1$. Intuitively, this means that within each cell defined by $v$, there are both treated and control units, provided the sample size is sufficiently large.
In Assumption 4, we require a slightly stronger condition than standard overlap, strict overlap. It says that the probabilities of treatment assignment given covariates, or propensity scores, must be uniformly bounded away from 0 and 1 by some positive constant $\eta$. Mechanically, this requirement avoids extreme values of propensity scores, thereby ensuring the identifiability of certain estimators, such as inverse propensity score weighting (IPW) and augmented inverse propensity score weighting (AIPW). It is also crucial for theoretical guarantees, including asymptotic normality of these estimators.
Now, we present two approaches for identifying the CME under unconfoundedness and strict overlap: (a) stratified difference-in-means and (b) IPW. When all covariates $V_{i}$ are discrete, these two methods are numerically equivalent (imbens2015causal)—see also the proof for the CME case in the Appendix. However, discussing both methods in this simple discrete setting is important, as they provide the foundation and intuition for more general cases.
Stratified difference-in-means.
Denote $\mu^{1}(v)=\mathbb{E}[Y_{i}\mid D_{i}=1,V_{i}=v]$ and $\mu^{0}(v)=\mathbb{E}[Y_{i}\mid D_{i}=0,V_{i}=v]$, which represent the conditional mean outcomes under treatment and control conditions, respectively, given covariate values $V_{i}=(X_{i},Z_{i})=(x,z)=v$. The limit of the stratified difference-in-means estimator can be written as:
$\Sigma_{z}\{(\mu^{1}(v)-\mu^{0}(v))\cdot\Pr(Z_{i}=z\mid X_{i}=x)\}$
[…]
Under unconfoundedness and overlap, the stratified difference-in-means identifies the CME by taking a weighted average CATE across all strata defined by the values of $v$ based on the conditional probability $\Pr(V_{i}=v\mid X=x)=\Pr(Z_{i}=z\mid X=x)$. The two conditional means, $\mu^{1}(v)$ and $\mu^{0}(v)$, are estimable by data given overlap and random sampling.
IPW.
As an alternative, we can identify the CME by weighting each observation by the inverse of the propensity score, which represents the probability of receiving the treatment given covariates, i.e., $\pi(v)=\Pr(D_{i}=1\mid V_{i}=v)$. The derivation follows:
$\mathbb{E}\Bigl[\tfrac{D_{i}\,Y_{i}}{\pi(V_{i})}-\tfrac{(1-D_{i})\,Y_{i}}{1-\pi(V_{i})}\,\Big|\,X_{i}=x\Bigr]=\mathbb{E}\Bigl[\tfrac{D_{i}\,Y_{i}}{\pi(V_{i})}\,\Big|\,X_{i}=x\Bigr]-\mathbb{E}\Bigl[\tfrac{(1-D_{i})\,Y_{i}}{1-\pi(V_{i})}\,\Big|\,X_{i}=x\Bigr]$
[…]
This expression recovers the same mean difference $\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid X_{i}=x]$ by appropriately weighting each observed $Y_{i}$ with $X_{i}=x$, ensuring that the estimate accounts for differences in treatment assignment probabilities.
The discussion so far has focused on the simple scenario of binary treatments and discrete covariates. However, researchers are often interested in estimating and visualizing the CME when $X$ is continuous. In fact, for the remainder of this book, we assume $X$ is a continuous variable, treating the case of discrete $X$ as a special case. While the intuition behind the identification results remains largely the same, the definition of the CME requires additional care when the treatment is continuous, as we discuss below.
1.3 Continuous Treatments
Following hirano2004propensity, we first adapt the potential outcomes framework to accommodate continuous treatments by considering $d$ as an interval $d=[d_{0},d_{1}]$. This formulation allows us to explore the unit-level dose-response function. We assume that each unit $i$’s potential outcome function, $Y_{i}(d)$, is twice continuously differentiable with respect to $d$ for all $d\in\mathcal{D}$, ensuring that the derivative $\partial Y_{i}(d)/\partial d$ exists and is well defined.
Moreover, we maintain Assumption 1 (random sampling) and Assumption 2 (SUTVA). To link the observed outcome with the potential outcomes, we assume $Y_{i}=Y_{i}(D_{i})$, with $D_{i}$ being the observed dose for unit $i$. For each unit $i$, given a vector of covariates $Z_{i}$ and a moderator $X_{i}$, we define the dose–response function $\mu(d,x,z)=\mathbb{E}\left[Y(d)\mid X=x,Z=z\right]$, which describes how the expected potential outcome varies with the treatment level $d$, conditional on the covariates. Let $\mu_{d}(d,x,z)=\partial\mu(d,x,z)/\partial d$ denote its derivative with respect to $d$. The CME with a continuous treatment is defined as the expected value of $\mu_{d}(d,x,z)$ among units with $X=x$, where the expectation is taken over the joint distribution of both $D$ and $Z$:
Definition 2 (CME with Continuous Treatments)
$\theta(x)\;=\;\mathbb{E}\!\big[\,\mu_{d}(D,X,Z)\ \big|\ X=x\big].$
Intuitively, $\mu_{d}(D_{i},X_{i},Z_{i})$ measures how the expected outcome for unit $i$ changes with a small increase in the treatment level around $D_{i}$, holding its covariates fixed. The CME, $\theta(x)$, averages these unit-specific rates of change across units with the same value of the moderator $x$. As before, we also need unconfoundedness and overlap assumptions to identify the CME.
In the context of a continuous treatment, the term “marginal effect” can be particularly ambiguous. In economics and political science, it often refers to the effect of an infinitesimal change in a regressor on the expected outcome, that is, the partial derivative of $\mathbb{E}[Y\mid X]$ with respect to a specific covariate. In this sense, “marginal” is analogous to “slope,” “trend,” or “velocity” (arel2024interpret). The term is especially common in nonlinear or discrete outcome models, where the marginal effect depends on the derivative of the link function. For example, in a probit model
$\Pr(Y=1\mid X)=\Phi(X\beta),$
where $\Phi(\cdot)$ is the standard normal CDF, the marginal effect of $x_{k}$ is
$\frac{\partial\Pr(Y=1\mid X)}{\partial x_{k}}=\phi(X\beta)\beta_{k},$
with $\phi(\cdot)$ denoting the standard normal density. Unlike the linear case, this effect varies with $X$ because $\phi(X\beta)$ depends on the predicted probability. Researchers sometimes report the estimated marginal effect evaluated at the sample mean, $\phi(\bar{X}\hat{\beta})\hat{\beta}_{k}$, as a key quantity of interest.
What adds to the confusion is that the two definitions coincide under the traditional linear interaction model, where the influence of $Z$ on $Y$ are additive and separated from those from $D$ and $Z$ (see details in the next chapter). However, our identification and estimation of the CME do not require linearity or additivity. In this Element, we follow the statistical tradition and use “marginal” to indicate that the expectation is taken over the joint distribution of all right-hand-side variables except $X$, integrating out the treatment $D$ and other covariates $Z$.
Assumption 5 (Unconfoundedness for continuous dose)
$Y_{i}(d)\perp\!\!\!\perp D_{i}\mid V_{i}=v,\quad\text{for all }d\in\mathcal{D}\text{ and }v\in\mathcal{V}$
This assumption, introduced as “weak unconfoundedness” in hirano2004propensity, requires only pairwise conditional independence between the treatment and each potential outcome. In contrast, the unconfoundedness in Assumption 3 imposes joint independence across all potential outcomes. In other words, weak unconfoundedness assumes that, conditional on covariates, treatment assignment is independent of each potential outcome separately rather than of the entire set of potential outcomes jointly.
Assumption 6 (Strict overlap for continuous $D$)
$\quad f_{D|V}(d\mid v)>0,\quad\text{for all }d\in\mathcal{D}\text{ and }v\in\mathcal{V}$
In the continuous treatment case, identifying the CME requires additional assumptions on functional form, which we will discuss in the subsequent chapters along with estimation strategies.
1.4 Approach and Organization
Now that we have clarified the primary estimand, the CME, and the key identification assumptions, the goal of this Element is to discuss identification results, estimation strategies, and inferential methods suitable for various empirical scenarios. These methods progressively relax functional form restrictions but require larger datasets. We will illustrate these approaches using empirical examples from political science.
Throughout this Element, we will primarily focus on unconfoundedness as the main identification strategy for estimating the effect of $D$ on $Y$, which is particularly well-suited for cross-sectional data. We make this choice because most panel data methods commonly used in the social sciences rely on fundamentally different identifying assumptions, such as parallel trends, and their formal treatment using the DML approach falls outside the scope of this Element. Nonetheless, we can accommodate panel data either by first differencing the outcome or by partialing out the fixed effects using the Frisch–Waugh–Lovell (FWL) theorem.
This Element is organized as follows: Chapter 2 reviews classic approaches to estimating the CME. We start with the widely used parametric linear interaction model. While this approach is versatile and easy to implement, it often suffers from issues such as lack of overlap and an inability to capture nonlinear CME, as pointed out by HMX (2019). To address these challenges, we first emphasize a design phase to improve overlap. We then review alternative methods proposed by HMX (2019), including the binning estimator as a diagnostic tool and the semiparametric kernel estimator, which relaxes the functional form assumption. We illustrate these challenges and solutions using several empirical examples. While the kernel estimator offers greater flexibility, it remains an outcome modeling approach that does not account for the treatment assignment process.
Chapter 3 introduces IPW and AIPW estimators for estimating the CME when $X$ is continuous, addressing limitations of traditional outcome modeling approaches. IPW estimates the CME by weighting outcomes using the inverse of the estimated propensity score, while AIPW combines outcome and propensity models to achieve double robustness. We detail how to construct signals from data using outcome modeling, IPW, and AIPW approaches, which serve as key building blocks for the DML methods developed in Chapter 4. To relax functional form assumptions, we incorporate basis expansions, Lasso regularization, and post-selection estimation, which we refer to as AIPW-Lasso. We extend this approach to continuous treatments using partially linear regression and a partialing-out strategy, constructing “denoised” variables before applying kernel or spline regression. We evaluate the performance of these estimators through simulated and empirical applications, demonstrating the advantages of each component in improving robustness, flexibility, and accuracy.
In Chapter 4, we formally introduce the DML framework, which extends AIPW to accommodate high-dimensional or nonlinear nuisance functions using machine learning methods. While AIPW leverages signals that satisfy Neyman orthogonality and achieves robustness against model misspecification, DML ensures $\sqrt{n}$-rate consistency even when nuisance functions are estimated with flexible machine learning models. The chapter outlines three key components of DML: (i) Neyman orthogonality, which ensures that small errors in estimating nuisance functions do not introduce first-order bias; (ii) high-quality machine learning methods, such as random forests, neural networks, and histogram gradient boosting, which approximate nuisance functions while controlling regularization bias; and (iii) sample splitting and cross-fitting, which mitigate overfitting by estimating nuisance functions and treatment effects on separate data folds.
We then apply DML to estimating CME under both binary and continuous treatments. For binary treatments, the estimator builds on the interactive regression model and uses an AIPW-style Neyman orthogonal score to estimate the CATE, which is then projected onto the moderator $X$ to obtain the CME. For continuous treatments, DML adopts a partially linear regression, using a residual-on-residual approach to partial out the influence of covariates before estimating the CME. The chapter concludes with empirical applications, demonstrating DML in both binary and continuous treatment settings. Comparisons with traditional estimators highlight DML’s advantages in capturing complex relationships while maintaining valid inference. Chapter 4 also addresses an important use case where the outcome variable $Y$ is discrete. We demonstrate that double nonlinearity—nonlinearity in the outcome and nonlinearity in the CME—can be better addressed using DML methods than parametric approaches.
An important unanswered question is under what conditions the DML approach outperforms the kernel estimator or AIPW-Lasso in common research settings. To explore this, Chapter 5 presents a series of Monte Carlo simulations. Our results show that even under fairly complex DGPs, AIPW-Lasso (with basis expansion and regularization) generally outperforms DML estimators in small to medium-sized datasets. DML performs better when the sample size is sufficiently large (e.g., at least 5,000 observations). These findings highlight a fundamental trade-off between model flexibility and the ability to handle smaller samples. Simply put, DML offers greater flexibility but requires larger datasets to perform well, while AIPW-Lasso provides a desirable middle ground in most practical applications.
In Chapter 6, the final chapter, we summarize our findings and provide practical recommendations for researchers, drawing on the insights gained throughout the Element. All the methods discussed in this Element can be implemented using the interflex package in R.
1.5 Summary
Understanding conditional relationships is a common objective in social science research. Researchers are often interested in how the effect of a treatment $D$ on an outcome $Y$ depends on a moderating variable $X$. While the methodological literature has made significant progress in this area, current approaches face several challenges, including a lack of clarity regarding the intended estimand, misinterpretation of estimation results, insufficient overlap in real-world data, rigid functional form assumptions, and difficulties in handling complex scenarios such as discrete outcomes. This Element introduces a simple but general framework that addresses these challenges while ensuring statistical validity.
Our key estimand is the conditional marginal effect (CME). When the treatment is binary, it is defined as
$\theta(x)=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid X_{i}=x],$
which represents the CATE marginalized over all covariates except $X$. When the treatment is continuous, the CME is defined as
$\theta(x)=\mathbb{E}[\mu_{d}(D,X,Z)\ \big|\ X=x],$
where $\mu_{d}(d,x,z)$ is the partial derivative of the dose–response function with respect to the treatment level $d$. In other words, $\theta(x)$ represents the average rate at which the expected potential outcome changes with the treatment intensity, after averaging over all covariates except $X$ and the observed treatment levels $D$. In addition, researchers may be interested in effect modification,
$\Delta\theta(x_{1},x_{2})=\theta(x_{1})-\theta(x_{2}),$
which quantifies the difference in treatment effects between two values of the moderator, $x_{1}$ and $x_{2}$. This provides a descriptive (rather than causal) measure of the extent to which treatment effects are heterogeneous.
In cross-sectional settings, the unconfoundedness and overlap assumptions are crucial for identifying the CME. When the treatment or covariates, including the moderator, contain continuous variables, additional functional form or smoothness assumptions are needed for causal identification. In the following chapters, we present a progression of methods for identifying and estimating these quantities, starting with the traditional linear interaction model and gradually moving toward more flexible approaches that accommodate more complex DGPs.
Chapter 2. The Classic Approaches
In this chapter, we discuss classic approaches to estimating and visualizing the CME. We begin with the linear interaction model, which remains widely used in social science research. We introduce the variance estimator for the CME based on this model, which forms the basis for hypothesis testing. We also address the issue of multiple comparisons and explain how uniform confidence intervals can be constructed using a bootstrap procedure to mitigate this concern.
Next, we examine two key challenges in applying the linear interaction model: violations of the overlap assumption (or lack of common support) and model misspecification (such as omitted interactions and nonlinearity). We discuss simple diagnostic tools and strategies to detect and address these issues.
Finally, to relax functional form assumptions, we explore the kernel estimator proposed by HMX (hainmueller2019much) and introduce several improvements, including a fully interacted specification and an adaptive kernel. Throughout this chapter, we emphasize the importance of clearly defining the estimand and explicitly stating identifying and modeling assumptions.
2.1 Linear Interaction Model
In the previous chapter, we considered the CME with a binary treatment $D\in\{0,1\}$, a discrete moderator $X$, and additional discrete covariates $Z$. Under the unconfoundedness assumption, the empirical estimand, which only involves observable quantities (lundberg2021your), of the CME at a particular value $X_{i}=x$ is:
$\mathbb{E}_{z}[\mathbb{E}[Y_{i}\mid D_{i}=1,X_{i}=x,Z_{i}=z]]-\mathbb{E}_{z}[\mathbb{E}[Y_{i}\mid D_{i}=0,X_{i}=x,Z_{i}=z]],$
where $\mathbb{E}[Y_{i}\mid D_{i}=d,X_{i}=x,Z_{i}=z]$, $d\in\{0,1\}$ is a conditional expectation function that could take an arbitrary functional form. As discussed previously, when $D$ is binary and the covariates are discrete and limited in number, it is feasible to estimate conditional expectations nonparametrically. However, as the number of covariates increases or continuous variables are introduced, nonparametric estimation becomes impractical. To address this, we can impose modeling assumptions that restrict the functional form of $\mathbb{E}[Y_{i}(d)\mid X_{i},Z_{i}]$. Combining with the unconfoundedness assumption, the functional form assumption simplifies both estimation and interpretation.
The linear interaction model is the most commonly used approach in political science for studying marginal effects. We formally incorporate this widely used specification as a functional form assumption:
Assumption 7 (Linear interaction model, or the linear estimator)
$Y_{i}(d)=\beta_{0}+\beta_{1}d+\beta_{2}X_{i}+\beta_{3}(d\cdot X_{i})+Z_{i}^{\top}\beta_{4}+\epsilon_{i}$
where $\epsilon_{i}$ represents an idiosyncratic error.
Given Assumption 7, the unconfoundedness assumption implies:
$\{\beta_{0}+\beta_{1}d+\beta_{2}X_{i}+\beta_{3}(d\cdot X_{i})+Z_{i}^{\top}\beta_{4}+\epsilon_{i}\}\perp\!\!\!\perp D_{i}\mid X_{i},Z_{i}$
which further implies that $\epsilon_{i}\perp\!\!\!\perp D_{i}\mid X_{i},Z_{i}$. Combined with the model specification, it leads to the zero-conditional-mean assumption: $\mathbb{E}[\epsilon_{i}\mid D_{i},X_{i},Z_{i}]=0$. Therefore, if the Assumption 7 is correct, the coefficients in the linear interaction model can be consistently estimated using Ordinary Least Squares (OLS) under unconfoundedness.
Given unconfoundedness, Assumption 7 (the linear interaction model) implies the following conditional expectation function for the observed outcome:
$\mathbb{E}[Y_{i}\mid D_{i},X_{i},Z_{i}]=\beta_{0}+\beta_{1}D_{i}+\beta_{2}X_{i}+\beta_{3}(D_{i}\cdot X_{i})+Z_{i}^{\top}\beta_{4}$ (3)
The binary treatment case.
When $D_{i}$ is binary, under the unconfoundedness assumption and the Assumption 7, Equation (3) is equivalent to the following conditional expectations:
$\mathbb{E}[Y_{i}\mid D_{i}=0,X_{i},Z_{i}]$ $=\beta_{0}+\beta_{2}X_{i}+Z_{i}^{\top}\beta_{4}$
$\mathbb{E}[Y_{i}\mid D_{i}=1,X_{i},Z_{i}]$ $=(\beta_{0}+\beta_{1})+(\beta_{2}+\beta_{3})X_{i}+Z_{i}^{\top}\beta_{4}$
These equations imply linear relationships between $X$ and the expected outcome, as well as between $Z$ and the expected outcome. In addition, the coefficients for $Z$ are constant across treatment values, and there are no interactions between $X$ and $Z$. The CME can be written as:
$\theta(x)$ $=\mathbb{E}_{z}[\beta_{0}+\beta_{1}+\beta_{2}x+\beta_{3}\cdot x+Z_{i}^{\top}\beta_{4}]-\mathbb{E}_{z}[\beta_{0}+\beta_{2}x+Z_{i}^{\top}\beta_{4}]$
$=(\beta_{0}+\beta_{1}+\beta_{2}x+\beta_{3}\cdot x+\beta_{4}^{\top}\mathbb{E}_{z}[Z_{i}])-(\beta_{0}+\beta_{2}x+\beta_{4}^{\top}\mathbb{E}_{z}[Z_{i}])$
$=\beta_{1}+\beta_{3}x$
which is a linear combination of $\beta_{1}$ and $\beta_{3}$. Therefore, $\hat{\theta}(x)=\hat{\beta}_{1}+\hat{\beta}_{3}x$. This logic can be easily extended to discrete $D_{i}$.
The continuous treatment case.
In the previous chapter, we defined the CME for continuous treatments as the expected partial derivative of the dose-response function, i.e., $\mu(d,x,z)=\mathbb{E}\left[Y(d)\mid X=x,Z=z\right]$, where $\mu_{d}(d,x,z)=\partial\mu(d,x,z)/\partial d$, and $\theta(x)\;=\;\mathbb{E}\!\big[\,\mu_{d}(D,X,Z)\ \big|\ X=x\big]$. Under Assumption 7, the potential outcome is linear in $d$ with a slope of $\beta_{1}+\beta_{3}X_{i}$. Consequently, the partial derivative with respect to $d$ is deterministic given $X=x$, and the CME simplifies to:
$\theta(x)=\beta_{1}+\beta_{3}x$
This yields the same functional form as the binary treatment case. Unconfoundedness is crucial here because it ensures $\epsilon_{i}\perp\!\!\!\perp D_{i}\mid X_{i},Z_{i}$, allowing us to use OLS to consistently estimate the key coefficients $\beta_{1}$ and $\beta_{3}$.
Inference and Hypothesis Testing
The above results suggest that, under the linear interaction model, the CME is a simple, linear function of $x$. For statistical inference, 95% pointwise confidence intervals are commonly used in empirical research. By 95% pointwise confidence interval, we mean that at each fixed value of $x$, 95% of such intervals would contain the true CME under repeated sampling. These intervals are calculated independently for each value of $x$, without considering the joint coverage probability across multiple values of $x$.
We can use one of the following two approaches to construct the pointwise confidence intervals, both implemented in interflex: (i) analytical asymptotic variance based on normal approximation or (ii) bootstrapping, which involves repeatedly resampling the data and re-estimating $\hat{\theta}(x)$. Analytically, the variance of $\hat{\theta}(x)$ can be estimated as follows:
$\widehat{\operatorname{Var}}(\hat{\theta}(x))=\widehat{\operatorname{Var}}(\hat{\beta}_{1})+x^{2}\widehat{\operatorname{Var}}(\hat{\beta}_{3})+2x\widehat{\operatorname{Cov}}(\hat{\beta}_{1},\hat{\beta}_{3}).$
in which $\widehat{\operatorname{Var}}(\hat{\beta}_{1})$, $\widehat{\operatorname{Var}}(\hat{\beta}_{3})$ and $\widehat{\operatorname{Cov}}(\hat{\beta}_{1}\hat{\beta}_{3})$ are often obtained using the Eicker-Huber-White robust estimator or the cluster-robust estimator (if a cluster structure exists).
In their seminal work, BCG (2006) recommend visualizing the CME, which they refer to as the marginal effects, using a plot where $x$ is placed on the x-axis and $\hat{\theta}(x)$, derived from the linear interaction model, is displayed on the y-axis. See Figure 1 in Chapter 1 for example. The plot also includes the 95% pointwise confidence interval to provide a visual representation of the uncertainty associated with the estimates. We provide the details of the example and R code to replicate this figure below.
Example 4
huddy2015expressive explores the effect of the expressive model of partisanship. Drawing on four survey studies, the authors argue that partisan identity drives campaign participation and strong emotional responses to ongoing campaign events. We replicate Figure 2A in the original paper, which suggests that strongly identified partisans are more likely to exhibit stronger emotional reactions—such as anger when threatened with electoral loss—compared to those with weaker partisan identities. The variables of interest include:
  • Outcome: level of anger (continuous, $\in[0,1]$)
  • Treatment: electoral loss threat or electoral win reassurance (binary, $\in\{0,1\}$)
  • Moderator: strength of partisan identity (continuous, $\in[0,1]$)
1# R code excerpt
2library(interflex)
3
4Y="totangry" # Anger
5D="threat" # Threat
6X="pidentity" # Partisan Identity
7Z <- c("issuestr2", "knowledge" , "educ" , "male" , "age10" )
8
9## linear interaction model
10out.linear <- interflex(estimator =
11"linear", data = df, Y = Y, D = D, X = X, Z = Z,
12vartype = "bootstrap", na.rm = TRUE)
13
14# plot the CME
15plot(out.linear)
Figure 2: Replicating huddy2015expressive Figure 2A
Notes: The black line represents the CME estimates based on the linear interaction model. The shaded area represent 95% pointwise confidence intervals. The dashed lines represent 95% uniform confidence intervals. The histograms at the bottom of the figure depict the distributions of $X$ across treatment (pink) and control (gray) groups.
As mentioned in Chapter 1, researchers are often interested in testing three types of hypotheses:
  • $H_{0a}:\theta(x)=0$ given a specific $x$. Researchers ask whether, at this particular value of $x$, the CME is statistically distinguishable from zero.
  • $H_{0b}:\Delta\theta(x_{1},x_{2})=0$ or $\Delta\theta(x_{1},x_{2})>0$ given specific values of $x_{1}$ and $x_{2}$. Researchers compare two CMEs and determine whether they are statistically distinguishable or whether one CME is larger than the other.
  • $H_{0c}:\theta(x)=0$ (or $\theta(x)\gtrless 0$), $x\in\mathcal{X}$, a pre-specified set. Researchers investigate whether the CME across a set or region of $x$ values is jointly statistically distinguishable from zero or jointly larger or smaller than zero.
To test $H_{0a}$, we can perform a $t$-test using $\hat{\theta}(x)$ and $\widehat{\operatorname{Var}}(\hat{\theta}(x))$. This is equivalent to visually inspecting whether the pointwise confidence interval in the marginal effect plot intersects the horizontal zero line. For instance, in Figure 2, at $x=0.5$, the pointwise confidence interval does not overlap with zero. This allows us to reject the null hypothesis that $\theta(0.5)=0$ at the 5% significance level.
To test $H_{0b}$, we first estimate $\Delta\theta(x_{1},x_{2})$ using the following approach:
$\widehat{\Delta\theta}(x_{1},x_{2})$ $=\hat{\theta}(x_{1})-\hat{\theta}(x_{2})$
$=(\hat{\beta}_{1}+\hat{\beta}_{3}x_{1})-(\hat{\beta}_{1}+\hat{\beta}_{3}x_{2})$
$=\hat{\beta}_{3}(x_{1}-x_{2}).$
The variance of this estimate is given by: $\widehat{\operatorname{Var}}(\widehat{\Delta\theta}(x_{1},x_{2}))=\widehat{\operatorname{Var}}(\hat{\beta}_{3})(x_{1}-x_{2})^{2}$. Because $(x_{1}-x_{2})$ is a constant, testing $H_{0b}$ is equivalent to testing whether $\beta_{3}=0$. Indeed, researchers often interpret the statistical significance of $\hat{\beta}_{3}$, the coefficient of the interaction term, as evidence for heterogeneous treatment effects.
To test $H_{0c}$, we need to construct uniform confidence intervals. Before doing so, we first discuss the issue of multiple comparisons.
Multiple Comparisons
esarey2018marginal highlight a multiple comparisons issue associated with CME plots. They argue that BCG (2016)’s suggestion to use a visual test—examining whether the (pointwise) confidence interval includes zero at a given value of the moderator $X$ to conclude that the treatment $D$ and outcome $Y$ are statistically related at that value—is overly optimistic. This approach does not account for the multiple comparisons problem, which arises when making inferences across a range of moderator values, leading to inflated Type I error (or uncontrolled false positive rates).
In the context of the linear interaction model, using an example with a binary moderator, esarey2018marginal point out that the visual test is equivalent to conducting hypothesis testing twice, once in each subsample defined by $X$, with each test carrying its own probability of a false positive. This issue becomes even more pronounced when the treatment $D$ or the moderator $X$ is continuous, as the number of comparisons increases significantly, further compounding the likelihood of false positives. esarey2018marginal recommend applying procedures that control the overall false discovery rate or familywise error rate, such as the sequential test procedure by benjamini1995controlling or the joint $F$-test proposed by franzese2009modeling.
In the interflex package, we address the multiple comparisons issue in estimating CME by incorporating uniform confidence intervals, also known as simultaneous confidence intervals. These intervals ensure a specified overall coverage level across all estimates, controlling the family-wise error rate and addressing the multiple comparisons problem. They are implemented using the sup-$t$ band approach introduced by montiel2017simultaneous using bootstrapping, which is a simulation-based method that involves repeatedly resampling the data with replacement to estimate the sampling distribution of an estimator. The procedure takes the following steps:
  1. 1.
    We discretize the pre-specified set of the moderator $X$, $\mathcal{X}$, into a grid of $k$ evaluation points, denoted as $\{x_{1},x_{2},\dots,x_{k}\}$. For each bootstrap replication $b\in\{1,\dots,B\}$, we compute the vector of CME estimates on this grid: $\hat{\theta}^{(b)}=(\hat{\theta}^{(b)}(x_{1}),\dots,\hat{\theta}^{(b)}(x_{k}))$.
  2. 2.
    For each grid point $j\in\{1,\dots,k\}$, we can obtain the empirical quantiles from the bootstrap estimates: let $Q_{j,\zeta}$ and $Q_{j,1-\zeta}$ denote the $\zeta$-th and $(1-\zeta)$-th quantiles of the distribution $\{\hat{\theta}^{(b)}(x_{j})\}_{b=1}^{B}$.
  3. 3.
    We determine the largest $\zeta$ such that the rectangular band formed by these quantiles covers the entire true CME curve with probability at least $\alpha:11$1The lower bound of $\zeta$, $\frac{\alpha}{2k}$, corresponds to the quantiles with the Bonferroni correction, while the upper bound, $\frac{\alpha}{2}$, corresponds to the quantile of the pointwise confidence interval.
    $\zeta^{*}=\sup\left\{\zeta\in\left[\frac{\alpha}{2k},\frac{\alpha}{2}\right]\middle|\frac{1}{B}\sum_{b=1}^{B}\mathbf{1}\left(\forall j\in\{1,\dots,k\}:\hat{\theta}^{(b)}(x_{j})\in[Q_{j,\zeta},Q_{j,1-\zeta}]\right)\geq 1-\alpha\right\}$
  4. 4.
    The uniform confidence intervals are constructed using this optimal $\zeta^{*}$, resulting in the band $\hat{Q}_{j,\zeta^{*}},\hat{Q}_{j,1-\zeta^{*}}]$ at each grid point $x_{j}$.
This method ensures that the confidence intervals are uniformly valid across all evaluation points in $\mathcal{X}$. In other words, the uniform confidence band covers the true CME function $\theta(x)$ simultaneously for all $x\in\mathcal{X}$ with probability at least $1-\alpha$. Consequently, if the uniform confidence band excludes zero at any point $x\in\mathcal{X}$ (i.e., if the horizontal zero line does not lie entirely within the band), we can reject the joint null hypothesis $H_{0c}:\theta(x)=0,\forall x\in\mathcal{X}$ at the $\alpha$ significance level.
In Figure 2, the dotted lines represent the uniform confidence intervals ($\mathcal{X}$ is set to the support of $X$), while the shaded area depicts the pointwise confidence intervals. The uniform confidence intervals are consistently wider, reflecting a more conservative approach to interval estimation that adjusts for multiple comparisons across the moderator’s range. We observe that the lower bounds of the uniform confidence intervals cross the horizontal zero line around $x=0.25$, suggesting that for values of the moderator $X$ greater than 0.25, the effect of the treatment becomes statistically distinguishable from zero at the 5% level, indicating a significant treatment effect. [^2] Using a uniform confidence band constructed over the entire support of $X$ to test hypotheses about a specific sub-region (e.g., $x>0.25$) is statistically valid but conservative. If the research hypothesis had specified this sub-region a priori, a narrower band could have been constructed specifically for that range, potentially increasing power (see esarey2018marginal, for a detailed discussion on controlling false positive rates in this context). Therefore, a visual test using the uniform confidence intervals provides a practical method for performing hypothesis testing for $H_{0c}$.
Using the linear interaction model to estimate the CME presents several challenges, primarily lack of common support and model misspecification, as discussed in HMX (hainmueller2019much). In the remainder of this chapter, we illustrate these challenges with examples and propose several solutions to address them.
2.2 Lack of Common Support
HMX (hainmueller2019much) highlight that the overlap assumption is often violated in applications, particularly in observational studies, a problem commonly referred to as a “lack of common support.” Researchers using the linear interaction model typically report the CME by substituting $x$ values into the conditional marginal effect formula, $\hat{\theta}(x)=\hat{\beta}_{1}+\hat{\beta}_{3}x$, without checking for overlap. This occurs because regression produces estimates regardless of whether the overlap assumption holds.
When $X$ is continuous, the overlap assumption requires two conditions: (1) a sufficient number of observations with values of the moderator $X$ close to $x_{0}$, and (2) sufficient variation in the treatment $D$ at or around $x_{0}$ for similar values of $Z$. For example, if all data points near $X=x_{0}$ belong to the treatment group ($D=1$), overlap is violated. If either of these conditions is not met, CME estimates must rely on extrapolation based on the assumed functional form, extending the model to regions with little or no data. In such cases, the CME estimates become highly model-dependent and unreliable. We provide such an example below.
Example 5
chapman2009audience examines the effect of international organization, specifically the authorizations granted by the U.N. Security Council, on public opinion of U.S. foreign policy. The authors use the size of rallies, which measures the short-term change in presidential approval ratings surrounding military disputes, as a proxy for support for foreign policy. We replicate the Figure 2 of the original paper, which suggests the effect of U.N. authorizations is conditional on similarity of preferences between the U.S and the U.N. Security Council: the public is more likely to favor policies that are explicitly approved by relatively conservative institutions, i.e., institutions with more heterogeneous preferences. The variables of interest include:
  • Outcome: size of rallies supporting the authorization (continuous, $\in[-16,33]$)
  • Treatment: granting of a U.N. authorization (binary, $\in\{0,1\}$)
  • Moderator: preference distance between the U.S. and the Security Council (continuous, $\in[-1,0]$)
1# R code excerpt
2
3library(interflex)
4
5Y="rally" # size of rallies
6D="unauth" # U.N. authorization
7X="S" # US affinity with UN Security Council
8Z=c("priorpop","bipart","statemnt","nextelec","nytcov","buscon","revorg","wardumk","majopp","allies","war","SCappeal","regorgact","ordsever","hostlvl")
9
10## linear estimator with uniform CI
11out.linear<-interflex(estimator = "linear",
12 Y=Y,D=D,X=X,Z=Z, data=d,
13 vartype = "bootstrap", na.rm = TRUE)
14plot(out.linear)
15
16## diagnostic plot
17out.raw <- interflex(estimator = "raw", data = d,Y=Y,D=D,X=X, na.rm = TRUE)
18plot(out.raw)
19
20## propensity score computation
21V<-append(Z, X)
22covariate_matrix <- d[V]
23covariate_matrix[] <- lapply(covariate_matrix, function(x) as.numeric(as.character(x)))
24covariate_matrix<-as.matrix(covariate_matrix)
25
26cf<-causal_forest(covariate_matrix, d\$rally, d\$unauth)
27e.hat <- cf\$W.hat
Figure 3: Replicating chapman2009audience Figure 2
(a) CME plot
(b) Outcome by treatment condition
(c) Propensity score
Notes: Treatment measures UN authorization, outcome is the size of rallies, measuring the short-term change in presidential approval ratings surrounding military disputes. (a) The CME plot. The black line represents the CME estimates based on a linear interaction model. The shaded area represents pointwise confidence intervals, while the dotted line indicates uniform confidence interval, both of which are calculated via bootstrap. (b) Raw outcome plot by treatment condition; (c) Propensity score, estimated with grf package, by treatment condition.
Figure 3(a) presents the CME plot based on the linear interaction model. The author interprets the positive estimated CME when the moderator is smaller than 0.5-0.5 as evidence that the positive effect of U.N. authorizations on U.S. public opinion is conditional on the U.S. policy position is far from those of the countries serving on the Security Council.
However, this finding appears to be entirely driven by extrapolation, and the lack of common support is easy to detect using simple plots. HMX (hainmueller2019much) recommend (i) adding a histogram at the bottom of the CME plot and (ii) using a diagnostic plot, such as Figure 3(b), that displays the raw data by treatment condition. When the treatment is binary, we also recommend plotting the propensity score, a standard approach for assessing overlap (heckman1999economics; imbens2025comparing).
Indeed, the histogram in Figure 3(b) reveals that there are almost no observations when the moderator is below 0.7-0.7. Even more concerning is the lack of treated units (with U.N. authorization), which are concentrated in a narrow range around 0.5-0.5. This indicates a severe lack of overlap in the regions where the author interprets the effect as positive. In fact, as shown in Figure 3(a), the downward slope of the CME estimates is almost entirely driven by six treated units through extrapolation. The propensity score plot in Figure 3(c) further confirms this issue, showing insufficient common support in these regions.
Design Phase to Improve Overlap
Later in this chapter and in the next two chapters, we introduce methods that relax Assumption 7. To address concerns about lack of overlap, we recommend incorporating a design phase prior to any outcome analysis. This approach, proposed by imbens2015causal, reflects a growing consensus in the causal inference literature that study design should precede analysis, even in non-experimental settings.
In observational studies, the design phase refers to a preparatory stage in which researchers structure or preprocess the data before analyzing outcomes. The goal is to approximate the conditions of a randomized controlled trial using only information on covariates and treatment assignment—without referencing the outcome variable. In other words, the design phase should be used to construct a balanced sample or subsamples of treated and control units that are comparable in their covariate distributions, ensuring that “within this selected subsample, inferences are most robust and credible” (imbens2015causal). The design phase directly addresses the lack of overlap by identifying—and, if necessary, removing—observations in regions of the covariate space where treated and control units do not sufficiently overlap. This prevents researchers from extrapolating treatment effects to areas unsupported by the data. Even when overlap is present, the design phase helps improve covariate balance and reduce potential confounding.
To conduct a design phase in practice, we encourage researchers to apply matching, trimming, and/or reweighting methods. Crucially, outcome data should not be used during this phase, and hypothesis testing should be avoided to reduce the risk of $p$-hacking and preserve the validity of subsequent analyses.
It is important to note, however, that the design phase does not resolve all identification or estimation challenges. First, it does not make an unlikely unconfoundedness assumption suddenly valid. Second, it does not address issues related to nonlinearity or heterogeneity in the CME. Even after achieving covariate balance, the relationship between treatment and outcome may still vary with the moderator or other covariates in complex ways. These challenges are addressed during the estimation phase using flexible modeling strategies
To illustrate the value of design-based preprocessing, we reference studies such as noble2024presidential, which show how these methods can effectively standardize treatment groups and improve the reliability of causal inference.
Example 6
noble2024presidential examines how Members of Congress in the United States leverage the president’s symbolic power. The study argues that legislators—particularly those from the out-party—strategically reference the president to nationalize debates, shape constituent opinions, increase in-party approval, and reduce incentives for compromise. The analysis draws on a dataset of nearly two million floor speeches from 1973 to 2016. Here, we replicate the top panel of Figure 2 from the original paper, which shows that out-party legislators reference the president more frequently—and that this difference narrows as constituency support for the president increases. The variables of interest are:
  • Outcome: Frequency of presidential references in Congressional speeches (continuous, $\in\left[0,258\right]$)
  • Treatment: Out-partisanship of lawmakers (binary, $\in\{0,1\}$).
  • Moderator: District past presidential vote margin (continuous, $\in\left[-46.06,46.99\right]$).
Figure 4: Replicating noble2024presidential Figure 2
(a) Full sample
(b) Trimmed sample
Notes: Treatment $D$ is the out-partisanship of lawmakers, outcome $Y$ is the frequency of presidential references in congressional speeches. In each figure, the black dashed line represents the CME estimates; the red points (and bars) represent the binning estimates (and 95% pointwise confidence intervals); the shaded area and dotted lines represent 95% pointwise and uniform confidence intervals, respectively.
Using a large observational dataset, lack of overlap arises naturally due to partisan sorting in congressional districts: districts that overwhelmingly support the president tend to elect fewer out-party legislators, leading to imbalance across covariates—particularly the moderator. Figure 4(a) shows CME estimates based on the linear estimator (left panel), overlaid with binning estimates, and the kernel estimator (right panel). At the bottom of each panel, we plot histograms of the moderator by treatment condition. The plots reveal limited common support at the extremes of the moderator (when $X\notin[-10,20]$). Despite this, the linear estimator still produces estimates in these regions, which are not credible as they rely heavily on extrapolation. In contrast, the kernel estimator reflects this lack of support through very wide confidence intervals. It is worth noting that the author is clearly aware of this issue and presents CME estimates over the range $x\in[-4,16]$, using the linear estimator, although the estimation is based on the full sample.
To ensure reasonable overlap, we trimmed the data based on the central 90% percentile range of the propensity score $\hat{\pi}(V)=\Pr(D=1\mid X,Z)$. After trimming, both estimators exhibit more stable and nearly linear patterns, as shown in Figure 4(a). The narrower confidence intervals and smoother trends reflect improved covariate balance between treated and control units in the trimmed sample. The resulting estimates support the original finding: as a constituency becomes more supportive of the president, the gap between out-party and in-party lawmakers in referencing the president in congressional speeches narrows. This example shows that data trimming during the design phase improves the credibility of the estimates and may even help justify simpler models.
2.3 Misspecification
The linear interaction model makes rigid parametric assumption and can lead to misleading conclusions when misspecified. Below, we discuss two key scenarios where this occurs: (1) missing interaction terms between $X$ and $Z$ and (2) a nonlinear CME.
Missing Interaction Terms
blackwell2022reducing and feigenberg2025omitted highlight that the linear interaction model (as specified in Assumption 7) may omit potential interactions between the moderator $X$ and other covariates $Z$, which are likely to be correlated in observational studies where neither $D$ nor $X$ is randomized. They refer to this specification as the single-interaction model, contrasting it with the fully-moderated model, where $X_{i}$ is interacted with all covariates:
$Y_{i}(d)=\beta_{0}+\beta_{1}d+\beta_{2}X_{i}+\beta_{3}(d\cdot X_{i})+Z_{i}^{\top}\beta_{4}+(X_{i}\cdot Z_{i})^{\top}\beta_{5}+\epsilon_{i}.$
The linear interaction model produces inconsistent estimates when (a) the treatment–moderator interaction is predictive of omitted interactions and (b) the omitted interactions significantly influence the outcome (beiser2020problems). This issue arises, for example, when the heterogeneous effect of $D$ on $Y$ depends not only on $X$ but also on $Z$, yet $Z$ is included only as a level term.
To address this issue, blackwell2022reducing recommend using a fully moderated model combined with a post-double-selection (PDS) procedure, originally introduced by belloni2014inference. This method involves three steps: (i) applying Lasso to select interactions between $X_{i}$ and $Z_{i}$ that predict the outcome $Y$, the treatment $D$, and the interaction $D\cdot X$; (ii) taking the union of the selected variables from each model; and (iii) running a post-Lasso regression using this union to estimate the causal effect. The key advantage of PDS-Lasso is that it mitigates regularization bias by using Lasso only for variable selection, while relying on unpenalized OLS in the final step. This preserves interpretability and guards against omitted variable bias without distorting parameter estimates.
The kernel estimator introduced in the next section is fully moderated and thus immune to this criticism, provided the number of additional covariates is limited.
In Chapter 3, we introduce Lasso-based methods that are similar in spirit but differ in two key ways. First, they incorporate more flexible basis expansions of $X$ and $Z$, as well as the interaction of the basis expansions, to accommodate higher-order terms and interactions among covariates. Second, for binary treatments, we propose an AIPW estimator, and for continuous treatments, we use a partialing-out approach—both of which go beyond the post-Lasso regression strategy described here.
Nonlinear CME
As discussed earlier, while the linear interaction model allows the effect of $D$ on $Y$ to vary with $X$, it imposes the restriction that $\theta(x)=\beta_{1}+\beta_{3}x$. When $D$ is continuous, this implies that the effect of $D$ on $Y$ changes with $X$ at a constant rate, determined by $\beta_{3}$. However, in many empirical settings, this assumption is unrealistic, as treatment effects are rarely strictly linear in a covariate and may not even be monotonic. HMX (hainmueller2019much) highlight this limitation and emphasize that violating this assumption can lead to misinterpretation of the CME.
Binning estimator as a diagnostic tool.
To diagnose potential nonlinearity in the CME, HMX (hainmueller2019much) propose the binning estimator, which partitions the range of the continuous moderator $X$ into intervals and estimates the CME at an evaluation point within each interval. For simplicity, consider dividing $X$ into three bins using cutoff points $\delta_{1/3}$ and $\delta_{2/3}$, representing the first and second terciles of the $X$ distribution. Define three dummy variables:
$G_{1,i}=\begin{cases}1&X_{i}<\delta_{1/3}\\ 0&\text{otherwise}\end{cases},\ G_{2,i}=\begin{cases}1&\delta_{1/3}\leq X_{i}<\delta_{2/3}\\ 0&\text{otherwise}\end{cases},\ G_{3,i}=\begin{cases}1&X_{i}\geq\delta_{2/3}\\ 0&\text{otherwise}.\end{cases}$
Within each bin, we select an evaluation point $x_{j}$ (such as the median of $X$ values in that bin) and model the outcome $Y$ as a piecewise linear function of $X$, interacting with the treatment $D$. We then estimate the following model using OLS:
$Y_{i}=\sum_{j=1}^{3}\left[\mu_{j}+\alpha_{j}D_{i}+\eta_{j}(X_{i}-x_{j})+\beta_{j}(X_{i}-x_{j})D_{i}\right]\cdot G_{j,i}+Z_{i}^{\top}\gamma+\epsilon_{i},$
where $\mu_{j},\alpha_{j},\eta_{j},\beta_{j}$ are unknown parameters to be estimated, $Z$ represents additional covariates, and $\epsilon$ is an error term with $\mathbb{E}[\epsilon_{i}|X_{i}=x,D_{i},Z_{i}]=0$. The key insight is that at the chosen evaluation points $x_{j}$, the CME simplifies to $\theta(x_{j})=\alpha_{j}$ since $(X_{i}-x_{j})=0$ at these points.
By allowing $\alpha_{j}$ to vary across bins, the binning estimator accommodates flexible CME patterns across different segments of the moderator. Unlike the linear interaction model, which imposes a single global linear relationship on the CME, this piecewise approach relaxes the linearity assumption on the CME and can capture more nuanced patterns. At the same time, the binning estimator remains easier to implement than fully nonparametric or semiparametric methods, as it relies on standard regression techniques and well-established inference methods. Standard errors for $\theta(x_{j})$ are obtained from the regression output, eliminating the need for additional estimation steps or complex inference procedures.
We want to emphasize that the binning estimator should be used primarily as a diagnostic tool. With a fixed, coarse partition, the estimates $\alpha_{j}$ provide a “local” test of the treatment effect at the chosen evaluation points $x_{j}$ (typically the bin medians), allowing us to visually assess whether the linear functional form is appropriate without imposing it globally.
Example 7
clark2006rehabilitating examine the effect of the temporal proximity of presidential elections on the number of political parties in legislative elections. They argue that presidential elections, as the most important election in a polity, are most likely to exert the strongest influence when held concurrently with legislative elections. The direction of this “coattails effect” is moderated by the number of presidential candidates: temporal proximity to presidential elections reduces the number of parties contesting legislative elections when the number of presidential candidates is small, because parties that are not competitive in presidential races are disadvantaged. We replicate Figure 2 in the original paper, the variables of interest are:
  • Outcome: number of parties compete in an election (continuous)
  • Treatment: temporal proximity of presidential elections (continuous, $\in[0,1]$)
  • Moderator Variable: number of presidential candidates. (continuous)
1# R code excerpt
2library(interflex)
3
4Y="enep1" #Effective number of electoral parties
5D="proximity1" #Temporal proximity of presidential elections
6X="enpres" #Number of presidential candidates"
7Z=c("uppertier", "logmag", "uppertier_eneg", "eneg", "logmag_eneg")
8
9## diagnostic plot
10cuts_raw <- c(0,1.64,3.28,4.92)
11out.raw <- interflex(estimator = "raw", data = d, Y=Y,D=D,X=X, na.rm = TRUE,
12ncols = 4, cutoffs = cuts_raw)
13plot(out.raw)
14
15## binning estimator with uniform CI
16cuts_bin <- c(0,0.1,3,4,7)
17out.binning <- interflex(estimator = "binning", data = d, Y=Y,D=D,X=X,Z=Z, data=d, vartype ="bootstrap", nbins = 4, cutoffs = cuts_bin)
18plot(out.binning)
Figure 5: Replicating clark2006rehabilitating Figure 2
(a) Raw outcome plot by moderator values
(b) CME plot
Notes: Treatment $D$ is presidential elections, outcome $Y$ is the effective number of electoral parties. (a) Plot of the raw outcome data in each bin of the moderator. (b) The CME estimates based on the linear interaction model (black line) and the binning estimator (red hollow circles). The gray ribbon and red bars show the 95% pointwise confidence intervals, respectively. The dashed lines represent 95% uniform confidence intervals of the linear interaction model.
Because a significant portion of the observations in the sample (59%) have $X=0$, as shown in Figure 5(a), we discretize $X$ into four bins: $\{0\}$, $(0,3]$, $(3,4]$, and $(4,6.57]$. The last three bins contain approximately equal numbers of observations. Figure 5(a) also reveals that the bivariate relationship between $D$ and $Y$ does not appear to increase with the moderator.
Figure 5(b) compares the CME estimates from the linear interaction model and the binning estimator. Contrary to the study’s claim of a positive interaction based on the linear interaction model, the binning estimator indicates highly nonlinear effects. Specifically, the effect is negative and small in the second bin ($X\in(0,3]$), turns negative and statistically significant in the third bin ($X\in(3,4]$), and then approaches zero in the last bin ($X\in(4,6.57]$). This pattern suggests that the linear interaction model is misspecified.
Moreover, the treatment has no variation in the first bin ($X=0$), making the treatment effect non-identifiable at $x=0$ due to a lack of common support. This further highlights the limitations of the linear interaction model in this case. Additionally, the uniform confidence interval for the second bin (M1) covers zero, meaning that we cannot reject the null hypothesis that $\theta(x)<0$ for all $x\leq 3$ at the 5% level, contradicting what the linear interaction model might have suggested.
While we typically do not assume that the CME changes with $X$ in a piecewise manner, this example demonstrates that the binning estimator is an easy-to-implement and powerful diagnostic tool for detecting potential nonlinearity in the CME. To achieve greater flexibility in modeling the functional form of the CME, we turn to more advanced methods, such as the kernel estimator, which we will discuss later in this chapter, and AIPW and DML estimators, which we will explore in the next two chapters.
2.4 The Kernel Estimator
While parametric methods rely on strong and often unrealistic functional form assumptions, and nonparametric models avoid such assumptions but suffer from the “curse of dimensionality,” semiparametric methods offer a middle ground. They relax functional form restrictions by allowing some components of the model to be flexible while maintaining parametric structure in others. HMX (hainmueller2019much) propose a semiparametric kernel estimator based on smooth varying-coefficient models (SVCM) (hastie2017generalized) to estimate the CME. In the remainder of the Element, we refer to this method as the kernel estimator. We introduce two improvements: (i) allowing the coefficient on $Z$ to vary with $X$, making the model fully moderated, and (ii) incorporating an adaptive kernel bandwidth to improve estimation accuracy across different regions of the data distribution.
Assumption 8
Smooth varying-coefficient model , SVCM
$Y_{i}(d)=f(X_{i})+g(X_{i})d+Z_{i}^{\top}\gamma(X_{i})+\epsilon_{i}$
in which $f(\cdot),g(\cdot)$ and $\gamma(\cdot)$ are smooth functions.
SVCM assumes that $Y$ is a function of $D$, $X$, and $Z$, while their impact on $Y$ is modeled as an unspecified smooth function of $X$. At each evaluation point $x_{0}$, each of them is approximated by a local linear regression: [^3] HMX (hainmueller2019much) assume a slightly more restrictive form of SVCM: $Y_{i}(d)=f(X_{i})+g(X_{i})d+Z_{i}^{\top}\gamma+\epsilon_{i}$, where $\gamma$ does not vary with $X$. Assumption 8 represents a fully moderated version of their model.
$f(X)_{|X-x_{0}|<\epsilon}$ $=\mu(x_{0})+\eta(x_{0})(X-x_{0})$
$g(X)_{|X-x_{0}|<\epsilon}$ $=\alpha(x_{0})+\beta(x_{0})(X-x_{0})$
$\gamma(X)_{|X-x_{0}|<\epsilon}$ $=\rho(x_{0})+\delta(x_{0})(X-x_{0})$
As before, given Assumption 8, the unconfoundedness assumption implies $\mathbb{E}[\epsilon_{i}|X_{i},D_{i},Z_{i}]=0$. The CME becomes:
$\theta(x)=\mathbb{E}_{Z}\left[f(x)+g(x)+Z_{i}^{\top}\gamma(x)\right]-\mathbb{E}_{Z}\left[f(x)+Z_{i}^{\top}\gamma(x)\right]=g(x).$
when $D$ is binary. CME for a continuous $D$ similarly yields:
$\theta(x)=\mathbb{E}_{Z}\left[\frac{\partial(f(x)+g(x)d+Z_{i}^{\top}\gamma(x))}{\partial d}\right]=g(x).$
Therefore, our goal is to obtain valid inference for $g(x)$. In SVCM, the intercept $f(x)$, treatment effect $g(x)$, and control coefficients $\gamma(x)$ are all flexible functions of the moderator $X$, allowing for nonlinear CME.
HMX (hainmueller2019much) propose a kernel smoothing estimator for the CME with SVCM. For each evaluation point $x_{0}$ in the support of $X$, $\hat{f}(x_{0})$, $\hat{g}(x_{0})$, and $\hat{\gamma}(x_{0})$ are estimated by minimizing the following weighted least-squares (WLS) objective function. With our two improvements, the kernel estimator that considers the influence of potentially all data points with greater emphasis on points closer to $x_{0}$ has the following weighted least-squares objective function:
[…]
in which:
$L=\sum_{i=1}^{N}\left\{\left[Y_{i}-\tilde{\mu}-\tilde{\alpha}D_{i}-\tilde{\eta}(X_{i}-x_{0})-\tilde{\beta}D_{i}(X_{i}-x_{0})-\tilde{\rho}^{\top}Z_{i}-\tilde{\delta}^{\top}Z_{i}(X_{i}-x_{0})\right]^{2}K\left(\frac{|X_{i}-x_{0}|}{h(x_{0})}\right)\right\}$
Kernel smoothing averages data points using a kernel function, with weights determined by proximity to the estimation point. The kernel function, $K(.)$, assigns higher weights to observations closer to the target point $x_{0}$. With the above kernel estimator, $\hat{f}(x_{0})=\hat{\mu}(x_{0})$ and $\hat{g}(x_{0})=\hat{\alpha}(x_{0})$. Smoothing is performed using a Gaussian kernel $K(\cdot)$, while the parameter of the bandwidth function $h(x_{0})$ is selected via least-squares cross-validation.
Adaptive Kernel
Classical bandwidth selection methods perform well only when the underlying density is approximately normal. They become problematic for long-tailed or multimodal distributions. A fixed-bandwidth kernel estimator performs well near the mode, where data density is high, but oversmooths in the tails, where observations are sparse, leading to biased estimates. To mitigate this issue, adaptive kernel density estimation adjusts the bandwidth based on the local density at each evaluation point, allowing for finer resolution in sparse regions while maintaining stability in denser areas.
With an adaptive kernel, the bandwidth $h$ is no longer constant but varies with the location of the evaluation point $x_{0}$, denoted as $h(x_{0})$. We take the following density-based adaptation (silverman2018density).
$h(x_{0})=h_{0}\times\left(\frac{\bar{\rho}^{GM}}{\rho(x_{0})}\right)^{\frac{1}{2}}$
where $h_{0}$ is a global baseline bandwidth, $\rho(x_{0})$ denotes the local design density at $x_{0}$, and $\bar{\rho}^{\mathrm{GM}}$ is the geometric mean of $\left\{\rho\left(x_{i}\right)\right\}$ over the sample. In practice, we first obtain a pilot estimate of $\rho(\cdot)$ (e.g., via kernel density estimation) and then plug it into the adaptation rule to produce bandwidths $h(x_{0})$ for the kernel regression. This construction makes $h(x_{0})$ inversely proportional to the local design density: in regions with sparse data, the bandwidth increases, providing broader smoothing across $X$; in denser regions, it decreases, enabling more localized smoothing and preserving finer features. The baseline bandwidth $h_{0}$ is chosen by 10-fold least-squares cross-validation. The sample is randomly partitioned into ten folds. For each fold $k$, we fit the estimator on the remaining nine folds and compute predictions $\widehat{m}^{(-k)}_{h^{(-k)}(x_{i})}(x_{i})$ for the held-out observations $i\in I_{k}$, where $h^{(-k)}(x_{i})$ denotes the adaptive bandwidth computed from the training data for that fold. The cross-validation criterion is
$\mathrm{CV}(h_{0})=\frac{1}{N}\sum_{k=1}^{10}\sum_{i\in I_{k}}\bigl[Y_{i}-\widehat{m}^{(-k)}_{h^{(-k)}(x_{i})}(x_{i})\bigr]^{2},$
and the selected $h_{0}$ minimizes $\mathrm{CV}(h_{0})$, thereby balancing bias and variance.
A common issue in semiparametric regression is boundary bias, where estimates near the edges of the data range tend to be less accurate due to fewer observations. As recommended by fan1995local, HMX (hainmueller2019much) address this problem by incorporating two additional terms, $\eta(X-x_{0})$ and $\beta D(X-x_{0})$, which capture the first partial derivative of $Y$ with respect to $X$ at each $x_{0}$. With a fully moderated version, $\delta Z(X-x_{0})$ is also included for each covariate $Z$. These terms adjust for the local slope of the response surface, helping to reduce bias at the boundaries of the support of $X$, where data are sparse and estimation uncertainty is higher.
Inference
Once we obtain the CME estimates through kernel-weighted least squares, the next step is to estimate the standard errors of the estimated parameters. This can be done using two approaches: (i) analytically, via the robust variance estimator, and (ii) empirically, using bootstrapping.
We can estimate pointwise variance analytically using the robust (sandwich) covariance matrix. The local linear kernel regression estimates $\hat{\alpha}(x_{0})$ and $\hat{\beta}(x_{0})$ are obtained as coefficients from a weighted least squares regression. Consequently, the variance of these estimates can be obtained directly from the robust variance-covariance matrix of the WLS estimator, without requiring additional Taylor expansions.
The confidence interval for $\hat{\theta}(x)=\hat{\alpha}(x)$ is calculated as:
[…]
An alternative approach is bootstrapping. Given a dataset $S$ with $n$ observations, we generate $B$ bootstrap samples $S_{1},S_{2},\dots,S_{B}$ by sampling with replacement from $S$. For each bootstrap sample $S_{b}$, we compute the kernel estimates and extract the CME $\hat{\theta}_{b}=\hat{\alpha}_{b}$. The standard deviation of these bootstrap estimates approximates the standard error:
$SE(\hat{\theta})\approx\sqrt{\frac{1}{B-1}\sum_{b=1}^{B}(\hat{\theta}_{b}-\bar{\theta}_{B})^{2}}$
where $\bar{\theta}_{B}$ is the mean of the bootstrap estimates.
Furthermore, we extend the bootstrap approach to construct uniform confidence intervals, similar to those used with the linear interaction model. Again, we implement the bootstrap-based sup-$t$ band method, ensuring valid inference across the entire range of $X$.
We apply the kernel estimator to Examples 4 and 7. The CME estimates from both examples are shown in Figure 6. Using data from huddy2015expressive, the kernel estimator yields CME estimates that are nearly identical to those from the linear interaction model, suggesting that Assumption 7 is reasonable in this setting.
1# R code excerpt
2
3library(interflex)
4### Huddy et. al 2015
5Y="totangry" #Angre
6D="threat" #Threat
7X="pidentity" #Partisan Identity
8Z<-c("issuestr2", "pidstr2", "knowledge" , "educ" , "male" , "age10" )
9
10out.kernel1 <- interflex(estimator = "kernel",
11 Y=Y,D=D,X=X,data=d,
12 vartype = "bootstrap", ## uniform CI
13 full.moderate = TRUE,
14 na.rm = TRUE)
15plot(out.kernel1)
16
17### Clark and Golder 2006
18Y="enep1" #Effective number of electoral parties
19D="proximity1" #Temporal proximity of presidential elections
20X="enpres" #Number of presidential candidates"
21Z=c("uppertier" ,"logmag" , "uppertier_eneg" ,"eneg", "logmag_eneg")
22
23out.kernel2 <- interflex(estimator = "kernel",
24 Y=Y,D=D,X=X,data=d,
25 vartype = "bootstrap", ## uniform CI
26 full.moderate = TRUE, na.rm = TRUE)
27plot(out.kernel2)
Figure 6: Applying the Kernel Estimator
Notes: The above figures replicate Figure 2A from huddy2015expressive (a) and Figure 2 from clark2006rehabilitating (b). In figure (b), treatment is presidential elections, outcome is the effective number of electoral parties. (a) Plot of the raw outcome data in each bin of the moderator. In each figure, the black dashed line represents the CME estimates based on the kernel estimator; shaded area and dotted lines represent pointwise and uniform confidence intervals, respectively, via nonparametric bootstrapping.
In contrast, using data from clark2006rehabilitating, the CME estimates differ substantially from those based on the linear interaction model. The first thing we notice is the absence of variation in the treatment variable when the moderator value is zero, indicating that the treatment effect is unidentifiable at this point due to lack of common support. This is not captured by the linear interaction model, but with kernel estimator, the confidence intervals blow up as the moderator approaches zero, given that there is no variation in the treatment variable at this point.
The kernel estimator also picks up on nonlinearity in the CME masked by the linear interaction model. Contrary to the anticipated positive interaction suggested by the study, the CME estimates are insignificant when $X$ is close to 0, shifts to a negative and statistically significant as the moderator ranges from approximately $[0.5,2]$, and then reverts nearly to indistinguishable from zero when $X$ is bigger than 2.
It is worth noting that in both examples, the patterns revealed by the kernel estimator are broadly consistent with the findings from the binning estimator, which we recommend as a diagnostic tool.
2.5 Summary
This chapter examines classic methods for estimating and visualizing the CME. We begin with the widely used linear interaction model, discussing its variance estimator and hypothesis testing procedures. To address the issue of multiple comparisons, we introduce uniform confidence intervals constructed via bootstrapping.
Next, we highlight two key challenges associated with the linear interaction model: lack of common support and model misspecification. We propose diagnostic tools, such as raw data plots and a binning estimator, to detect these issues and discuss strategies to mitigate them, including a design phase to improve overlap in observational studies.
Finally, we introduce a semiparametric kernel estimator to relax the functional form assumptions of linear interaction models. Building on HMX (hainmueller2019much), we propose two key improvements: a fully moderated specification and an adaptive kernel for more accurate estimation across different data regions. We also implement uniform confidence intervals for the kernel estimator to ensure valid inference.
Below, we summarize the challenges associated with the linear interaction model, along with diagnostic methods and potential solutions.
  • Challenge: Multiple comparisons
    • Solution: Uniform confidence intervals
  • Challenge: Lack of common support
    • Diagnostics tools: Raw outcome plot, propensity score plot
    • Solution: The design phase
  • Challenge: Rigid functional form
    • Diagnostics tools: Raw outcome plot, the binning estimator
    • Solution: Fully moderated model, the kernel estimator, AIPW and extensions (Chapter 3), double/debiased machine learning (Chapter 4)
Chapter 3. AIPW and Extensions
So far, our analysis has primarily focused on modeling the conditional expectation function of the potential outcome, $\mathbb{E}[Y(d)\mid X,Z]$. For example, the linear interaction model posits that
$\mathbb{E}[Y(d)\mid X,Z]=\beta_{0}+\beta_{1}d+\beta_{2}X+\beta_{3}(d\cdot X)+\beta_{4}^{\prime}Z,$
while a semiparametric model assumes
$\mathbb{E}[Y(d)\mid X,Z]=f(X)+g(X)\,d+Z^{\prime}\,\gamma(X).$
Combining with the unconfoundedness assumption, we can use the OLS or the local linear kernel regression to identify the parameter(s) of interest, such as the $\beta_{3}$ in the linear interaction model or the $g(X)$ in the semiparametric model.
However, these approaches have important limitations. First, they do not explicitly account for the treatment assignment process. In the binary-treatment setting, for instance, one might model the propensity score $\mathbb{E}[D\mid X,Z]=\Pr[D=1\mid X,Z]$ to improve robustness when estimating the CME. Second, although we allow for nonlinearities in $X$, we have largely overlooked potential nonlinearities and high-dimensional interactions among covariates $Z$. The kernel estimator permits a nonlinear relationship between $Z$ and $Y$ only through the varying coefficients $\gamma(X)$; it does not flexibly capture the nonlinear structure of $Z$ itself or accommodate higher-order interactions among its components.
In this chapter, we address these limitations using the augmented inverse probability weighting (AIPW) approach and its extensions. We begin with the binary-treatment case. After revisiting the IPW estimator from Chapter 1—now allowing for a continuous moderator—we introduce the AIPW framework. The IPW estimator identifies the CME solely through the propensity score, without specifying an outcome model. The AIPW estimator further combines an outcome model with the propensity-score model, achieving double robustness: it remains consistent provided that either the outcome model or the propensity-score model is correctly specified. This AIPW framework for CME estimation has been systematically discussed by knaus2022double and kennedy2023towards.
When the treatment $D$ is continuous, the CME can be estimated analogously. We first “partial out” the covariates from both the outcome and the treatment, yielding residualized (“partialled out”) versions of $Y$ and $D$. The CME can then be identified from the relationship between these adjusted variables. This residual-based estimator was developed by nie2021quasi.
The AIPW and “partial-out” frameworks lay the foundation for the double/debiased machine learning (DML) framework introduced in the next chapter. The key difference here is that we use regularized linear models (e.g., Lasso) rather than more advanced machine-learning algorithms to estimate the outcome and propensity-score models. We also use a simple bootstrap for inference in this chapter, rather than the cross-fitting and Gaussian multiplier bootstrap common in the DML literature. We view this chapter as a bridge between the quasi-linear models introduced earlier and the more advanced DML methods introduced in the next chapter.
3.1 Outcome, IPW and AIPW Signals for the CME
This section formalizes how to estimate the CME for a binary treatment using three complementary identification strategies—outcome modeling, inverse probability weighting (IPW), and augmented IPW (AIPW). We establish identification for each approach and discuss their robustness and variance trade-offs. Each method maps the data for each observation into a corresponding signal. Under the respective model assumptions, the conditional expectation of these signals given $X=x$ equals the CME, $\theta(x)$.
In Chapter 1, we have defined the conditional mean outcome by treatment condition $\mu^{d}(v)=\mathbb{E}[Y_{i}\mid D_{i}=d,V_{i}=v]$ for $d\in\{0,1\}$, and the propensity score $\pi(v)=\Pr(D_{i}=1\mid V_{i}=v)$. We showed that, when $D$ is binary and all covariates in $V$ are discrete, under unconfoundedness and overlap assumptions, the stratified difference-in-means identifies the CME:
$\sum_{v}\left\{\left(\mu^{1}(v)-\mu^{0}(v)\right)\cdot\Pr(V=v\mid X=x)\right\}=\theta(x),$
which naturally extends to cases where $Z$ includes continuous variables by replacing the sum with an integral:
$\int\left(\mu^{1}(v)-\mu^{0}(v)\right)\cdot f_{V\mid X}(v\mid x)\,dv=\theta(x),$
where $f_{V\mid X}(v\mid x)$ denotes the conditional density of $V$ given $X=x$. We can therefore use expectation notation to cover both cases:
$\mathbb{E}\!\left[\mu^{1}(V_{i})-\mu^{0}(V_{i})\mid X_{i}=x\right]=\theta(x).$ (4)
Moreover, in Chapter 1 we also showed that the IPW estimator identifies the CME:
$\mathbb{E}\!\left[\left.\frac{D_{i}Y_{i}}{\pi(V_{i})}-\frac{(1-D_{i})Y_{i}}{1-\pi(V_{i})}\right|\,X_{i}=x\right]=\theta(x).$ (5)
Both expectations are taken over the conditional distribution of $V$ given $X$.
Building on the strengths of both outcome modeling and IPW, robins1994estimation introduced the AIPW estimator, which combines the outcome model and the propensity-score model to improve robustness and efficiency in estimating the CME. The AIPW estimator for the CME can be constructed as
$\mathbb{E}\!\left[\mu^{1}(V_{i})-\mu^{0}(V_{i})\mid X_{i}=x\right]\;+$
$\mathbb{E}\!\left[\frac{D_{i}}{\pi(V_{i})}\bigl(Y_{i}-\mu^{1}(V_{i})\bigr)-\frac{1-D_{i}}{1-\pi(V_{i})}\bigl(Y_{i}-\mu^{0}(V_{i})\bigr)\,\Big|\,X_{i}=x\right]\;=\;\theta(x).$ (6)
The proof in the appendix shows that the equality holds (i.e., AIPW is consistent) when either the outcome models $\mu^{1}(V_{i})$ and $\mu^{0}(V_{i})$ or the propensity-score model $\pi(V_{i})$ is correctly specified, but not necessarily both. This property is known as double robustness.
Asymptotically, beyond double robustness, AIPW exhibits lower variance than IPW when both the outcome and propensity-score models are correctly specified. robins1994estimation show that AIPW attains the smallest asymptotic variance within the class of inverse probability–weighted estimators. However, when only the outcome model(s) are correctly specified, AIPW remains consistent but can have higher variance than a purely outcome-based estimator.
In finite samples, AIPW does not necessarily dominate the outcome model or IPW. li2019addressing demonstrate via simulations that its finite-sample performance depends heavily on overlap in propensity scores. When overlap is poor—that is, many estimated propensity scores $\hat{\pi}(V_{i})$ lie near 0 or 1, producing very large weights—AIPW can perform worse than an outcome-based estimator.
Motivated by this finite-sample concern, we adopt propensity-score clipping to mitigate instability from extreme weights. IPW and AIPW estimators can be sensitive when the estimated propensity score $\hat{\pi}(V_{i})$ is close to 0 or 1. To address this, we replace $\hat{\pi}(V_{i})$ with
$\tilde{\pi}(V_{i})=\min\!\left\{\max\!\bigl(\hat{\pi}(V_{i}),\alpha\bigr),1-\alpha\right\},$
for a small threshold $\alpha>0$. Clipping reduces variance and improves finite-sample stability, at the cost of a small bias when scores are modified (cole2008constructing). In the remainder of this chapter, we set $\alpha=10^{-2}$ in our implementations.
Signals for the CME.
We now introduce signals, defined as per-observation transformations whose conditional expectations equal the CME. We define outcome, IPW, and AIPW signals as
$\Lambda_{i}^{(\mathrm{outcome})}$ $=\mu^{1}(V_{i})-\mu^{0}(V_{i}),$
$\Lambda_{i}^{(\mathrm{ipw})}$ $=\frac{D_{i}}{\pi(V_{i})}\,Y_{i}-\frac{1-D_{i}}{1-\pi(V_{i})}\,Y_{i},$
$\Lambda_{i}^{(\mathrm{aipw})}$ $=\mu^{1}(V_{i})-\mu^{0}(V_{i})+\frac{D_{i}}{\pi(V_{i})}\bigl(Y_{i}-\mu^{1}(V_{i})\bigr)-\frac{1-D_{i}}{1-\pi(V_{i})}\bigl(Y_{i}-\mu^{0}(V_{i})\bigr).$
Therefore, each of Equations (4)–(6) implies an identification result for a corresponding signal under the correctly specified model: [^4] By “correctly specified,” we mean that the working model coincides with the true data generating process on the support of $V$: for all $v$, $\mu^{d}(v)=\mathbb{E}[Y(d)\mid V=v]$ for $d\in\{0,1\}$ and/or $\pi(v)=\Pr(D=1\mid V=v)$, with positivity $0<\pi(v)<1$. For estimated models $\hat{\mu}^{d},\hat{\pi}$, this can be read as large-sample convergence to the truth.
Equation (4) $\Rightarrow\quad\mathbb{E}\!\left[\Lambda_{i}^{(\mathrm{outcome})}\mid X_{i}=x\right]=\theta(x)\quad\text{if }\mu=(\mu^{1},\mu^{0})\text{ is correctly specified},$
Equation (5) $\Rightarrow\quad\mathbb{E}\!\left[\Lambda_{i}^{(\mathrm{ipw})}\mid X_{i}=x\right]=\theta(x)\quad\text{if }\pi\text{ is correctly specified},$
Equation (6) $\Rightarrow\quad\mathbb{E}\!\left[\Lambda_{i}^{(\mathrm{aipw})}\mid X_{i}=x\right]=\theta(x)\quad\text{if either }\mu=(\mu^{1},\mu^{0})\text{ or }\pi\text{ is correctly specified}.$
In other words, taking the conditional expectation of any of the outcome, IPW, or AIPW signals given $X_{i}=x$, under their respective modeling assumptions, identifies the CME.
Signals for the CME on the treated.
Although we focus primarily on the CME (a special case of the conditional average treatment effect, CATE), the same logic extends to other estimands. A commonly used estimand is the conditional average treatment effect on the treated (CATT), $\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid D_{i}=1,V_{i}=v]$. Analogously, restricting attention to heterogeneity in the moderator yields the conditional marginal effect on the treated (CMET),
$\mathrm{CMET}(x)=\mathbb{E}\!\left[Y_{i}(1)-Y_{i}(0)\mid D_{i}=1,X_{i}=x\right].$
Outcome, IPW, and AIPW signals for the CMET are
$\Xi_{i}^{(\mathrm{outcome})}$ $=\frac{Y_{i}-\mu^{0}(V_{i})}{\Pr(D_{i}=1\mid X_{i})}\,D_{i},$
$\Xi_{i}^{(\mathrm{ipw})}$ $=\frac{Y_{i}\bigl(D_{i}-\pi(V_{i})\bigr)}{\Pr(D_{i}=1\mid X_{i})\bigl(1-\pi(V_{i})\bigr)},$
$\Xi_{i}^{(\mathrm{aipw})}$ $=\frac{1}{\Pr(D_{i}=1\mid X_{i})}\bigl(Y_{i}-\mu^{0}(V_{i})\bigr)\!\left(D_{i}-\frac{\pi(V_{i})(1-D_{i})}{1-\pi(V_{i})}\right).$
Here, $\Pr(D_{i}=1\mid X_{i})$ is the probability of treatment given the moderator; it can be estimated parametrically or nonparametrically. [^5] It is worth noting that $\Pr(D_{i}=1\mid X_{i})$ is the marginal propensity score, which is conditioned only on $X$, and needs to be estimated separately from the full propensity score $\pi\left(V_{i}\right)$, which is conditioned on $V=(X,Z)$. Under unconfoundedness and overlap, we establish in the appendix that under correctly specified models:
$\mathbb{E}\!\left[\Xi_{i}^{(\mathrm{outcome})}\mid X_{i}=x\right]=\mathbb{E}\!\left[\Xi_{i}^{(\mathrm{ipw})}\mid X_{i}=x\right]=\mathbb{E}\!\left[\Xi_{i}^{(\mathrm{aipw})}\mid X_{i}=x\right]=\mathrm{CMET}(x).$
The AIPW signal possesses this property when either the outcome model $\mu^{0}(\cdot)$ or the propensity-score model $\pi(\cdot)$ is correctly specified. Notably, estimation of $\mathrm{CMET}(x)$ does not require modeling the treated potential-outcome function $\mu^{1}(\cdot)$.
3.2 Estimation Strategies
This subsection is slightly more technical than the preceding material. Readers new to weighting and doubly robust methods can first skim the three-step “recipe” in estimating CME using AIPW and the worked examples, then return to details as needed; those already familiar with IPW and AIPW can treat this as a compact reference.
Implementing the outcome, IPW, and AIPW estimators for the CME proceeds in three steps. First, estimate the propensity score $\Pr(D=1\mid X,Z)$ and the outcome models $\mathbb{E}[Y\mid D=0,X,Z]$ and $\mathbb{E}[Y\mid D=1,X,Z]$. We refer to these as nuisance parameters: they are not of direct substantive interest but are essential for identifying and estimating the CME.
[…]
Finally, estimate the CME as a function of $X$ using the constructed signals. When $X$ is discrete, compute subgroup means by averaging the signals within each level of $X$. When $X$ is continuous, regress the signals $\{\hat{\Lambda}_{i}\}$ on a flexible representation of $X$ (e.g., a B-spline basis expansion) to obtain a smooth estimate of the CME. Alternatively, we can use the semiparametric kernel regression, as introduced in Chapter 2, to estimate this relationship.
Steps 1 and 2: Constructing Signals
We begin by discussing how to construct signals from data using outcome modeling, IPW, and AIPW.
Signals from outcome modeling.
With the outcome-modeling approach, we can fit two separate linear regressions for data with $D_{i}=1$ and $D_{i}=0$. Specifically, we regress $Y_{i}$ on $(1,V_{i})$ separately within each treatment group, yielding the following estimates:
$\hat{\mu}^{0}(V_{i})=\hat{\beta}_{0}^{0}+V_{i}\hat{\beta}_{V}^{0}\;,\qquad\hat{\mu}^{1}(V_{i})=\hat{\beta}_{0}^{1}+V_{i}\hat{\beta}_{V}^{1}\;,$
where $(\hat{\beta}_{0}^{0},\hat{\beta}_{V}^{0\prime})$ and $(\hat{\beta}_{0}^{1},\hat{\beta}_{V}^{1\prime})$ are the OLS coefficients from these two regressions. For each observation $i$, we then construct the signal from the outcome model:
$\hat{\Lambda}_{i}^{(\mathrm{outcome})}=\hat{\mu}^{1}(V_{i})-\hat{\mu}^{0}(V_{i})=(\hat{\beta}_{0}^{1}-\hat{\beta}_{0}^{0})+V_{i}(\hat{\beta}_{V}^{1}-\hat{\beta}_{V}^{0})\,$,
which represents the predicted treatment effect for observation $i$.
IPW signals.
For the IPW estimator, we first fit a propensity-score model using a logistic regression of $D$ on $V$. For each observation $i$, the estimated propensity score is
$\hat{\pi}(V_{i})=\widehat{\Pr}\bigl(D_{i}=1\mid V_{i}\bigr)=\frac{\exp\bigl(\hat{\gamma}_{0}+V_{i}\hat{\gamma}_{V}\bigr)}{1+\exp\bigl(\hat{\gamma}_{0}+V_{i}\hat{\gamma}_{V}\bigr)}\,$,
where $(\hat{\gamma}_{0},\hat{\gamma}_{V}^{\prime})$ are coefficients from the logistic regression. [^6] As discussed previously, we can clip the estimated propensity score to $\alpha,1-\alpha]$ with $\alpha=0.01$ to improve the stability of the estimator.
For each observation $i$, we then construct the IPW signals, which are the IPW-adjusted outcomes:
$\hat{\Lambda}_{i}^{(\mathrm{ipw})}=\frac{D_{i}}{\hat{\pi}(V_{i})}\,Y_{i}-\frac{1-D_{i}}{1-\hat{\pi}(V_{i})}\,Y_{i}.$
This adjustment reweights the treated and control outcomes by $\frac{1}{\hat{\pi}(V_{i})}$ and $\frac{1}{1-\hat{\pi}(V_{i})}$, respectively:
$\hat{\Lambda}_{i}^{(\mathrm{ipw})}=\frac{Y_{i}}{\hat{\pi}(V_{i})}\,\quad\text{if }D_{i}=1,\qquad\hat{\Lambda}_{i}^{(\mathrm{ipw})}=-\frac{Y_{i}}{1-\hat{\pi}(V_{i})}\,\quad\text{if }D_{i}=0.$
This reweighting creates a pseudo-population in which covariates $V_{i}$ are (asymptotically) balanced between treated and control groups via the propensity score, ensuring that a simple difference in means in this pseudo-population can recover the treatment effect. Specifically, among treated units, those with smaller estimated propensity scores receive more weight, while among control units, those with higher estimated propensity scores receive more weight. This adjustment corrects for selection bias induced by $V$ by aligning covariate distributions across treatment groups.
AIPW signals.
The AIPW estimator combines elements of both outcome modeling and IPW. We first fit two linear regression models, $\hat{\mu}^{1}(V_{i})$ and $\hat{\mu}^{0}(V_{i})$, as well as the propensity-score model $\hat{\pi}(V_{i})$. We then construct the signal as follows:
$\hat{\Lambda}_{i}^{(\mathrm{aipw})}=\hat{\mu}^{1}(V_{i})-\hat{\mu}^{0}(V_{i})+\frac{D_{i}}{\hat{\pi}(V_{i})}\bigl(Y_{i}-\hat{\mu}^{1}(V_{i})\bigr)-\frac{1-D_{i}}{1-\hat{\pi}(V_{i})}\bigl(Y_{i}-\hat{\mu}^{0}(V_{i})\bigr).$
Note that
$\hat{\Lambda}_{i}^{(\mathrm{aipw})}$ $=\hat{\mu}^{1}(V_{i})-\hat{\mu}^{0}(V_{i})+\frac{1}{\hat{\pi}(V_{i})}\bigl(Y_{i}-\hat{\mu}^{1}(V_{i})\bigr)$ $\text{if }D_{i}=1,$
$\hat{\Lambda}_{i}^{(\mathrm{aipw})}$ $=\hat{\mu}^{1}(V_{i})-\hat{\mu}^{0}(V_{i})-\frac{1}{1-\hat{\pi}(V_{i})}\bigl(Y_{i}-\hat{\mu}^{0}(V_{i})\bigr)$ $\text{if }D_{i}=0.$
For both treated and control units, the first term, $\hat{\mu}^{1}(V_{i})-\hat{\mu}^{0}(V_{i})$, represents the predicted treatment effect for unit $i$ based on the outcome model. The second term, $\frac{1}{\hat{\pi}(V_{i})}\bigl(Y_{i}-\hat{\mu}^{1}(V_{i})\bigr)$ for treated units and $\frac{1}{1-\hat{\pi}(V_{i})}\bigl(Y_{i}-\hat{\mu}^{0}(V_{i})\bigr)$ for control units, adjusts for discrepancies between observed outcomes and their outcome modeling-based predictions by reweighting residuals from the outcome model. Similar to the IPW estimator, treated units with smaller estimated propensity scores receive more weight, while control units with larger estimated propensity scores are weighted more heavily. However, instead of reweighting the raw outcome $Y_{i}$, AIPW reweights the residuals $\bigl(Y_{i}-\hat{\mu}^{D_{i}}(V_{i})\bigr)$, which helps reduce sensitivity to misspecification in the outcome model.
It is worth noting that fixed effects can be included in outcome models to control for unobserved unit-invariant or time-invariant confounders; however, we do not recommend incorporating fixed effects into the propensity-score estimation. Since propensity-score models typically use nonlinear specifications (e.g., logit or probit), including a large number of fixed effects can lead to the incidental parameters problem (neyman1948consistent). This problem arises when the number of fixed-effect parameters increases with the number of observations, resulting in biased and inconsistent maximum likelihood estimates when the number of observations per group is small.
Next, we use a simulated example to illustrate the advantages of IPW and AIPW over a purely outcome-modeling approach. In this example, the CME is nonlinear, but the researcher misspecifies the outcome model as a linear interaction model. Fortunately, the researcher correctly specifies the propensity-score model. This scenario highlights the limitations of relying solely on outcome modeling, particularly when using rigid parametric models. Later, we introduce basis expansions to relax parametric assumptions and enhance estimation flexibility.
Example 8 (A Simulated Sample with a Misspecified Outcome Model)
The sample consists of 1,000 observations, with a moderator $X_{i}$ uniformly distributed on [2,2][-2,2] and two covariates $Z_{i1}$ and $Z_{i2}$ uniformly distributed on $[0,1]$. Let $V_{i}=[X_{i},Z_{i1},Z_{i2}]$ be the set of all confounders. Treatment is assigned with probability
$\Pr(D_{i}=1)=\frac{\exp(0.5X_{i}+0.5Z_{i1})}{1+\exp(0.5X_{i}+0.5Z_{i1})}.$
The outcome $Y$ is generated via
$Y_{i}=1+X_{i}+D_{i}-X_{i}^{2}\,D_{i}+Z_{i1}+\epsilon_{i},$
where $\epsilon_{i}\sim\mathcal{N}(0,1)$. Unconfoundedness holds: $D_{i}\perp\!\!\!\perp(Y_{i}(0),Y_{i}(1))\mid V_{i}$. The CME is nonlinear in $x$, i.e., $\theta(x)=1-x^{2}$. Covariate $Z_{i2}$ is redundant since it does not directly affect $D_{i}$ or $Y_{i}$. However, the researcher correctly specifies the propensity-score model but misspecifies the outcome model as
$Y_{i}=1+X_{i}+D_{i}+X_{i}D_{i}+Z_{i1}+\epsilon_{i}.$
Figure 7(d)(a) plots the raw outcome $Y_{i}$ against the moderator $X_{i}$, using dark blue for treated units ($D_{i}=1$) and light blue for control units ($D_{i}=0$). In the treated group, the relationship between $X_{i}$ and $Y_{i}$ is nonlinear.
Figure 7(d)(b) plots the signals from the outcome model, $\hat{\Lambda}_{i}^{(\mathrm{outcome})}$. Because the model is misspecified as linear, the signals—representing predicted treatment effects—are linear in $X$. Here, we fit a B-spline regression to these signals to obtain a smooth estimate of the CME with respect to $X$ (the orange line). These misspecified signals lead to misleading CME estimates, which are also linear in $X$ and differ from the true CME shown in red.
Figure 7: Constructed Signals
(a)   Raw Outcomes
(b)   Outcome Model Signals & $\hat{\theta}(x)$
(c)   IPW Signals & $\hat{\theta}(x)$
(d)   AIPW Signals & $\hat{\theta}(x)$
Note: In (a), the dark blue and light blue dots represent outcome values from the treatment and control groups, respectively; the dark blue and light blue lines represent the LOESS fits for the two groups. In (b)–(d), the dark blue and light blue dots represent signals of the observations in the treatment and control groups, respectively; the red line represents the true CME, while the orange line represents the CME estimates based on the signals after smoothing.
Figure 7(d)(c) plots the signals from the IPW estimator, $\hat{\Lambda}_{i}^{(\mathrm{ipw})}$. Compared to the corresponding raw outcomes, treated observations (dark blue) with small $X$ values are relatively upweighted, as are control observations (light blue) with large $X$ values. This occurs because $D$ (and the estimated propensity score) is positively correlated with $X$. Since the propensity-score model is correctly specified, this adjustment mitigates selection bias and helps reveal the true CME after smoothing.
In Figure 7(d)(d), we plot the AIPW signals, $\hat{\Lambda}_{i}^{(\mathrm{aipw})}$. They behave similarly to the IPW signals and successfully recover the true CME after smoothing even when the outcome model is misspecified, illustrating the doubly robust property. Notably, their spread is much smaller than that of IPW because the misspecified outcome model still retains some predictive power for the outcome. Moreover, they exhibit fewer extreme values—particularly for control units—than the IPW signals.
Step 3: Smoothing
So far, we have glossed over the smoothing step, which reduces the signals to a one-dimensional CME along $X$; we now formally discuss it. Once we obtain the signals $\hat{\Lambda}_{i}$ from outcome modeling, IPW, or AIPW, the final step is to approximate the CME by fitting a flexible function $f(X)$ that represents the conditional mean of these signals given $X$. When the moderator $X$ is discrete, this reduces to averaging the signals within each group defined by $X$. For a continuous moderator, we consider two approaches for estimating $f(X)$: one using kernel regression, which parallels the kernel estimator discussed in Chapter 2, and another using B-spline regression.
Kernel Regression
One can employ the kernel-weighted local polynomial smoother (LOESS). For each evaluation point $x_{0}$, LOESS constructs a locally weighted regression of $\hat{\Lambda}$ on $X$ within a neighborhood around $x_{0}$. Users can specify a span parameter, $\alpha$, which determines the size of this local neighborhood as a fraction of the overall range of $X$. Let
$W_{i}(x_{0};\alpha)\;=\;K\!\left(\frac{|X_{i}-x_{0}|}{\alpha\,\mathrm{range}(X)}\right),$
where $K(\cdot)$ is a kernel weight function (often the tricube weight in standard LOESS). One then solves a weighted least squares problem:
$\bigl(\hat{\beta}_{0}(x_{0}),\hat{\beta}_{1}(x_{0})\bigr)\;=\;\operatorname*{arg\,min}_{\beta_{0},\,\beta_{1}}\sum_{i=1}^{n}\Bigl[\hat{\Lambda}_{i}-\beta_{0}-\beta_{1}\bigl(X_{i}-x_{0}\bigr)\Bigr]^{2}\,W_{i}(x_{0};\alpha),$
where $\hat{\Lambda}_{i}$ denotes the signals. The coefficient $\hat{\beta}_{0}(x_{0})$ represents the local intercept at $x_{0}$, while $\hat{\beta}_{1}(x_{0})$ captures the local slope in the neighborhood of $x_{0}$. Therefore, the fitted function at $x_{0}$ is
$\hat{f}(x_{0})\;=\;\hat{\beta}_{0}(x_{0}).$
This approach ensures that data points closer to $x_{0}$ receive higher weight, while those farther away have less influence. If a higher-order polynomial fit is desired (e.g., quadratic), include additional terms such as $\beta_{2}(X_{i}-x_{0})^{2}$, $\beta_{3}(X_{i}-x_{0})^{3}$, etc., with the kernel weights $W_{i}(x_{0};\alpha)$ unchanged.
To select an optimal $\alpha$, we employ $K$-fold cross-validation. For each candidate $\alpha$ in a predefined grid, compute the mean squared error (MSE) over the folds:
$\operatorname{MSE}(\alpha)\;=\;\frac{1}{K}\sum_{k=1}^{K}\frac{1}{n_{k}}\sum_{i\in\mathrm{Fold}_{k}}\Bigl(\hat{\Lambda}_{i}-\hat{f}_{-k}(X_{i};\alpha)\Bigr)^{2},$
where $\hat{f}_{-k}$ denotes the LOESS fit excluding observations in the $k$-th fold, and $n_{k}$ is the number of observations in that fold. We then choose
$\alpha^{*}\;=\;\operatorname*{arg\,min}_{\alpha}\operatorname{MSE}(\alpha),$
and refit LOESS on the entire dataset using $\alpha^{*}$. Predicting over the evaluation grid yields the final estimate $\hat{f}(x)$, which in turn provides the CME at each $x$.
B-Spline Regression
A B-spline is a piecewise polynomial function that ensures smoothness at the junctions of polynomial segments (the knots). In our setting, for the signals $\hat{\Lambda}_{i}$, we fit a linear model of the form
$\hat{\Lambda}_{i}=\beta_{0}+\sum_{j=1}^{K}\beta_{j}\,B_{j}(X_{i})+\epsilon_{i},$
where $B_{j}(\cdot)$ is the $j$-th B-spline basis function of chosen degree (spline_degree) and dimension (spline_df). By selecting an adequate number of knots and a suitable polynomial degree, [^7] In implementation, we set spline_degree to 3 (cubic spline) and spline_df to 6; users may choose spline_degree and spline_df based on other criteria such as AIC or BIC. the B-spline expansions $\{B_{j}(X)\}_{j=1}^{K}$ capture nonlinearities in the signal as a function of $X$. After estimating $\{\beta_{j}\}$ via least squares, we predict $\hat{f}(x)$ over an evaluation grid to obtain $\theta(x)$. Formally,
$\hat{f}(x)\;=\;\hat{\beta}_{0}+\sum_{j=1}^{K}\hat{\beta}_{j}\,B_{j}(x).$
Figure 8: Kernel vs. B-Spline Regressions for Smoothing
(a)   Kernel Regression
(b)   B-Spline Regression
Note: Dark blue and light blue dots represent AIPW signals for the treatment and control groups, respectively. The red line represents the true CME, while the orange line represents the estimated CME.
Both kernel and B-spline regression are flexible modeling strategies for capturing nonlinear relationships, but they differ in how smoothness is regulated and how tuning parameters (knot configuration vs. $\alpha$-span) are selected. B-spline regression employs a polynomial basis with continuity constraints at knot boundaries, whereas kernel/LOESS methods adaptively weight data within localized neighborhoods to achieve smooth fits. In either case, the estimated function $\hat{f}(x)$ captures the shape of $\hat{\Lambda}$ as a function of $X$, delivering pointwise CME estimates across the range of the moderator. Figure 8(b) compares these two approaches (orange lines) in practice, showing that they generally produce similar fits when applied to $\hat{\Lambda}_{i}^{(\mathrm{aipw})}$. In all subsequent CME estimations in this chapter, we continue to use B-spline regression to obtain the CME from the constructed signals due to its computational efficiency.
3.3 Basis Expansion and Variable Selection
One way to increase the flexibility of the outcome and propensity-score models is to expand the set of regressors $V=(X,Z)$ beyond their raw forms. Specifically, each covariate—whether the moderator $X$ or an additional covariate $Z_{j}\in Z$—can be replaced with a set of basis functions that capture nonlinearities, and interactions among them can also be incorporated. Common approaches include polynomial expansions (e.g., second degree or higher) and B-spline expansions, which partition the support of a variable into segments while enforcing smoothness across boundaries. By enriching the model space, the estimated relationships can better adapt to complex patterns that might otherwise be inadequately captured by purely linear terms.
Concretely, suppose $V\in\mathbb{R}^{1+p}$, where $p$ is the dimension of the covariates. For each variable $L$ in $V$, including $X$, we can generate a set of transformed variables $\{\phi_{k}(L)\}$, where each function represents a polynomial or B-spline basis expansion. Interaction terms can be constructed by multiplying these expansions for $X$ with those for $Z$, allowing the model to capture effect modification across covariates. We denote the full set of expansions by $\psi(V)$.
Once expanded, these regressors can be incorporated into both the outcome model $\mathbb{E}[Y\mid D,X,Z]$ and the propensity-score model $\mathbb{E}[D\mid X,Z]$. In the outcome model, basis expansions enable more flexible relationships between $X$, $Z$, and $Y$. If the true relationship is nonlinear, involves threshold effects, or exhibits heterogeneous slopes across different regions of $X$ or $Z$, the expanded basis helps approximate these complexities. In the propensity-score model, these expansions allow the probability of treatment to vary nonlinearly with the covariates. With a sufficiently rich expansion—and appropriate regularization, such as Lasso—these flexible specifications can reduce bias by better approximating the unknown data-generating process.
A practical rule of thumb is to prefer flexible yet parsimonious bases. For polynomial expansions, prefer low-degree polynomials (quadratic or cubic) and avoid high-degree polynomials. For B-splines, use cubic (degree 3) or natural splines for better boundary behavior, and employ a small number of interior knots (e.g., 3–6) placed at empirical quantiles (de1978practical; hastie2009elements). In principle, these hyperparameters can be selected by cross-validation or information criteria. Because we pair these choices with Lasso regularization tuned by cross-validation in the next section, we do not additionally cross-validate them here, to conserve computational resources. In the remaining chapters and simulations, we use a B-spline expansion with spline_degree = 3 and spline_df = 6 (placing three internal knots at 25th, 50th, and 75th percentiles). We also include interaction terms between the expanded basis of $X$ and the expanded basis of each $Z$, as well as interactions between the expanded bases of different $Z$ variables.
We illustrate the benefit of basis expansions using another simulated example. We apply outcome modeling, IPW, and AIPW, showing results for each, with and without basis expansion.
Example 9 (A Simulated Example with Complex Nonlinear Relationships)
We simulate $V=(X_{i},Z_{i1},Z_{i2},Z_{i3})$ with $X_{i}\sim\mathrm{Unif}[-2,2]$ and each $Z_{ij}\sim\mathrm{Unif}[0,1]$ for $j=1,2,3$. The treatment is generated by
$D_{i}\sim\mathrm{Bernoulli}\Bigl(\mathrm{logit}^{-1}\bigl(-1+0.5Z_{i1}+|X_{i}|-X_{i}\cdot Z_{i2}+Z_{i3}^{2}\bigr)\Bigr),$
and the outcome is generated by
$Y_{i}=1+X_{i}+D_{i}-D_{i}X_{i}^{2}+Z_{i1}^{2}+\sin(Z_{i2})+0.5\,Z_{i1}\exp(Z_{i3})+\varepsilon_{i},$
where $\varepsilon_{i}\sim\mathcal{N}(0,1)$. Under this DGP, the true CME is $\theta(x)=1-x^{2}$. Unconfoundedness holds: $D_{i}\perp\!\!\!\perp(Y_{i}(0),Y_{i}(1))\mid V_{i}$. However, the researcher is unaware of the functional form of either the propensity-score model or the outcome model.
Figure 9: With and Without Basis Expansion
(a) Outcome Modeling: without and with basis expansion
(b) IPW: without and with basis expansion
(c) AIPW: without and with basis expansion
Note: Dark blue and light blue dots represent signals for the treatment and control groups, respectively. The red line represents the true CME, while the orange line represents the estimated CME. In each subfigure, the left and right panels show results without and with basis expansion, respectively.
Figure 9 shows the results. The left panel in each subfigure presents the signals and CME estimates without basis expansion, and the right panel includes basis expansions on both $X$ and $Z$. The left panels in Figure 9 demonstrate that (i) the linear outcome model fails to capture the nonlinearity; (ii) the IPW estimator also misses the true shape and occasionally produces large outliers in the estimated signals; and (iii) the AIPW estimator performs somewhat better even with a linear outcome model, likely because the two models compensate for each other’s deficiencies.
The right panels of Figure 9 show that both the outcome and AIPW estimators capture the inverse-$U$ shape of the CME, $\theta(x)=1-x^{2}$, more accurately and outperform their counterparts without basis expansions. However, the IPW estimator remains sensitive to propensity scores near 0 or 1, which results in large inverse weights and unstable estimates. In contrast, the AIPW signals are far more stable and concentrated within a narrow band, yielding more precise CME estimates.
Although basis expansions can enhance model flexibility, they also introduce more parameters to estimate, potentially leading to greater variance and a higher risk of overfitting, particularly when the number of covariates is large. Next, we show that regularized methods, such as Lasso, are essential for managing this high-dimensional setting and ensuring stable, reliable inference.
Variable Selection
When the number of covariates or expanded basis functions is large, conventional OLS or logistic regression may overfit and yield unstable estimates. To address this limitation, one can apply Lasso to achieve both regularization and variable selection. Lasso imposes an $\ell_{1}$-penalty on the regression coefficients, shrinking many of them toward zero and effectively selecting a sparse subset of relevant predictors. In its canonical form for a linear model with response $\{y_{i}\}_{i=1}^{n}$ and predictors $\{x_{i}\}_{i=1}^{n}\in\mathbb{R}^{p}$, the Lasso estimate $\hat{\beta}_{\lambda}$ is defined as
$\hat{\beta}_{\lambda}=\arg\min_{\beta\in\mathbb{R}^{p}}\Biggl\{\frac{1}{n}\sum_{i=1}^{n}\bigl(y_{i}-x_{i}^{\top}\beta\bigr)^{2}+\lambda\sum_{j=1}^{p}|\beta_{j}|\Biggr\}.$
The $\ell_{1}$-penalty shrinks many coefficients to zero, and the trade-off between sparsity and goodness of fit is governed by the penalty parameter $\lambda\geq 0$. A straightforward application of Lasso in this context is to estimate separate Lasso regressions for each outcome model (one for the treated and one for the control subsample) and for the propensity-score model. Let
$I_{1}=\{\,i:D_{i}=1\,\},\quad\text{and}\quad I_{0}=\{\,i:D_{i}=0\,\},$
and, with some abuse of notation, let $X_{i}$ denote the vector of predictors—including basis expansions and their interactions—for observation $i$. Then one solves:
$\hat{\beta}_{1,\lambda_{1}}$ $=\arg\min_{\beta_{1}}\left\{\frac{1}{|I_{1}|}\sum_{i\in I_{1}}\bigl(Y_{i}-X_{i}^{\top}\beta_{1}\bigr)^{2}+\lambda_{1}\|\beta_{1}\|_{1}\right\}\quad\Longrightarrow\quad\hat{\mu}^{1}(V_{i})=X_{i}^{\top}\hat{\beta}_{1,\lambda_{1}},$
$\hat{\beta}_{0,\lambda_{0}}$ $=\arg\min_{\beta_{0}}\left\{\frac{1}{|I_{0}|}\sum_{i\in I_{0}}\bigl(Y_{i}-X_{i}^{\top}\beta_{0}\bigr)^{2}+\lambda_{0}\|\beta_{0}\|_{1}\right\}\quad\Longrightarrow\quad\hat{\mu}^{0}(V_{i})=X_{i}^{\top}\hat{\beta}_{0,\lambda_{0}},$
and, using all $n$ observations for the propensity score,
$\hat{\gamma}_{\lambda_{p}}$ $=\arg\min_{\gamma}\left\{-\frac{1}{n}\sum_{i=1}^{n}\left[D_{i}X_{i}^{\top}\gamma-\log\bigl(1+\exp(X_{i}^{\top}\gamma)\bigr)\right]+\lambda_{p}\|\gamma\|_{1}\right\}$
$\Longrightarrow\quad\hat{\pi}(V_{i})=\frac{1}{1+e^{-X_{i}^{\top}\hat{\gamma}_{\lambda_{p}}}}.$
Here, $\lambda_{1}$, $\lambda_{0}$, and $\lambda_{p}$ are penalty parameters for each model that can be selected via cross-validation. The estimated coefficients are then used to construct the outcome or propensity-score models and, subsequently, to build various signals as discussed previously.
Post-Selection Lasso.
A well-known issue with Lasso is that standard errors for its coefficients are not readily available, and inference is not straightforward. Because Lasso produces sparse solutions, one can define the model selected by Lasso as the set of predictors with nonzero coefficients:
$\widehat{\mathcal{A}}_{1}=\left\{j:\hat{\beta}_{1,\lambda_{1},j}\neq 0\right\},\quad\widehat{\mathcal{A}}_{0}=\left\{j:\hat{\beta}_{0,\lambda_{0},j}\neq 0\right\},\quad\widehat{\mathcal{A}}_{p}=\left\{j:\hat{\gamma}_{\lambda_{p},j}\neq 0\right\}.$
Post-selection coefficients are obtained by refitting the model using the appropriate unpenalized estimator—OLS for linear models and logistic regression for logit models—on the variables selected by Lasso. Specifically, the post-selection coefficients for $\beta_{1,\lambda_{1}}$, denoted $\tilde{\beta}_{1,\lambda_{1}}$, are estimated by running a standard OLS regression on the treated units using only the predictors in $\widehat{\mathcal{A}}_{1}$; similarly, one obtains $\tilde{\beta}_{0,\lambda_{0}}$ and $\tilde{\gamma}_{\lambda_{p}}$. The various signals are then constructed based on these refined estimates.
belloni2012sparse and belloni2013least provide results showing that predictions based on post-selection coefficients perform at least as well as those based on penalized coefficients. Under the conditions that guarantee valid predictions by Lasso, the post-selection coefficients often yield slightly improved performance, and in other cases, they perform comparably.
Example 10 (Many Covariates and Complex Nonlinear Relationships)
We simulate a sample of 1,000 units with $V_{i}=(X_{i},Z_{i1},Z_{i2},\dots,Z_{i,8})$, where $X_{i}\sim\mathrm{Unif}[-2,2]$ and each $Z_{j}\sim\mathrm{Unif}[0,1]$ for $j=1,2,\dots,8$. The treatment and outcome are generated in the same way as in Example 9, meaning that covariates $Z_{i4},\dots,Z_{i,8}$ are redundant. The true CME is $\theta(x)=1-x^{2}$. Unconfoundedness holds: $D_{i}\perp\!\!\!\perp(Y_{i}(0),Y_{i}(1))\mid V_{i}$.
Figure 10(b) presents the AIPW signals and CME estimates without and with Lasso selection. Because the additional variables $Z_{i4},\dots,Z_{i,10}$ are included despite having little influence on the treatment or outcome, identifying the relevant covariates becomes nontrivial. Figure 10(b) illustrates how Lasso improves estimation by selecting a sparse, informative subset of predictors, leading to drastically more stable signals and more accurate CME estimates. Without the variable selection, the basis expansions and interaction terms over $(X,Z)$ make the estimated signals noisy and the CME estimation unstable.
Figure 10: AIPW with Basis Expansion: w/o and w/ Variable Selection
(a)   Without variable selection
(b)   With variable selection
Note: Dark blue and light blue dots represent AIPW signals for the treatment and control groups, respectively. The red line represents the true CME, while the orange line represents the estimated CME.
In summary, the outcome-modeling, IPW, and AIPW estimators for the CME follow a common three-step procedure: (1) model the outcome and/or the propensity score; (2) construct signals; and (3) smooth the signals over the moderator $X$. IPW relies solely on propensity scores to balance treated and control units—an approach that is simple but can suffer from high variance when scores are near 0 or 1. AIPW combines both models and offers double robustness: as long as either the outcome or propensity-score model is correctly specified, the CME estimates remain consistent. It is important to emphasize that valid identification and inference across all methods require unconfoundedness and overlap assumptions; if these assumptions fail—e.g., due to unobserved confounding—the resulting CME estimates can be biased and/or unstable, regardless of modeling choices.
In the remainder of the Element, we refer to the AIPW estimator with basis expansion and post-Lasso selection as AIPW-Lasso, and refer to the post-Lasso estimators based on outcome signals and IPW signals as outcome-Lasso and IPW-Lasso, respectively. We summarize the AIPW-Lasso procedure in Algorithm 1.
Algorithm 1 AIPW-Lasso
1:Inputs:
2:  Data: $\{(X_{i},Z_{i},D_{i},Y_{i})\}$ for $i=1,\dots,n$
3:    $D_{i}\in\{0,1\}$ is a binary treatment
4:    $V_{i}:=(X_{i},Z_{i})$ are covariates, of which $X_{i}$ is the moderator
5:    $Y_{i}$ is the outcome
6:  Choice of basis functions $\psi(\cdot)$ and smoothing method
7:Outputs:
8:  Estimated function $\hat{\theta}(x)=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid X_{i}=x]$
9:1) Estimate outcome model coefficients
10:  * Using treated units $\{i:D_{i}=1\}$:
11:    - Run Lasso regression with basis expansion $\psi(V_{i})$ to select active set $\widehat{\mathcal{A}}_{1}$
12:    - Run OLS on selected covariates $\widehat{\mathcal{A}}_{1}$ to obtain coefficients $\hat{\beta}_{1}$
13:  * Using control units $\{i:D_{i}=0\}$:
14:    - Run Lasso regression with basis expansion $\psi(V_{i})$ to select active set $\widehat{\mathcal{A}}_{0}$
15:    - Run OLS on selected covariates $\widehat{\mathcal{A}}_{0}$ to obtain coefficients $\hat{\beta}_{0}$
16:2) Estimate propensity-score coefficients
17:  * Using all units:
18:    - Run logit Lasso regression with basis expansion $\psi(V_{i})$ to select active set $\widehat{\mathcal{A}}_{p}$
19:    - Run unpenalized logit regression on $\widehat{\mathcal{A}}_{p}$ to obtain coefficients $\hat{\gamma}$
20:3) Predict nuisance parameters for all observations
21:  * For all $i=1,\dots,n$, using the estimated coefficients and basis $\psi(V_{i})$:
22:    $\hat{\mu}^{1}(V_{i})\leftarrow\text{predict using }\hat{\beta}_{1}$
23:    $\hat{\mu}^{0}(V_{i})\leftarrow\text{predict using }\hat{\beta}_{0}$
24:    $\hat{\pi}(V_{i})\leftarrow\text{predict using }\hat{\gamma}$
25:4) Compute AIPW signals
26:  $\hat{\Lambda}_{i}=\hat{\mu}^{1}(V_{i})-\hat{\mu}^{0}(V_{i})+\frac{D_{i}}{\hat{\pi}(V_{i})}(Y_{i}-\hat{\mu}^{1}(V_{i}))-\frac{1-D_{i}}{1-\hat{\pi}(V_{i})}(Y_{i}-\hat{\mu}^{0}(V_{i}))$
27:5) Project AIPW signals onto $X$ to obtain $\hat{\theta}(x)$
28:  Regress $\hat{\Lambda}_{i}$ on $X_{i}$, using kernel or B-spline regression (with cross-validation for hyperparameter selections) $\Rightarrow\hat{\theta}(x)$.
29:Return:
30:  The function $\hat{\theta}(x)$ = estimated CME as a function of $x$.
3.4 Inference
Constructing confidence intervals analytically for the CME is challenging due to the multi-step nature of the estimation procedure. Our aim is to capture uncertainty arising from two main sources: (1) the estimation of the outcome and propensity-score models, which includes uncertainty from Lasso-based variable selection; and (2) the smoothing step that derives the CME from the resulting signals. We do not account for uncertainty from tuning or model-selection decisions—such as the choice of knots in the B-spline expansion, the selection of the Lasso penalty parameter $\lambda$, or the choice of the span parameter in kernel regression.
To approximate the sampling distribution of the CME estimates $\hat{\theta}(x)$, we implement a nonparametric bootstrap procedure that replicates the main estimation pipeline. In each bootstrap iteration, we resample the data with replacement and re-run the Lasso variable-selection step—using the same fixed penalty parameter as in the original estimation—to identify the active covariates. We then refit the outcome and propensity-score models using OLS and logistic regression on the selected covariates. The CME is subsequently estimated via kernel regression, using the span parameter fixed at its original cross-validated value. The B-spline expansion (including knot placement and polynomial degree) is also held fixed across bootstrap iterations.
Pointwise confidence intervals for $\theta(x)$ at each evaluation point $x$ can be constructed using either a percentile-based method or a normal approximation based on the empirical bootstrap variance. Similarly, uniform confidence intervals that ensure simultaneous coverage across the entire range of $X$ can be constructed following the approaches outlined in previous chapters.
Figure 11: Estimated CME with Simulated Data: Kernel vs. Post-Lasso Methods
(a)   Kernel Estimator
(b)   Outcome-Lasso
(c)   IPW-Lasso
(d)   AIPW-Lasso
Note: In each figure, the red line represents the true CME; the black line represents the estimated CME. The shaded area and the dashed lines represent the 95% pointwise and uniform confidence intervals, respectively. The histograms at the bottom of the figure depict the distributions of $X$ across treatment (pink) and control (gray) groups.
In Figure 11, we display the estimated CME curves along with pointwise and uniform confidence intervals on the simulated data from Example 10, using the kernel estimator introduced in Chapter 2, the outcome-modeling approach (post-Lasso), the IPW-Lasso estimator, and the AIPW-Lasso estimator. The outcome-modeling approach yields an estimated CME that deviates slightly from the true CME when $X$ is relatively small. The IPW-Lasso estimator produces noisier estimates and wider confidence intervals. The AIPW-Lasso estimator provides an estimated CME that more closely aligns with the truth than the other post-Lasso estimators. Interestingly, the kernel estimator performs fairly well, too.
Finally, we apply four methods, including the kernel estimator, outcome-Lasso, IPW-Lasso, and AIPW-Lasso, to the data from Example 6. The last three methods all employ basis expansion and post-selection-Lasso regression. The code snippet below demonstrates how to implement all three procedures using the interflex package:
1D <- "opp"
2Y <- "pres_ref"
3X <- "pres_vote_margin"
4Z <- c("prev_vote","majority","leadership","seniority","tot_speech")
5d <- d[which(d[,X]<=16.5 & d[,X]>=-4.1),]
6
7## Use AIPW Signals
8
9est.aipw<-interflex(estimator=’lasso’,signal = ’aipw’,
10 data = d, nboots = 2000,
11 Y=Y,D=D,X=X, Z = Z,
12 treat.type = "discrete",
13 reduce.dimension = ’kernel’,
14 na.rm = TRUE)
15
16## Use IPW Signals
17
18est.ipw<-interflex(estimator=’lasso’,signal = ’ipw’,
19 data = d, nboots = 2000,
20 Y=Y,D=D,X=X, Z = Z,
21 treat.type = "discrete",
22 reduce.dimension = ’kernel’,
23 na.rm = TRUE)
24
25## Use Outcome Modeling Signals
26
27est.outcome<-interflex(estimator=’lasso’,signal = ’outcome’,
28 data = d, nboots = 2000,
29 Y=Y,D=D,X=X, Z = Z,
30 treat.type = "discrete",
31 reduce.dimension = ’kernel’,
32 na.rm = TRUE)
Figure 12: Replicating noble2024presidential Figure 2: Kernel vs Post-Lasso-Methods
(a)   Kernel Estimator
(b)   Outcome-Lasso
(c)   IPW-Lasso
(d)   AIPW-Lasso
Note: The treatment is out-partisanship of lawmakers, the outcome is frequency of presidential references in Congressional speeches. In each figure, the black line represents the estimated CME. The shaded area and the dashed lines represent the pointwise and uniform confidence intervals, respectively. The histograms at the bottom of the figure depict the distributions of $X$ across treatment (pink) and control (gray) groups.
In Figure 12, we display the estimated CME along with their pointwise and uniform confidence intervals. The gray band represents the pointwise confidence intervals, while the dashed lines indicate the uniform confidence intervals that ensure simultaneous coverage across all values of $x$. The uniform intervals are wider because they capture the uncertainty of the entire estimated curve rather than at individual points. Overall, the figure demonstrates that the main conclusion from noble2024presidential remains robust across various estimators. The kernel estimator selects a large bandwidth via cross-validation, resulting in a nearly linear estimated CME. The outcome modeling approach also shows a decreasing CME, although the rate of decrease is smaller than that estimated by the kernel method. The IPW estimator exhibits a clear decreasing pattern but comes with noisy estimates and wide confidence intervals. Finally, the AIPW estimator produces narrower confidence intervals than the outcome modeling approach and generally shows a weakly decreasing trend; however, when the local vote margin is not lopsided (i.e., $5\leq X\leq 5$), the estimated CME does not appear strictly monotonic.
3.5 Continuous Treatment
We now consider a continuous treatment $D$ and still focus on the estimation of conditional marginal effect:
$\operatorname{CME}(x)\;=\;\mathbb{E}\!\left[\mu_{d}(D,X,Z)\,\middle|\,X=x\right],$
where $\mu(d,x,z)\equiv\mathbb{E}[Y(d)\mid X=x,Z=z]$ denotes the expected potential outcome and $\mu_{d}(d,x,z)\equiv\partial\mu(d,x,z)/\partial d$ is its derivative with respect to $d$. For ease of identification, we adopt a partially linear regression model (PLRM), first introduced by robinson1988root. This model assumes the potential outcome $Y(d)$ is a linear function of $d$, but the coefficient on $d$ can vary with $X$.
Assumption 9 (Partially Linear Regression Model with $X$ as the Moderator)
Let $V\equiv(X,Z)$. There exist functions $\theta:\mathcal{X}\to\mathbb{R}$, $g:\mathcal{V}\to\mathbb{R}$, and $m_{D}:\mathcal{V}\to\mathbb{R}$ such that the potential outcome $Y_{i}(d)$ and the treatment assignment $D_{i}$ satisfy:
$Y_{i}(d)\;=\;\theta(X_{i})\,d\;+\;g(V_{i})\;+\;\zeta_{i},$
$D_{i}\;=\;m_{D}(V_{i})\;+\;\varepsilon_{i},\qquad\mathbb{E}\!\left[\varepsilon_{i}\mid V_{i}\right]=0,$
Under this assumption, the observed outcome is $Y_{i}=Y_{i}(D_{i})=\theta(X_{i})D_{i}+g(V_{i})+\zeta_{i}$. Combined with unconfoundedness (which implies $\mathbb{E}[\zeta_{i}\mid D_{i},V_{i}]=0$), the conditional potential outcome function is:
$\mu(d,x,z)\equiv\mathbb{E}[Y(d)\mid X=x,Z=z]=\theta(x)d+g(x,z)$
The derivative with respect to $d$ is simply:
$\mu_{d}(d,x,z)\equiv\partial\mu(d,x,z)/\partial d=\theta(x)$
Therefore, the CME simplifies to $\mathbb{E}\!\left[\mu_{d}(D,X,Z)\,\middle|\,X=x\right]=\mathbb{E}\!\left[\theta(X)\,\middle|\,X=x\right]=\theta(x)$. Our estimation target is therefore the function $\theta(x)$ itself. This function precisely captures the marginal effect of $D$ on $Y$, and the model assumes this effect varies only with $X$.
To elucidate the identification strategy, we use the “partialing out” approach. Note that:
$\mathbb{E}[Y_{i}\mid V_{i}]=\mathbb{E}[\theta(X_{i})\,D_{i}\mid V_{i}]+\mathbb{E}[g(V_{i})\mid V_{i}]+\mathbb{E}[\zeta_{i}\mid V_{i}].$
By our assumptions, $\mathbb{E}[g(V_{i})\mid V_{i}]=g(V_{i})$, $\mathbb{E}[\zeta_{i}\mid V_{i}]=0$ (by unconfoundedness), and $\mathbb{E}[\theta(X_{i})\,D_{i}\mid V_{i}]=\theta(X_{i})\,\mathbb{E}[D_{i}\mid V_{i}]$ since $X_{i}$ is a component of $V_{i}$. We also have $\mathbb{E}[D_{i}\mid V_{i}]=m_{D}(V_{i})$. Substituting these yields:
$\mathbb{E}[Y_{i}\mid V_{i}]=\theta(X_{i})\,m_{D}(V_{i})+g(V_{i}).$
Now, define the residualized outcome $\tilde{Y}_{i}\equiv Y_{i}-\mathbb{E}[Y_{i}\mid V_{i}]$ and residualized treatment $\tilde{D}_{i}\equiv D_{i}-\mathbb{E}[D_{i}\mid V_{i}]$. We have
$\tilde{Y}_{i}$ $=\left(\theta(X_{i})\,D_{i}+g(V_{i})+\zeta_{i}\right)-\left(\theta(X_{i})\,m_{D}(V_{i})+g(V_{i})\right)$
$=\theta(X_{i})\,(D_{i}-m_{D}(V_{i}))+\zeta_{i}$
$=\theta(X_{i})\,\tilde{D}_{i}+\zeta_{i}.$
In this final equation, $\tilde{Y}_{i}$ and $\tilde{D}_{i}$ represent the residualized outcome and treatment after removing the effects of all covariates in $V_{i}$. chernozhukov2015post gives a thorough discussion of this “partialing out” procedure. Intuitively, we are only removing the confounding variation in $D$ attributable to $V=(X,Z)$; the CME is the slope of $Y$ with respect to the remaining, orthogonalized part of $D$, indexed by $X$.
In practice, we estimate $\mathbb{E}[Y\mid V_{i}]$ and $\mathbb{E}[D_{i}\mid V_{i}]$ using flexible methods—such as basis expansions combined with post-selection Lasso—to “denoise” $Y_{i}$ and $D_{i}$, yielding $\hat{\tilde{Y}}_{i}$ and $\hat{\tilde{D}}_{i}$. We then estimate the CME by regressing $\hat{\tilde{Y}}_{i}$ on $\hat{\tilde{D}}_{i}$, relying on the key property that $\zeta_{i}$ remains orthogonal to $\tilde{D}_{i}$ (i.e., $\mathbb{E}[\zeta_{i}\mid\tilde{D}_{i}]=0$).
For implementation, one approach is to fit the following B-spline regression:
$\hat{\theta}\ \in\ \arg\min_{\theta\in\Theta_{\mathrm{bs}}}\mathbb{E}\Bigl[\bigl(\hat{\tilde{Y}}_{i}-\theta(X_{i})\,\hat{\tilde{D}}_{i}\bigr)^{2}\Bigr],$
Under the sieve representation, $\theta(x)\approx\beta^{T}\phi(x)$, where $\phi(x)$ is a set of B-spline basis functions.
Alternatively, one can use a kernel-based local linear estimator introduced in Chapter 2. From our derivation, the population model for the residuals is:
$\tilde{Y}_{i}=\theta(X_{i})\tilde{D}_{i}+\zeta_{i},\quad\mathbb{E}[\zeta_{i}\mid X_{i},\tilde{D}_{i}]=0.$
To estimate $\theta(X=x_{0})$, we can use a local linear approximation for $\theta(x)$:
$\theta(X_{i})\approx\alpha(x_{0})+\beta(x_{0})(X_{i}-x_{0})\quad\text{for }X_{i}\text{ near }x_{0}.$
The local minimization problem is then:
$\left(\hat{\alpha}(x_{0}),\hat{\beta}(x_{0})\right)=\arg\min_{\tilde{\alpha},\,\tilde{\beta}}\,\sum_{i=1}^{N}\left[\hat{\tilde{Y}}_{i}-\left(\tilde{\alpha}+\tilde{\beta}(X_{i}-x_{0})\right)\hat{\tilde{D}}_{i}\right]^{2}K\left(\frac{X_{i}-x_{0}}{h(x_{0})}\right).$
The estimate of the CME at $x_{0}$ is $\hat{\theta}(x_{0})=\hat{\alpha}(x_{0})$. Researchers can select the bandwidth function $h(x_{0})$ via cross-validation. We summarize this partial-out Lasso (PO-Lasso) procedure in Algorithm 2, which is modified based on chernozhukov2015post.
Algorithm 2 PO-Lasso for CME
1:Inputs:
2:  Data: $\{(X_{i},Z_{i},D_{i},Y_{i})\}$ for i=1,,ni=1,...,n
3:    $D_{i}$ is a continuous treatment
4:    $V_{i}:=(X_{i},Z_{i})$ are covariates, in which $X_{i}$ is the moderator
5:    $Y_{i}$ is the outcome
6:  Choice of basis functions (e.g., B-splines) $\psi(V_{i})$
7:Outputs:
8:  Estimated function $\hat{\theta}(x)=\widehat{\operatorname{CME}}(x)$
9:1) Basis Expansion
10:  * Conduct basis expansion of $V_{i}$, obtain $\psi(V_{i})$.
11:2) Partial out covariates from the outcome (Post-Lasso)
12:  * Run a Lasso regression of $Y_{i}$ on $\psi(V_{i})$ to select the active set of basis functions, $\widehat{\mathcal{A}}_{y}$.
13:  * Run an OLS regression of $Y_{i}$ on the selected basis functions indexed by $\widehat{\mathcal{A}}_{y}$.
14:  * Compute fitted values $\hat{m}_{Y}(V_{i})=\hat{\mathbb{E}}[Y_{i}\mid V_{i}]$ from this OLS regression.
15:  * Compute residuals $\hat{\tilde{Y}}_{i}=Y_{i}-\hat{m}_{Y}(V_{i})$ for all $i$.
16:3) Partial out covariates from the treatment (Post-Lasso)
17:  * Run a Lasso regression of $D_{i}$ on $\psi(V_{i})$ to select the active set of basis functions, $\widehat{\mathcal{A}}_{d}$.
18:  * Run an OLS regression of $D_{i}$ on the selected basis functions indexed by $\widehat{\mathcal{A}}_{d}$.
19:  * Compute fitted values $\hat{m}_{D}(V_{i})=\hat{\mathbb{E}}[D_{i}\mid V_{i}]$ from this OLS regression.
20:  * Compute residuals $\hat{\tilde{D}}_{i}=D_{i}-\hat{m}_{D}(V_{i})$ for all $i$.
21:4) Estimate $\theta(x)$ using residuals
22:  * Estimate the varying-coefficient model $\hat{\tilde{Y}}_{i}=\theta(X_{i})\hat{\tilde{D}}_{i}+\zeta_{i}$ using the estimated residuals $(\hat{\tilde{Y}}_{i},\hat{\tilde{D}}_{i})$ and the moderator $X_{i}$.
23:  * This estimation can use B-splines (interacting $\hat{\tilde{D}}_{i}$ with basis of $X_{i}$) or local linear regression (regressing $\hat{\tilde{Y}}_{i}$ on $\hat{\tilde{D}}_{i}$ with kernel weights).
24:Return:
25:  The function $\hat{\theta}(x)$ = estimated CME as a function of $x$.
We provide a simulated example below.
Example 11 (Complex DGP with a Continuous Treatment)
We simulate a sample of $1000$ observations. The moderator $X_{i}\sim\mathrm{Unif}[-2,2]$. Four additional covariates $\{Z_{i1},Z_{i2},Z_{i3},Z_{i4}\}$ are sampled from $\mathcal{N}(0,1)$. The continuous treatment $D_{i}$ is generated as:
$D_{i}=\underbrace{0.5Z_{i1}+X_{i}^{2}}_{m_{D}(V_{i})}+\varepsilon_{i}.$
where $\varepsilon_{i}\sim\mathcal{N}(0,1)$. The outcome $Y_{i}$ is generated as:
$Y_{i}=1+1.5X_{i}+D_{i}-D_{i}X_{i}^{2}+2X_{i}\exp(1+Z_{i1})+2\,\mathbf{1}\{Z_{i2}>0\}Z_{2}+\zeta_{i}.$
with $\zeta_{i}\sim\mathcal{N}(0,1)$. We can rewrite the outcome equation to match the PLRM form $Y_{i}=\theta(X_{i})D_{i}+g(V_{i})+\zeta_{i}$:
$Y_{i}=\underbrace{(1-X_{i}^{2})}_{\theta(X_{i})}D_{i}+\underbrace{\left(1+1.5X_{i}+2X_{i}\exp(1+Z_{i1})+2\,\mathbf{1}\{Z_{i2}>0\}Z_{2}\right)}_{g(V_{i})}+\zeta_{i}.$
The true CME is therefore $\theta(x)=1-x^{2}$. The challenge comes from $g(V_{i})$, a highly complex, non-linear function of $X$ and $Z$ that acts as a confounder.
In Figure 13(b), we compare two estimation strategies using data from Example 11. In panel (a), we apply the kernel estimator discussed in Chapter 2. Because it cannot effectively accommodate the strong nonlinearity in $Z_{i}$ (i.e., the $g(V_{i})$ term), the resulting black curve diverges from the true effect (red), especially in the tails of $X_{i}$. In contrast, panel (b) shows the result from PO-Lasso using kernel regression for smoothing. Here, the black curve closely aligns with the true effect (red) across most of the domain. This illustrates how partialing out can correct for distortions caused by nonlinearities and covariate interactions in the DGP, yielding a substantially more accurate estimate of the CME.
Figure 13: CME Estimates for Example 4: Continuous Treatment
(a)   Kernel Estimator
(b)   PO-Lasso
Note: In each figure, the red line represents the true CME; the black line represents the estimated CME. The shaded area and the dashed lines represent the 95% pointwise and uniform confidence intervals, respectively. The histograms at the bottom of the figure depict the distributions of $X$.
Finally, we illustrate PO-Lasso using an empirical example from political science.
Example 12
adiguzel2023out examines the impact of the incumbent’s public health provision on electoral outcomes. The authors use a reform in Turkey that significantly altered the geographic distribution of health clinics. Here, we replicate Figure 4A in the study. The authors find that reduced congestion levels significantly increased the AKP’s vote share, especially in poorer communities which exhibited a more substantial response to the improvements in health care access. The variables of interest are:
  • Outcome: Change in the vote share for the incumbent AKP party (continuous, $\in[-33.9,39.5]$)
  • Treatment: Change in congestion (continuous, $\in[-28.8,20.8]$); congestion is measured by the number of patients per doctor at the nearest clinic, reducing congestion indicating improved service quality
  • Moderator: Logarithm of median property value (continuous, $\in[-0.2,7.7]$)
We apply the kernel estimator from Chapter 2 and the PO-Lasso estimator. We use the B-spline regression in the last step of estimating $\theta(x)$ as it is less time consuming. The code snippet below demonstrates how to implement the PO-Lasso estimator using the interflex package. We trim the data to exclude moderator values where the distribution is sparse.
1D <- "Dpop1000_asm_dr_3"
2Y <- "Dakp_3"
3X <- "lograyic09"
4Z <- c(’Duniversity_3’, ’Dodr_3’, ’Dydr_3’, ’Dpopulation_3’, ’Dhospital_pri_3’, ’Dhospital_pub_3’)
5
6df<-df %>%
7 dplyr::filter(lograyic09 <=7) %>%
8 dplyr::filter(lograyic09 >=3)
9
10est.lasso<-interflex(estimator=’lasso’,
11 data = df, nboots = 2000,
12 Y=Y,D=D,X=X, Z = Z,
13 na.rm = TRUE)
In Figure 14(b), we display the estimated CME along with their 95% pointwise and uniform confidence intervals. The gray band represents pointwise intervals, while the dashed lines indicate uniform intervals that ensure simultaneous coverage across all values of $x$. Both the kernel estimator and PO-Lasso yield broadly consistent findings: poorer neighborhoods with $X<4$ respond more strongly to improvements in health care, while the effect is essentially nonexistent in other areas. Although the main argument in adiguzel2023out remains supported, the more flexible estimators reveal a more nuanced pattern than the original linear model.
Figure 14: Replicating adiguzel2023out Figure 4A
(a)   Kernel Estimator
(b)   PO-Lasso
Note: The treatment is the change in congestion level, the outcome is the change in the vote share for the incumbent AKP party. In each figure, the black line represents the estimated CME. The shaded area and the dashed lines represent the 95% pointwise and uniform confidence intervals, respectively. The histograms at the bottom of the figure depict the distributions of the moderator.
3.6 Summary
This chapter introduces the AIPW-Lasso and PO-Lasso estimators. We begin by discussing the limitations of existing approaches, such as neglecting the treatment assignment mechanism and potential nonlinearities or interactions in covariates.
To address these issues, we first review the outcome-modeling and IPW approaches in the continuous moderator setting. The main limitation of IPW is its sensitivity to extreme weights and specification of the propensity score model. We then introduce the AIPW estimator, which integrates outcome modeling with propensity score weighting. This combination offers the double robustness property—consistency provided at least one of these two models is correctly specified. By residualizing outcomes with respect to both treatment and covariates, AIPW reduces bias and variance compared to IPW alone. We illustrate these properties using simulated data, showing that AIPW signals, compared to those from outcome modeling or IPW alone, yield more accurate CME estimates even under model misspecification.
Next, we emphasize improving model flexibility through basis expansions and discuss regularization via Lasso to manage high-dimensional covariates. Using simulated and empirical examples, we demonstrate that basis expansion and variable selection through Lasso significantly improve the accuracy and stability of CME estimates. We call this method AIPW-Lasso. We also discuss implementing inference for CMEs using a bootstrap procedure that accounts for estimation uncertainty.
Lastly, we discuss the estimation of CME under a continuous treatment variable through partial linear regression models. This extension involves “partialing out” Lasso-selected covariates from the treatment and outcome, leading to more reliable CME estimation with complex DGPs. We therefore refer to this method as PO-Lasso.
Chapter 4. Double Machine Learning
In previous chapters, we introduced IPW and AIPW methods for estimating the CME. With AIPW, we emphasized the idea of combining both outcome and treatment models to obtain a doubly robust estimator that remains valid under moderate misspecification of either model. In this chapter, we formally introduce the concept of Neyman orthogonality and show that the AIPW signals satisfy this condition. Neyman orthogonality is a central property underlying recent advances in doubly robust estimation and high-dimensional inference literature. Intuitively, it ensures that small modeling errors in the nuisance functions do not introduce first-order bias in the final effect estimate, such as the CME.
In this Element, we focus on the CME $\theta(x)$ as the primary causal estimand. Here, $\theta(\cdot)$ is a target function of a single moderator $X$. A key step in identifying the CME is estimating the nuisance functions, denoted by $\eta_{0}(V)$, which depend on the full set of high-dimensional covariates $V$. These nuisance components—such as the propensity score $\Pr(D_{i}=1\mid V_{i})$, the outcome model $\mathbb{E}[Y_{i}(d)\mid V_{i}]$, or conditional variances—can exhibit complexities that exceed the capacity of parametric or semiparametric models to capture accurately. Although approaches like IPW, AIPW, and post-Lasso expansions handle moderate nonlinearity or moderate-dimensional covariates, they may struggle with truly high-dimensional or strongly nonlinear data structures.
Flexible machine learning methods address this challenge by approximating complex relationships in $\eta$ without strong assumptions on functional form. Techniques such as random forests (RF), neural networks (NN), histogram gradient boosting (HGB), and hybrid or ensemble approaches offer considerable flexibility but typically achieve slower convergence rates than parametric models. Intuitively, flexible models require stronger regularization to control complexity and avoid overfitting, thereby introducing regularization bias. While regularization reduces variance and practically balances the bias–variance trade-off, it systematically biases the estimated nuisance functions. Naively “plugging in” these biased machine learning estimates (e.g., the estimated outcome model) into the estimation formula for $\theta(\cdot)$ risks losing the $\sqrt{n}$-consistency required for asymptotic normality and valid hypothesis testing. This is the specific problem that Neyman orthogonality is designed to solve (chernozhukov2018double; semenova2021debiased).
Double/debiased machine learning (DML), also known as “double machine learning” or “orthogonalized machine learning,” provides the formal framework to solve this regularization problem. It builds on the same principle of Neyman orthogonality found in AIPW but generalizes it to accommodate high-dimensional nuisance functions $\eta$ estimated via machine learning. The DML framework mitigates the specific biases introduced by flexible estimation through two key strategies: orthogonalization, which corrects for regularization bias, and cross-fitting, which corrects for overfitting bias. In modern high-dimensional inference, the DML framework was first developed by chernozhukov2018double for estimating fixed-dimensional parameters. semenova2021debiased extend it by allowing the target parameter to be a function, which is an infinite-dimensional parameter, such as the CME $\theta(x)$.
In its canonical form, DML operates through a two-stage procedure, designed primarily for scalar causal parameters such as the ATE. In the first stage, one uses machine learning methods with cross-fitting to estimate the nuisance functions. In the second stage, one constructs a Neyman-orthogonal score from nuisance estimates and solves the corresponding moment condition to obtain the target parameter, enjoying $\sqrt{n}$-rate convergence and asymptotic normality under mild assumptions. However, if the estimand is a function rather than a scalar, such as the CME $\theta(x)$, an additional smoothing step is required. Similar to the smoothing step introduced in the previous chapter, after building the orthogonal signals, one projects the signal onto a set of basis functions that scales with the sample size for dimension reduction (semenova2021debiased). Thus, while the core DML machinery—orthogonality and cross-fitting—handles the nuisance estimation and removes first-order bias, we employ a final smoothing step to “collapse out” the remaining covariates $Z$ and recover CME as a function of $X=x$.
The remainder of the chapter unfolds as follows: first, we introduce the key components of the general DML framework, allowing $\theta_{0}\in\Theta$ to be either a finite-dimensional vector or an element of an infinite-dimensional function space. Next, we examine causal models of CME under various scenarios, including the interactive regression model (IRM) for binary treatments and the partially linear regression model (PLRM) for continuous treatments. We then elaborate on constructing Neyman-orthogonal score functions, employing cross-fitting procedures, and implementing smoothing techniques within these causal models. We illustrate how to implement these estimators with empirical examples from political science.
4.1 Key Ingredients of DML
The general DML framework includes three key ingredients: (i) Neyman orthogonality, (ii) high-quality machine learning methods, and (iii) sample-splitting and cross-fitting strategies. We first lay out the necessary assumption for the properties of DML estimator to hold for fixed dimensional (scalar) parameter, then lay out the regularity condition to build the pointwise and uniform asymptotic properties of the DML estimator for infinite-dimensional parameter.
Table 3: Key Ingredients of DML
Component Purpose
Neyman orthogonal score Ensures that the estimation of the parameter of interest $\theta_{0}$ is robust against estimation biases of nuisance parameters $\hat{\eta}$.
High-quality machine learning method Ensures the estimated nuisance parameter, $\hat{\eta}$, converges to the true parameter, $\eta$, at a sufficiently fast rate, specifically $o_{p}(n^{-1/4-\delta})$ for some $\delta\geq 0$.
Sample splitting and cross-fitting Utilizes sample splitting to mitigate overfitting bias by independently estimating $\eta$ and $\theta$ using different data subsets.
Neyman Orthogonality
In the previous chapter, we saw that AIPW constructs an estimator for the CME that remains consistent if either the outcome model or the propensity model is correctly specified. This double robustness arises primarily from Neyman orthogonality: the AIPW signals are constructed so that small errors in nuisance functions induce only second-order biases, preventing first-order bias in the CME estimate. Recall that, in the binary-treatment setting, the AIPW signals takes the form:
$\Lambda_{i}=\mu^{1}(V_{i})-\mu^{0}(V_{i})+\frac{D_{i}}{\pi(V_{i})}\bigl(Y_{i}-\mu^{1}(V_{i})\bigr)-\frac{1-D_{i}}{\,1-\pi(V_{i})\,}\bigl(Y_{i}-\mu^{0}(V_{i})\bigr),$ (7)
[…]
$\left.\frac{\partial}{\partial\eta}\Bigl(\mathbb{E}[\Lambda_{i}(\eta)]\Bigr)\right|_{\eta=\eta_{0}}=0,$
where $\eta_{0}$ denotes the true nuisance functions $(\mu^{0}_{0},\mu^{1}_{0},\,\pi_{0})$. This derivative is taken with respect to local perturbations of true $\eta$. We provide proofs in the Appendix showing that the partial derivative with respect to the nuisance functions is zero, meaning that small errors in estimating $\eta$ do not affect the target parameter $\theta(\cdot)$ to the first order. In the AIPW formula, Neyman orthogonality underlies the double robustness property: even if one nuisance function is misspecified, the main parameter $\theta(\cdot)$ remains consistently estimated.
DML builds on the idea of Neyman orthogonality. When estimating the target function $\theta_{0}$ in the presence of high-dimensional nuisance functions $\eta$, a key concern is the potential influence of regularization bias in $\eta$ on the estimation of $\theta_{0}$. Neyman orthogonality, introduced in neyman1959optimal and neyman1979c, offers a structured approach to reduce the sensitivity of the final estimator to certain classes of errors in the nuisance function estimates. We begin with a generic score $\psi(W_{i},\theta,\eta)$, where $W_{i}=(Y_{i},D_{i},X_{i},Z_{i})$ represents data, to specify a population moment condition:
$\mathbb{E}[\psi(W_{i};\theta_{0},\eta_{0})]=0,$
where $\theta_{0},\eta_{0}$ denote the true CME function and nuisance parameters. Neyman orthogonality requires the score function to satisfy the following condition:
Definition 3 (Neyman orthogonality)
$\frac{\partial}{\partial\eta}\,\mathbb{E}[\psi(W_{i};\theta_{0},\eta)]\bigg|_{\eta=\eta_{0}}=0$
In other words, infinitesimal changes around $\eta_{0}$ do not alter the main moment condition to first order. This is exactly what we observed in the classic AIPW formula: each component involving $\mu^{1}$, $\mu^{0}$, and $\pi$ is constructed so that small estimation errors cancel out. The score function is linear in the parameter of interest as well.
To formalize the Neyman orthogonal property in the context of DML, chernozhukov2018double propose the following assumption, which we restate here in abbreviated form. Let $\psi(W;\theta,\eta)$ be a score function, where $\theta$ is a scalar parameter.
Assumption 10 (Linear scores with approximate Neyman Orthogonality)
Suppose:
  1. (1)
    Identification via moment condition. The true parameter $\theta_{0}$ solves $\mathbb{E}\bigl[\psi(W;\theta_{0},\eta_{0})\bigr]=0.$
  2. (2)
    Linearity in $\theta$. The score is linear in $\theta$. Formally, there exist functions $\psi^{a}(\cdot)$ and $\psi^{b}(\cdot)$ such that
    $\psi(W;\theta,\eta)=\psi^{a}(W;\eta)\,\theta+\psi^{b}(W;\eta).$
  3. (3)
    Smoothness in $\eta$. The map $\eta\mapsto\mathbb{E}\bigl[\psi(W;\theta,\eta)\bigr]$ is twice Gateaux-differentiable in a neighborhood around the true $\eta_{0}$.
  4. (4)
    (Near) Neyman Orthogonality. Small perturbations $\eta-\eta_{0}$ have only a second-order effect on the moment condition. Formally, for a sequence $\delta_{n}\to 0$,
    $\sup_{\eta\in\mathcal{T}_{n}}\left\|\partial_{\eta}\mathbb{E}[\psi(W;\theta_{0},\eta_{0})][\eta-\eta_{0}]\right\|\leq\delta_{n}\,n^{-1/2}.$
  5. (5)
    Identification condition. The matrix $J_{0}=\mathbb{E}\bigl[\psi^{a}(W;\eta_{0})\bigr]$ has singular values bounded away from 0 and $\infty$. This is a standard invertibility requirement to identify $\theta_{0}$.
In practice, constructing the score function begins with a naive moment function that would correctly identify $\theta_{0}$ if $\eta_{0}$ were known exactly. This function is then adjusted—or “de-biased”—to satisfy the orthogonality condition. We will show this process in more details in the context of CME estimation with binary and continuous treatments in the following subsections.
High-quality Machine Learning Method
The double robustness property of the DML approach ensures resilience to mild errors in $\hat{\eta}_{0}$. The key requirement is that the estimated nuisance parameter $\hat{\eta}$ converges to the true parameter $\eta_{0}$ at a sufficiently fast rate, specifically $o_{p}(n^{-1/4-\delta})$ for some $\delta\geq 0$: $\|\hat{\eta}-\eta_{0}\|=o_{p}(n^{-1/4-\delta})$, meaning that the nuisance estimation error vanishes in probability faster than $n^{-1/4-\delta}$.
Assumption 11 (Score regularity and quality of ML estimator)
Suppose the following conditions hold with high probability:
  1. (1)
    Good neighborhood for $\hat{\eta}$. With high probability, the machine learning nuisance estimator $\hat{\eta}$ falls into some set $T_{N}$, which is a neighborhood around $\eta_{0}$. Formally, $\hat{\eta}\in T_{N}$, where $T_{N}$ includes $\eta_{0}$ and is constrained by conditions below.
  2. (2)
    Bounded score moments. There exists an integer $q\geq 2$ and constant $c_{1}$ such that for all $\eta\in T_{N}$,
    $\bigl(\mathbb{E}\bigl[\|\psi(W;\theta_{0},\eta)\|^{q}\bigr]\bigr)^{1/q}\;\leq\;c_{1},\quad\bigl(\mathbb{E}\bigl[\|\psi^{a}(W;\eta)\|^{q}\bigr]\bigr)^{1/q}\;\leq\;c_{1}.$
    This ensures $\psi$ and its derivatives have finite moments, allowing standard limit laws to apply.
  3. (3)
    Controlled rates. For $\eta\in T_{N}$, define quantities (such as $r_{N}$, $r^{\prime}_{N}$, $\lambda_{N}$, and $\lambda^{\prime}_{N}$) capturing how much $\psi$ and $\psi^{a}$ can change when $\eta$ deviates slightly from $\eta_{0}$. We assume these differences remain small, typically on the order of $\delta_{N}\to 0$ as $N$ grows. Concretely:
    $\sup_{\eta\in T_{N}}\left\|\mathbb{E}[\psi^{a}(W;\eta)]-\mathbb{E}[\psi^{a}(W;\eta_{0})]\right\|\;\leq\;\delta_{N},$
    and similarly for $\psi(W;\theta_{0},\eta)$. This guarantees “smoothness” in $\eta$ and that small estimation errors in $\eta$ do not blow up the score.
  4. (4)
    Non-degenerate variance. The variance/covariance of $\psi$ under $\eta_{0}$ is bounded away from zero. Formally, all eigenvalues of
    $\mathbb{E}\bigl[\psi(W;\theta_{0},\eta_{0})\,\psi(W;\theta_{0},\eta_{0})^{\top}\bigr]$
    exceed a positive constant $c_{0}$. This ensures the asymptotic variance is well-defined.
This condition can be satisfied when $\eta_{0}$ is effectively approximated using methods such as decision trees or advanced neural network architectures (chernozhukov2018double). The interflex package implements three machine learning approaches: NN, RF, and HGB.
It is important to note that no single machine learning algorithm performs best across all scenarios. Therefore, no algorithm strictly dominates in all applications; performance varies widely depending on factors such as the specific task and the dimensionality of the dataset. In Table LABEL:tab:ml_techniques, we outline some broadly accepted pros and cons of each machine learning method.
Table 4: Comparison of Machine Learning Techniques
Method Pros Cons
NN • Excel in processing large, complex datasets, particularly those involving unstructured forms such as images, text, and audio.
• Use multiple layers to capture intricate patterns and dependencies within data, facilitating high levels of data abstraction.
• Require careful tuning of many hyperparameters (architecture, learning rate, regularization) to achieve optimal performance.
• Require significantly more data than traditional algorithms to generalize well, making them unsuitable for small datasets.
RF • Effective for classification and regression, leveraging bagging to reduce variance and control overfitting.
• Generally robust against overfitting compared to single decision trees and requires less hyperparameter tuning than boosting.
• Performance can suffer with high-dimensional sparse data, as random feature selection may fail to identify informative splits.
• Cannot extrapolate trends beyond the training data range and struggles with linear relationships compared to parametric models.
HGB • Efficiently handles large datasets by binning continuous features, offering faster training speeds than standard gradient boosting.
• Natively handles missing values and categorical features without extensive preprocessing.
• Effective for imbalanced datasets by focusing on hard-to-classify examples.
• More sensitive to noise and outliers than Random Forests, as the algorithm iteratively corrects errors, potentially overfitting to anomalies.
• Requires more careful tuning of hyperparameters (learning rate, tree depth) than Random Forests to prevent overfitting.
Sample Splitting and Cross-Fitting
Sample splitting serves as a crucial technique alongside the use of orthogonal score functions and machine learning methods for estimating nuisance components. Overfitting bias can arise when the nuisance parameters are estimated on the same sample used to estimate the parameter of interest $\theta$. Sample splitting helps mitigate this bias by ensuring independence between the data used to estimate the nuisance functions ($\eta_{0}$, training data) and the data used to estimate the causal parameter ($\theta_{0}$, holdout data).
Cross-fitting, as an efficient form of data splitting, maximizes the use of available data by alternating the roles of training and holdout sets in a cross-validated framework and combining the estimates from multiple splits to improve efficiency. The process can be outlined as follows:
  1. 1.
    Randomly partition the dataset into $K$ folds, $(I_{k})^{K}_{k=1}$, where each fold $I_{k}$ contains $n=N/K$ observations and $I^{c}_{k}=\{1,\ldots,N\}\setminus I_{k}$ denotes the complement set used for training.
  2. 2.
    For each fold $k$, estimate the nuisance parameters $\hat{\eta}_{0,k}$ using the data in $I^{c}_{k}$. [^8] Strictly speaking, to preserve the fold-level independence (i.e., $\hat{\eta}_{0,k}$ being a function only of the training sample), any hyperparameter tuning—such as selecting the Lasso penalty via cross-validation—should be carried out using only observations in $I_{k}^{c}$, typically through an “inner” cross-validation restricted to $I_{k}^{c}$, followed by refitting the nuisance learner on all of $I_{k}^{c}$ at the chosen hyperparameters. In practice, many implementations tune hyperparameters globally (e.g., cross-validate once on the full sample and reuse the selected hyperparameters across folds) to reduce computational burden.
  3. 3.
    Compute the fold-specific estimator $\hat{\theta}_{0,k}$ using the empirical expectation over observations in $I_{k}$, and average the results across all $K$ folds to obtain the final estimate $\hat{\theta}_{0}$.
The cross-fitting procedure leverages the full dataset while enhancing the robustness of the estimation process by reducing dependence on any single data partition.
The three key ingredients, namely Neyman orthogonality, the high-quality machine learning method, and sample-splitting and cross-fitting strategy, when effectively integrated, lead to the robust properties of the canonical DML estimator.
Theorem 4.1 (Properties of the DML)
Suppose that Assumption 10 and Assumption 11 hold. In addition, suppose that $\delta_{n}\geq n^{-1/2}$ for all $n\geq 1$. Then the DML estimator $\hat{\theta}_{0}$ concentrates in a $1/\sqrt{n}$ neighborhood of $\theta_{0}$ and is approximately linear and centered Gaussian (chernozhukov2018double):
$\sqrt{n}\sigma^{-1}(\hat{\theta}_{0}-\theta_{0})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\bar{\psi}(W_{i})+O_{P}(\rho_{n})\leadsto N(0,I_{d}),$ (8)
uniformly over $P\in\mathcal{P}_{n}$, where the size of the remainder term obeys
$\rho_{n}:=n^{-1/2}+r_{n}+r^{\prime}_{n}+n^{1/2}\lambda_{n}+n^{1/2}\lambda^{\prime}_{n}\leq\delta_{n},$ (9)
$\bar{\psi}(\cdot):=-\sigma^{-1}J_{0}^{-1}\psi(W;\theta_{0},\eta_{0})$ is the influence function, and the approximate variance is:
$\sigma^{2}:=J_{0}^{-1}\mathbb{E}_{P}[\psi(W;\theta_{0},\eta_{0})\psi(W;\theta_{0},\eta_{0})^{\top}](J_{0}^{-1})^{\top}.$ (10)
Equation (8) states that $\hat{\theta}_{0}$ is asymptotically normal with mean $\theta_{0}$ and covariance $\sigma^{2}/n$. Equation (9) defines the remainder term, which captures the cumulative effect of small errors in the final asymptotic expansion. Intuitively, the terms $r_{n}$, $r^{\prime}_{n}$, $\lambda_{n}$, and $\lambda^{\prime}_{n}$ reflect the quality of nuisance function estimation and the degree to which orthogonality mitigates the influence of these errors on the target parameter. Equation (10) gives the asymptotic covariance matrix, where $J_{0}$ denotes the Jacobian of the score function, capturing its sensitivity to deviations of $\theta$ from $\theta_{0}$.
Taken together, the theorem establishes that the DML estimator for the scalar parameter $\theta_{0}$ achieves $\sqrt{n}$-rate convergence and is approximately normally distributed. Both the convergence rate and the asymptotic distribution hold uniformly over a broad class of underlying data-generating processes.
Estimation of Function-Valued Parameters
When $\theta_{0}$ is a function of $x\in\mathcal{X}$, such as the CME, we need an additional smoothing step to handle the infinite-dimensional target. The standard approach is to project $\theta(x)$ onto a finite but growing set of basis functions. For instance, with B-spline expansions, we write:
$\theta_{0}(x)\approx p(x)^{\top}\beta_{0},$
where $p(x)\in\mathbb{R}^{d}$ is a vector of $d$ basis functions (e.g., splines, polynomials). After constructing the orthogonal signal $\hat{\Lambda}_{i}$ (using the cross-fitting procedure), we regress $\hat{\Lambda}_{i}$ on $p(X_{i})$ to obtain the coefficients $\hat{\beta}$. The final estimator of the function takes the form $\hat{\theta}(x)=p(x)^{\top}\hat{\beta}$.
When $\theta_{0}$ is function-valued, DML methods can still achieve desirable asymptotic properties, but the set of required assumptions necessarily expands to govern the choice of $d$. To ensure that $\hat{\theta}(x)$ retains the $\sqrt{n}$-rate and satisfies both pointwise and uniform asymptotic normality, we must manage the trade-off between approximation error and estimation variance. Specifically, semenova2021debiased establish that $d$ must grow large enough to reduce the approximation bias to a negligible level, yet remain small enough relative to the sample size $N$ to control variance and allow for uniform inference. We lay out these specific constraints on identification, basis growth, and approximation error below.
Assumption 12 (Identification and approximation)
Suppose that:
  1. (1)
    Eigenvalues bounded. Let $Q=\mathbb{E}[p(X)p(X)^{\prime}]$ be the population covariance of $p(X)$. We require that the eigenvalues of $Q$ are bounded away from zero and infinity uniformly in $d$. Specifically, there exist constants $0<C_{min}<C_{max}<\infty$ such that $C_{min}\leq\lambda_{min}(Q)\leq\lambda_{max}(Q)\leq C_{max}$.
  2. (2)
    Approximation error. We write $g(x)=p(x)^{\prime}\beta_{0}+r_{g}(x),$ with $r_{g}(x)$ being the approximation error; this assumption requires that both the mean squared and uniform sizes of $r_{g}(x)$ are controlled (see Assumption 3.3 in semenova2021debiased).
When we approximate $\theta_{0}(x)$ by projecting the signal onto $d$ basis functions, each observation contributes to estimating $d$ different coefficients. If $d$ grows too quickly in relation to sample size $N$, the moment-based estimation exhibit high variance. Hence we imposes constraints on the basis growth:
Assumption 13 (Growth condition)
Suppose that:
  1. (1)
    Basis growth. Define $\xi_{d}=\sup_{x\in\mathcal{X}}\|p(x)\|$. This assumption imposes that $\frac{\xi_{d}^{2}\,\log N}{N}=o(1)$ as $N\to\infty$. In words, the dimension $d$ can grow, but not so quickly that $\xi_{d}^{2}\,\log N$ outstrips $N$.
  2. (2)
    Lipschitz Basis Regularity. Define $\alpha(x)=\frac{p(x)}{\|p(x)\|}.$ We say the basis is Lipschitz regular if there exist an integer $m>2$ and constants $\xi^{L}_{d}$ such that:
    1. (a)
      $\xi^{L}_{d}:=\sup_{x,x^{\prime}\in\mathcal{X},\,x\neq x^{\prime}}\frac{\|\alpha(x)-\alpha(x^{\prime})\|}{\|x-x^{\prime}\|}<\infty$
    2. (b)
      $(\xi^{L}_{d})^{\frac{2m}{m-2}}\cdot\frac{\log N}{N}\leq 1,\quad\text{and}\quad\log\xi^{L}_{d}\leq\log d$.
To handle certain empirical-process bounds and smoothness arguments, basis functions must also be “well-behaved” in a Lipschitz sense. In words, $\alpha(x)$ is Lipschitz-continuous in $x$, so the basis does not vary too abruptly over $\mathcal{X}$. Bounding $\xi^{L}_{d}$ in this manner ensures uniform convergence arguments remain valid. Such control in Assumption (2) is crucial for establishing uniform convergence (confidence bands) in high dimensions, whereas simpler growth conditions suffice for pointwise results.
Assumption 14 (Small bias and error)
Suppose that:
  1. (1)
    Error bound on stochastic error. Let U=Yg(X)U=Y-g(X). We assume $\sup_{x\in\mathcal{X}}\mathbb{E}[U^{2}\mid X=x]\leq 1$. The conditional variance is uniformly bounded.
  2. (2)
    Small bias in nuisance estimation. We have cross-fitted nuisance parameters $\hat{\eta}_{k}$. With high probability, $\hat{\eta}_{k}$ lies in a shrinking neighborhood $T_{N}$ around $\eta_{0}$ such that the projection bias of the signal $Y(\eta)$ vanishes sufficiently fast (specifically, $B_{N}=o(1)$ and $\Lambda_{N}=o(1)$ as defined in semenova2021debiased).
  3. (3)
    Tail bounds. There exists $m>2$ such that $\sup_{x\in\mathcal{X}}\mathbb{E}[|U|^{m}\mid X=x]\leq 1.$ This moment condition facilitates strong approximation for uniform inference.
Similarly, Assumption (3) helps ensure stronger concentration inequality, facilitating uniform convergence of the DML estimator.
These assumptions jointly control three sources of error: approximation bias from the sieve representation, estimation error from the nuisance components, and uniform stochastic fluctuations of the orthogonal signal. Together, they allow the $\sqrt{n}$-consistency and asymptotic normality guarantees—typically established for finite-dimensional parameters—to be extended to the infinite-dimensional target $\theta(x)$.
4.2 Binary Treatment
In the case of binary treatment, semenova2021debiased show that the Neyman-orthogonal signal for the CME can be constructed using the following expression:
$\Lambda_{i}(\eta)=\mu(1,V_{i})-\mu(0,V_{i})+\frac{D_{i}(Y_{i}-\mu(1,V_{i}))}{\pi(V_{i})}-\frac{(1-D_{i})(Y_{i}-\mu(0,V_{i}))}{1-\pi(V_{i})}$ (11)
where $\mu(d,v)=\mathbb{E}[Y_{i}\mid D_{i}=d,V_{i}=v]$ is the (observed) outcome model and $\pi(v)=\Pr(D_{i}=1\mid V_{i}=v)$ is the propensity score. We denote the true value of these nuisance parameters as $\eta_{0}(v):=\{\pi_{0}(v),\mu_{0}(1,v),\mu_{0}(0,v)\}$.
[…]
Under the unconfoundedness assumption introduced in Chapter 1 (Assumption 3), the CME can be expressed as the conditional expectation of this observable signal. Specifically, when $Y_{i}(d)\perp\!\!\!\perp D_{i}\mid V_{i}=v$, we have
$\mathbb{E}[\Lambda_{i}(\eta_{0})\mid X_{i}=x]=\theta(x).$
In the DML framework, we do not require the functional forms of the nuisance models to be known parametrically. Instead, we rely on machine learning models to consistently estimate the nuisance components as the sample size becomes sufficiently large.
With binary treatment, we make the following modeling assumption.
Assumption 15 (Interactive Regression Model, IRM)
We assume the potential outcomes and treatment assignment satisfy:
$Y_{i}(d)$ $=g_{0}(d,V_{i})+u_{i},$ $\mathbb{E}[u_{i}$ $\mid V_{i}]=0$
$D_{i}$ $=m_{0}(V_{i})+\epsilon_{i},$ $\mathbb{E}[\epsilon_{i}$ $\mid V_{i}]=0$
where $g_{0}(d,v):=\mathbb{E}[Y(d)\mid V_{i}=v]$ is the structural conditional mean of the potential outcome, and $m_{0}(v):=\Pr(D_{i}=1\mid V_{i}=v)$ is the propensity score.
Under the unconfoundedness assumption ($Y(d)\perp D\mid V$), the structural function $g_{0}(d,v)$ coincides with the observational regression function $\mu_{0}(d,v):=\mathbb{E}[Y\mid D=d,V=v]$. This identification result justifies using machine learning methods to estimate $\mu_{0}$ from observational data.
This model is termed “interactive” because it allows the potential outcomes $Y_{i}(1)$ and $Y_{i}(0)$ to follow arbitrarily different functional forms in $V_{i}$, unlike partially linear models which assume the treatment effect is constant or independent of controls. With the IRM, the CME is given by:
$\theta(x)=\mathbb{E}[g_{0}(1,x,Z)-g_{0}(0,x,Z)\mid X=x].$
The IRM is general, encompassing both the linear interaction model and the smooth varying coefficient model (SVCM) as special cases. Specifically, the linear interaction model assumes
$\mathbb{E}[Y_{i}(d)\mid X_{i},Z_{i}]=\beta_{0}+\beta_{1}d+\beta_{2}X_{i}+\beta_{3}(d\cdot X_{i})+\beta_{4}Z_{i},$
which constrains $g_{0}(\cdot)$ to be linear. Similarly, the SVCM assumes
$\mathbb{E}[Y_{i}(d)\mid X_{i},Z_{i}]=f(X_{i})+q(X_{i})d+Z_{i}\gamma(X_{i}),$
which allows for smooth variation over $X$.
In the binary treatment case, we first estimate the nuisance functions using machine learning methods with cross-fitting. We then construct the AIPW-style Neyman-orthogonal signal for each unit $i$. If $X$ is discrete, the CME can be recovered by taking sample averages of $\{\hat{\Lambda}_{i}\}$ within each group $X_{i}=x$. If $X$ is continuous, we regress $\hat{\Lambda}_{i}$ on basis functions of $X$ (as described in the previous section) to estimate the curve $\theta(x)$.
Algorithm 3 DML Estimation of CME (Binary $D$)
1:Inputs:
2:  Data: $\{(X_{i},Z_{i},D_{i},Y_{i})\}$ for i=1,,ni=1,...,n
3:    $D_{i}\in\{0,1\}$ is binary treatment
4:    $V_{i}:=(X_{i},Z_{i})$ are controls ($X_{i}$ is the moderator)
5:    $Y_{i}$ is the outcome
6:  Number of folds $K$ (e.g., $K=5$)
7:  Choice of basis functions $p(X)$ (e.g., B-splines) for the final projection
8:Outputs:
9:  Estimated function $\hat{\theta}(x)=\widehat{\mathbb{E}}[Y(1)-Y(0)\mid X=x]$
10:1) Partition data into $K$ folds
11:  Randomly split the indices $\{1,...,n\}$ into $K$ disjoint sets: $I_{1},...,I_{K}$.
12:  Let $I_{k}$ be the “holdout” set and $I_{k}^{c}=\{1,...,n\}\setminus I_{k}$ be the “training” set.
13:2) Estimate nuisance functions (cross-fitting)
14:for $k=1$ to $K$ do
15:    (a) On the training set $I_{k}^{c}$:
16:      * Fit outcome models: $\hat{\mu}_{1}(v)\approx\mathbb{E}[Y\mid D=1,V=v]$, $\hat{\mu}_{0}(v)\approx\mathbb{E}[Y\mid D=0,V=v]$
17:      * Fit propensity model: $\hat{\pi}(v)\approx\Pr(D=1\mid V=v)$
18:    (b) For each observation $i$ in the holdout set $I_{k}$:
19:      * Compute predictions: $\hat{\mu}_{1}^{(-k)}(V_{i}),\hat{\mu}_{0}^{(-k)}(V_{i}),\hat{\pi}^{(-k)}(V_{i})$
20:end for
21:3) Compute AIPW-style Neyman-orthogonal signal
22:for each fold $k$, for each $i$ in $I_{k}$ do
23:    Construct the pseudo-outcome $\Lambda_{i}$:
24:    $\Lambda_{i}=\hat{\mu}_{1}^{(-k)}(V_{i})-\hat{\mu}_{0}^{(-k)}(V_{i})$
25:        $\frac{D_{i}(Y_{i}-\hat{\mu}_{1}^{(-k)}(V_{i}))}{\hat{\pi}^{(-k)}(V_{i})}-\frac{(1-D_{i})(Y_{i}-\hat{\mu}_{0}^{(-k)}(V_{i}))}{1-\hat{\pi}^{(-k)}(V_{i})}$
26:end for
27:4) Project signal onto $X$ to obtain $\hat{\theta}(x)$
28:  Run a regression (e.g., OLS) of the signal $\Lambda_{i}$ on the basis functions $p(X_{i})$:
29:  $\hat{\beta}=\arg\min_{b}\sum_{i=1}^{n}\left(\Lambda_{i}-p(X_{i})^{\top}b\right)^{2}$
30:  Construct estimator: $\hat{\theta}(x)=p(x)^{\top}\hat{\beta}$
31:Return:
32:  The function $\hat{\theta}(x)$ = estimated CME.
We now demonstrate how to implement this procedure using the interflex package. Using data from Example 6 based on noble2024presidential, we compare the results from AIPW-Lasso to those from the DML estimators.
1# R code excerpt
2
3library(interflex)
4D="opp" # Out-partisanship of Lawmakers
5Y="pres_ref" # Frequency of Presidential References
6X="pres_vote_margin" #Presidential Vote Margain
7Z=c("prev_vote","majority","leadership","seniority","tot_speech")
8
9## For code to generate plot (a), see last chapter
10
11## DML estimator with neural network machine learning method
12out.dml.nn<-interflex(estimator=’DML’,
13 data = d, model.y="nn", model.t = "nn",
14 Y=Y,D=D,X=X, Z = Z,
15 treat.type = "discrete", na.rm = TRUE,
16 vartype = "bootstrap", parallel = TRUE)
17plot(out.dml.nn)
18
19## DML estimator with random forest machine learning method
20out.dml.rf<-interflex(estimator=’DML’,
21 data = d, model.y="rf", model.t = "rf",
22 Y=Y,D=D,X=X, Z = Z,
23 treat.type = "discrete", na.rm = TRUE,
24 vartype = "bootstrap", parallel = TRUE)
25plot(out.dml.rf)
26
27## DML estimator with histogram gradient boosting machine learning method
28out.dml.hgb<-interflex(estimator=’DML’,
29 data = d, model.y="hgb", model.t = "hgb",
30 Y=Y,D=D,X=X, Z = Z,
31 treat.type = "discrete", na.rm = TRUE,
32 vartype = "bootstrap", parallel = TRUE)
33plot(out.dml.hgb)
Figure 15: Replicating noble2024presidential Figure 2: AIPW-Lasso vs DML Estimators
(a) AIPW-Lasso
(b) DML-NN
(c) DML-RF
(d) DML-HGB
Notes: The treatment $D$ is out-partisanship of lawmakers, the outcome $Y$ is frequency of presidential references in Congressional speeches. (a) CME estimates from AIPW-Lasso; (b)-(d) CME estimates from the DML estimator, with nuisance parameter estimated with NN, RF, and HGB, respectively. The gray ribbon show the 95% pointwise confidence intervals, respectively. The dashed lines represent 95% uniform confidence intervals.
Figure 15 compares CME estimates from the AIPW-Lasso estimator (a) and the DML estimators (b)–(d). As before, we trim the sample based on the moderator $X$, restricting it to the central 95% quantile range to improve overlap. AIPW-Lasso yields monotonically decreasing CME estimates in $x$, with fairly narrow confidence intervals. The DML estimators produce broadly similar patterns but with much wider confidence intervals. Substantively, both methods suggest that the presidential-reference treatment has a modestly positive effect on the outcome in areas of low constituency partisanship. AIPW-Lasso shows a statistically significant treatment effect for $X\in[-5,10]$, while the DML estimators, despite greater uncertainty, indicate a significant effect for out-party legislators when $X\in[0,5]$.
4.3 Continuous Treatment
In the general case of continuous treatment, semenova2021debiased derive a Neyman-orthogonal score that is robust to small estimation errors in the nuisance parameters $\eta_{0}=\{\mu_{0}(d,v),\,f_{D\mid V}(d\mid v)\}$. Here, $\mu_{0}(d,v)=\mathbb{E}[Y\mid D=d,V=v]$ represents the (observed) outcome model, and $f_{D\mid V}$ is the conditional density of the treatment. The general Neyman-orthogonal signal takes the form:
$\Lambda_{i}(\eta)=-\partial_{d}\log f_{D\mid V}(D_{i}\mid V_{i})\,[Y_{i}-\mu(D_{i},V_{i})]+\partial_{d}\mu(D_{i},V_{i}).$
Under standard unconfoundedness and overlap assumptions, the CME is identified as the conditional expectation of this signal:
$\mathbb{E}\!\big[\Lambda_{i}(\eta_{0})\mid X_{i}=x\big]\;=\;\mathbb{E}\!\big[\partial_{d}\mu_{0}(D_{i},V_{i})\mid X_{i}=x\big]\;=\;\theta(x).$
We provide the proof in the Appendix.
While theoretically robust, this general approach requires fully nonparametric estimation of the conditional density $f_{D\mid V}$ and its derivative, which can be highly unstable in high dimensions. To address this practical challenge, we adopt the Partially Linear Regression Model (PLRM) (robinson1988root) as our framework.
Assumption 16 (Partially Linear Regression Model, PLRM)
We assume the outcome and treatment assignment satisfy:
$Y_{i}(d)$ $=\theta(X_{i})\cdot d+g(V_{i})+\epsilon_{i},\quad\mathbb{E}[\epsilon_{i}\mid V_{i}]=0$
$D_{i}$ $=m(V_{i})+\xi_{i},\quad\mathbb{E}[\xi_{i}\mid V_{i}]=0$
where $Y_{i}(d)$ is the potential outcome; $\theta(X_{i})$ is the conditional marginal effect of interest; and $m(V_{i}):=\mathbb{E}[D_{i}\mid V_{i}]$ is the conditional mean of the treatment.
This assumption imposes a key restriction: the outcome $Y_{i}$ must depend linearly on the treatment $D_{i}$, although the slope $\theta(X_{i})$ is allowed to vary arbitrarily with the moderator $X_{i}$. This linearity assumption enables us to replace the complex density-based score with the “partialing out” transformation. By defining the residuals $\tilde{Y}_{i}=Y_{i}-\mathbb{E}[Y_{i}\mid V_{i}]$ and $\tilde{D}_{i}=D_{i}-\mathbb{E}[D_{i}\mid V_{i}]$, the relationship simplifies to:
$\tilde{Y}_{i}=\theta(X_{i})\tilde{D}_{i}+\epsilon_{i}.$
This corresponds to a Neyman-orthogonal moment condition based solely on conditional means $\eta=\{g(V),m(V)\}$:
$\psi(W;\theta,\eta)=\bigl((Y-g(V))-\theta(X)(D-m(V))\bigr)(D-m(V)).$
This score allows us to estimate the CME without estimating densities, making it robust to small estimation errors in $g(V)$ and $m(V)$ while satisfying DML requirements. However, this stability comes at the cost of the linearity assumption. If the true relationship involves higher-order effects (e.g., $Y$ depends on $D^{2}$), the PLRM is misspecified, and the estimator may fail to capture the true marginal effect.
Traditional parametric methods can be viewed as restricted forms of this partially linear model. For instance, if $\theta(\cdot)$ is assumed to be constant, we recover the classic Robinson model. If $\theta(X)$ is linear (i.e., $\theta(X)=\beta_{1}+\beta_{3}X$) and the baseline $g(V)$ is linear in controls, we recover a standard linear interaction model. The semiparametric smooth varying coefficient model (SVCM) is also a special case where $\theta(X)$ is smooth but the baseline effect of controls is restricted to be linear, i.e., $g(V)=f(X)+Z\gamma(X)$.
The key innovation in this framework is the integration of this model with DML. Unlike the PO-Lasso approach discussed in the previous chapter, which typically trains and evaluates residuals on the same sample, DML employs cross-fitting. We partition the data into folds, training the high-dimensional nuisance functions $\hat{m}(V)$ and $\hat{g}(V)$ on training folds and evaluating the orthogonal residuals on the holdout fold. This procedure ensures that the nuisance estimation errors (converging at $o_{P}(n^{-1/4})$) do not bias the final inference. We provide pseudo-code below to clarify this process.
Algorithm 4 DML Estimation of CME (Continuous $D$)
1:Inputs:
2:  Data: $\{(X_{i},Z_{i},D_{i},Y_{i})\}$ for i=1,,ni=1,...,n
3:    $D_{i}\in\mathbb{R}$ is a continuous treatment
4:    $V_{i}:=(X_{i},Z_{i})$ are covariates ($X_{i}$ is the moderator)
5:    $Y_{i}$ is the outcome
6:  Number of folds $K$ (e.g., $K=5$)
7:  Choice of basis functions $p(X)$ (e.g., B-splines)
8:Outputs:
9:  Estimated function $\hat{\theta}(x)=\widehat{\mathbb{E}}[\partial_{d}Y(d)\mid X_{i}=x]$
10:1) Partition data into $K$ folds
11:  Randomly split the indices $\{1,...,n\}$ into $K$ disjoint sets: $I_{1},...,I_{K}$.
12:  Let $I_{k}$ be the “holdout” set and $I_{k}^{c}=\{1,...,n\}\setminus I_{k}$ be the “training” set.
13:2) Estimate nuisance functions (cross-fitting)
14:for $k=1$ to $K$ do
15:    (a) On the training set $I_{k}^{c}$:
16:      * Fit outcome model: $\hat{g}(v)\approx\mathbb{E}[Y\mid V=v]$
17:      * Fit treatment mean model: $\hat{m}(v)\approx\mathbb{E}[D\mid V=v]$
18:    (b) For each observation $i$ in the holdout set $I_{k}$:
19:      * Compute predictions: $\hat{g}^{(-k)}(V_{i}),\hat{m}^{(-k)}(V_{i})$
20:end for
21:3) Form orthogonal residuals
22:for each fold $k$, for each $i$ in $I_{k}$ do
23:    Residualize Outcome: $\tilde{Y}_{i}=Y_{i}-\hat{g}^{(-k)}(V_{i})$
24:    Residualize Treatment: $\tilde{D}_{i}=D_{i}-\hat{m}^{(-k)}(V_{i})$
25:end for
26:  Collect the full set of residuals $\{(\tilde{Y}_{i},\tilde{D}_{i},X_{i})\}$ for $i=1,\dots,n$.
27:4) Project residuals onto $X$ to obtain $\hat{\theta}(x)$
28:  To estimate the varying coefficient $\theta(X)$, regress $\tilde{Y}_{i}$ on the interaction $\tilde{D}_{i}\cdot p(X_{i})$:
29:  $\hat{\beta}=\arg\min_{b}\sum_{i=1}^{n}\left(\tilde{Y}_{i}-\tilde{D}_{i}\cdot(p(X_{i})^{\top}b)\right)^{2}$
30:  Construct estimator: $\hat{\theta}(x)=p(x)^{\top}\hat{\beta}$
31:Return:
32:  The function $\hat{\theta}(x)$ = estimated CME curve.
We provide an R code excerpt to implement the estimation using the interflex package. Using data from Example 12, we compare the results from the PO-Lasso estimator with those from the DML estimators.
1# R code excerpt
2
3library(interflex)
4
5D="Dpop1000_asm_dr_3" #Change in Congestion
6Y="Dakp_3" #Change in AKP vote share
7X="lograyic09" #Property values (2009)
8Z=c(’Duniversity_3’, ’Dodr_3’, ’Dydr_3’, ’Dpopulation_3’, ’Dhospital_pri_3’, ’Dhospital_pub_3’)
9
10## For code to generate plot (a), see last chapter
11
12## DML estimator with neural network machine learning method
13out.dml.nn<-interflex(estimator=’DML’,
14 data = d, model.y="nn", model.t = "nn",
15 Y=Y,D=D,X=X, Z = Z, treat.type = "continuous",
16 na.rm = TRUE,
17 vartype = "bootstrap",parallel = TRUE)
18plot(out.dml.nn)
19
20## DML estimator with random forest machine learning method
21out.dml.rf<-interflex(estimator=’DML’,
22 data = d, model.y="rf", model.t = "rf",
23 Y=Y,D=D,X=X, Z = Z, treat.type = "continuous",
24 na.rm = TRUE,
25 vartype = "bootstrap",parallel = TRUE)
26plot(out.dml.rf)
27
28## DML estimator with histogram gradient boosting machine learning method
29out.dml.hgb<-interflex(estimator=’DML’,
30 data = d, model.y="hgb", model.t = "hgb",
31 Y=Y,D=D,X=X, Z = Z,treat.type = "continuous",
32 na.rm = TRUE,
33 vartype = "bootstrap",parallel = TRUE)
34plot(out.dml.hgb)
Figure 16: Replicating adiguzel2023out
PO-Lasso vs DML Methods
(a) PO-Lasso
(b) DML-NN
(c) DML-RF
(d) DM-HGB
Notes: The treatment $D$ measures change in congestion level, the outcome $Y$ measures change in the vote share for the incumbent AKP party. (a) CME estimates from PO-Lasso; (b)-(d) CME estimates from the DML estimator, with nuisance parameter estimated with NN, RF, and HGB, respectively. The gray ribbon show the 95% pointwise confidence intervals, respectively. The dashed lines represent 95% uniform confidence intervals.
Figure 16 displays CME estimates of the effect of changes in congestion on AKP vote share across communities with varying median property values. The estimation sample is trimmed to the range $3\leq\log(\text{property value})\leq 7$ due to sparse observations outside that interval. We have previous shown the CME estimates from PO-Lasso in the last chapter, now displayed in panel (a). Panels (b)-(d) show that the DML estimator produces similar patterns, with slightly bigger estimates when $X$ is small. They also yield wider confidence intervals.
4.4 Inference
Under the DML regularity conditions, the DML estimator $\hat{\theta}(x)=p(x)^{\top}\hat{\beta}$ is $\sqrt{N}$-consistent and asymptotically normal pointwise:
$\sqrt{N}\,\big(\hat{\theta}(x)-\theta_{0}(x)\big)\ \leadsto\ N\!\big(0,\ \sigma^{2}(x)\big).$
Let $q_{i}$ denote the vector of regressors used in the final projection step (e.g., $q_{i}=p(X_{i})$ for binary treatment or $q_{i}=\tilde{D}_{i}\cdot p(X_{i})$ for continuous treatment). Let $s_{i}$ be the orthogonal signal (the pseudo-outcome $\hat{\Lambda}_{i}$ or $\tilde{Y}_{i}$). Finally, let $u_{i}=s_{i}-q_{i}^{\top}\beta_{0}$ denote the second-stage residual.
We define the population design matrix $J_{0}:=\mathbb{E}[q_{i}q_{i}^{\top}]$ and the score variance matrix $\Omega_{0}:=\mathbb{E}[q_{i}q_{i}^{\top}u_{i}^{2}]$. The influence function for the scalar parameter $\theta(x)=p(x)^{\top}\beta_{0}$ is given by $\psi_{i}(x)=p(x)^{\top}J_{0}^{-1}q_{i}u_{i}$. Consequently, the asymptotic sandwich variance is:
$\sigma^{2}(x)\;=\;p(x)^{\top}\,J_{0}^{-1}\,\Omega_{0}\,J_{0}^{-1}\,p(x).$
This extends the scalar DML variance formula ($J_{0}^{-2}\mathbb{E}[\psi^{2}]$) to the estimation of a function via projection. Using cross-fitting, we construct the plug-in estimators:
$\hat{J}\;=\;\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in I_{k}}q_{i}q_{i}^{\top},\qquad\hat{\Omega}\;=\;\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in I_{k}}q_{i}q_{i}^{\top}\hat{u}_{i}^{2},$
where $\hat{u}_{i}$ are the empirical residuals. The pointwise variance estimator at $x$ is:
$\hat{\sigma}^{2}(x)\;=\;p(x)^{\top}\,\hat{J}^{-1}\,\hat{\Omega}\,\hat{J}^{-1}\,p(x).$
Let $\hat{\sigma}(x)=\sqrt{\hat{\sigma}^{2}(x)}$. An approximate $(1-\alpha)$ pointwise confidence interval for the CME at $x$ is:
$\left[\,\hat{\theta}(x)\ \pm\ \Phi^{-1}(1-\alpha/2)\ \frac{\hat{\sigma}(x)}{\sqrt{N}}\,\right].$
To obtain a confidence band that is valid simultaneously over a set $\mathcal{X}$, we apply a Gaussian multiplier bootstrap (chernozhukov2013gaussian) to the projected influence function. Let $\{\xi_{i}^{\,b}\}_{i=1}^{N}$ be i.i.d. standard normal bootstrap weights, independent of the data. For each bootstrap draw $b=1,\dots,B$, we compute the process:
$t^{*,b}(x)\;=\;\frac{1}{\sqrt{N}}\sum_{k=1}^{K}\sum_{i\in I_{k}}\xi_{i}^{\,b}\,\frac{p(x)^{\top}\hat{J}^{-1}(q_{i}\hat{u}_{i})}{\hat{\sigma}(x)}.$
Let $c_{1-\alpha}$ be the empirical $(1-\alpha)$-quantile of the maximal statistic $\{\sup_{x\in\mathcal{X}}|t^{*,b}(x)|\}_{b=1}^{B}$. The resulting simultaneous $(1-\alpha)$ confidence band is:
$\left[\,\hat{\theta}(x)\ \pm\ c_{1-\alpha}\ \frac{\hat{\sigma}(x)}{\sqrt{N}}\,\right]\qquad\text{for all }x\in\mathcal{X}.$
4.5 Discrete Outcome
Finally, we explore the application of AIPW and DML estimators to models with discrete outcomes. Traditional parametric approaches for binary or count data typically rely on link functions—such as logit, probit, Poisson, or negative binomial—to capture nonlinearity between predictors and the outcome. This introduces two key challenges. First, it becomes difficult to disentangle the nonlinearity imposed by the link function from the intrinsic nonlinearity of the CME function $\theta(x)$. Second, when the treatment $D$ is continuous, the marginal effect $\frac{\partial\mathbb{E}[Y]}{\partial D}$ depends on both $X$ and $D$ through the link function derivative, requiring an additional averaging step to recover the CME.
When the treatment is binary, the link function is naturally handled by the nuisance learners (e.g., using logistic regression for the propensity score and outcome model). However, when the treatment is continuous, the standard DML framework relying on the PLRM does not explicitly model the link function. In this setting, DML serves as a linear approximation strategy—analogous to the Linear Probability Model—treating the composite structure $\mathbb{E}[Y|D,V]$ as a single nuisance component while preserving Neyman orthogonality for the target parameter.
Link Functions in Discrete Outcome Models
In parametric GLM approaches, the conditional mean $\mu=\mathbb{E}[Y\mid D,V]$ is connected to a predictor $m(D,V)$ via an inverse link function $g^{-1}(\cdot)$:
$\mathbb{E}[Y\mid D,V]=g^{-1}\bigl(m(D,V)\bigr).$
Common choices for the link function $g(\cdot)$ include:
  • Logit Link (Binary): $\mu=\frac{1}{1+\exp\{-m(D,V)\}}$.
  • Probit Link (Binary): $\mu=\Phi\bigl(m(D,V)\bigr)$, where $\Phi$ is the standard normal CDF.
  • Poisson/Log Link (Count): $\mu=\exp\{m(D,V)\}$.
Classically, the predictor is modeled as a linear interaction:
$m(D,X,Z)=\beta_{0}+\beta_{1}D+\beta_{2}X+\beta_{3}(D\times X)+\gamma^{\top}Z.$
Challenges in Estimating the CME
For a binary treatment, the CME is simply the difference in expectations, $\mathbb{E}[Y(1)-Y(0)\mid X]$. However, for a continuous treatment, the CME is defined as the expected partial derivative of the dose-response function. Under the composite model $\mathbb{E}[Y\mid D=d,V=v]=g^{-1}(m(d,v))$, the chain rule yields:
$\frac{\partial}{\partial d}\mathbb{E}[Y\mid d,v]\;=\;g^{-1\,\prime}\!\big(m(d,v)\big)\cdot\frac{\partial m(d,v)}{\partial d}.$
Using the linear specification for $m(\cdot)$ above, this becomes:
$\frac{\partial\mathbb{E}[Y]}{\partial d}=g^{-1\,\prime}\!\big(\beta_{0}+\beta_{1}d+\beta_{2}x+\beta_{3}dx+\dots\big)\cdot(\beta_{1}+\beta_{3}x).$
Consequently, the CME at $X=x$ requires averaging this derivative over the distribution of the treatment $D$ and other covariates $Z$:
$\theta(x)\;=\;\mathbb{E}\!\left[\,g^{-1\,\prime}\!\big(m(D,x,Z)\big)\cdot(\beta_{1}+\beta_{3}x)\ \Big|\ X=x\right].$
This highlights the complexity: the marginal effect depends on the baseline probability or count (via $g^{-1\,\prime}$), meaning that even with a linear latent predictor, the observed marginal effect varies with $D$.
DML Strategy for Discrete Outcomes
The DML framework handles these challenges by focusing on the nuisance components rather than the link function parameters themselves.
Binary Treatment: The target $\mathbb{E}[Y(1)-Y(0)\mid X]$ is estimable using the standard AIPW signal. We simply modify the nuisance learners:
  • For the propensity score $\pi(V)$, we use a classification learner (e.g., regularized logistic regression, random forest classifier).
  • For the outcome models $\mu(d,V)$, we use classification learners to ensure predictions $\hat{Y}\in[0,1]$.
Continuous Treatment: The focus shifts to $\mathbb{E}\left[\left.\frac{\partial Y}{\partial D}\right\rvert X=x\right]$. While the true DGP may be a GLM (e.g., Poisson), we apply the DML estimator based on the PLRM.
  • Approximation: By using the PLRM score, we effectively estimate the “best linear approximation” to the conditional mean function. This approach is often robust and interpretable, recovering the average partial effect without requiring the unstable estimation of density derivatives
  • Estimation: We use probabilistic classifiers (or regressions) to estimate the conditional mean $\mathbb{E}[Y\mid V]$ and regressors to estimate $\mathbb{E}[D\mid V]$. We then form the residuals $\tilde{Y}$ and $\tilde{D}$ and regress $\tilde{Y}$ on $\tilde{D}$, interacting with basis functions of $X$.
Note that although we can use classification algorithms to learn the nuisance $\mathbb{E}[Y\mid V]$ to respect outcome bounds, the final projection step remains a linear regression of residuals, as dictated by the orthogonality of the Robinson score.
4.6 Summary
This chapter introduces DML as a robust approach for estimating the CME in the presence of high-dimensional or complex nuisance functions. The central concept underlying DML is Neyman orthogonality, which ensures that small estimation errors in nuisance functions, such as outcome models or propensity scores, do not bias the CME estimate to first order. This property generalizes the double robustness of AIPW introduced in Chapter 3 and permits the use of flexible machine learning methods, such as neural networks (NN), random forests (RF), and histogram gradient boosting (HGB), to approximate complex nuisance relationships. While orthogonalization yields estimators that are insensitive to regularization bias, cross-fitting mitigates overfitting by estimating nuisance components and causal effects on separate data splits. Together, these strategies support valid inference at the $\sqrt{n}$- rate and can be extended to infinite-dimensional parameters via smoothing methods such as kernel and B-spline regressions.
We also discuss applications of DML in binary and continuous treatment settings. For binary treatments, DML constructs orthogonal signals using an AIPW-style formula with cross-fitting, isolating moderators from additional covariates. For continuous treatments, it adopts a partial linear regression framework, orthogonally residualizing the treatment and outcome with respect to confounders. With discrete outcome, link functions complicate CME interpretation, especially with continuous treatments, where marginal effects vary with both $X$ and $D$. DML and AIPW estimators sidestep explicit link modeling by learning the nuisance components flexibly and, with continuous treatments, recovering an orthogonal best linear approximation through residual-on-residual regression. We provide empirical examples using the interflex package, illustrating its implementation in political science.
Chapter 5. Monte Carlo Evidence
In prior discussions, we examined various methods for estimating the CME, including the linear interaction model, the semiparametric kernel estimator, the AIPW-Lasso estimator—by which we mean AIPW with basis expansion and double-selection Lasso—and the more recently developed DML methods. This chapter uses Monte Carlo simulations to compare the performance of these methods in a binary treatment setting. Specifically, we aim to assess the relative advantages of the kernel estimator, AIPW-Lasso estimator, and DML estimators in terms of accuracy and associated computational cost under different scenarios.
Our simulation results highlight that for DML to perform well, a sufficiently large sample size and the presence of rich covariates are essential. Without enough observations, its advantages—such as handling high-dimensional data and capturing complex relationships—may not fully materialize, and it can underperform compared to the kernel estimator. In contrast, AIPW-Lasso, which relies on simpler learners, performs well even with small to medium-sized samples but can underperform compared to the DML methods under very complex data-generating processes. Furthermore, while the kernel estimator can adeptly capture nonlinearities in the CME, AIPW-Lasso and DML offer additional strengths in addressing nonlinearities among the other covariates. Researchers are therefore encouraged to consider AIPW-Lasso and DML when working with large datasets and when nonlinear relationships extend beyond the moderator.
We also compare the performance of three DML learners—neural networks (NN), random forests (RF), and histogram gradient boosting (HGB)—using both default and fine-tuned hyperparameters through cross-validation and grid search. Although fine-tuning can improve model fit and CME estimates, it is time-consuming and may require prior knowledge to set up appropriate hyperparameter grids.
In the remainder of this chapter, we present four data generating processes (DGPs), from simple to complex. First, we use a simple DGP to compare the kernel estimator, AIPW-Lasso, and various DML estimators in a setting where covariates affect the outcome linearly. Second, we introduce nonlinear relationships between covariates and the outcome to assess how well these methods adapt. Then we consider a more complex DGP to evaluate the performance of NN, RF, and HGB under both default and tuned hyperparameters. Lastly, we introduce high-order interactions and discontinuities in a substantially more complex DGP to compare the performance of the AIPW-Lasso and DML with tuned hyperparameters. In the simulations, we find the loss functions from the outcome and treatment assignment models are good predictors of the estimator’s performance, so researchers can use these reported metrics to guide their choice of estimators.
5.1 Simulation Study 1: Linear Covariate Effects
In the first simulation study, the covariate affects the outcome linearly. In this setting, the kernel estimator (with an appropriately chosen bandwidth) can approximate the true marginal effect accurately. Our simulations show that, with sufficiently large sample sizes, DML methods perform comparably to the kernel estimator with cross-validated bandwidth. However, for smaller samples, DML methods can be more sensitive to noise, often resulting in underfitting of their machine learning models.
To properly evaluate these methods, we define the outcome $Y_{i}$ as
$Y_{i}=\theta(X_{i})D_{i}+g(V_{i})+\varepsilon,$
where $\theta(X_{i})$ is the CME of $D_{i}$ on $Y_{i}$, $g(V_{i})=\mathbb{E}[Y_{i}\mid D_{i}=0,V]$ is the model for the untreated counterfactual outcome, and $\varepsilon\sim N(0,1)$ is the i.i.d. error term. The covariates $V_{i}=[X_{i},Z_{i}]$ are i.i.d drawn from a uniform distribution: $X_{i},Z_{i}$ are i.i.d drawn from $\textrm{Unif}(-\sqrt{3},\sqrt{3})$. The CME is quadratic in $x$: $\theta(x)=x^{2}$, and $V_{i}$ enters the outcome model $g(V_{i})$ linearly: $g(V_{i})=1+X_{i}+0.5Z_{i}$.
The propensity score $\Pr[D_{i}=1\mid V_{i}]$ follows a logistic model:
$\pi(V_{i})=\Pr[D_{i}=1\mid V_{i}]=\frac{\exp(0.5X_{i}+0.5Z_{i})}{1+\exp(0.5X_{i}+0.5Z_{i})}=\textrm{logit}^{-1}(0.5X+0.5Z),$
and the treatment variable $D_{i}$ is sampled from $D_{i}\sim\text{Bernoulli}\bigl(\pi(V_{i})\bigr).$
Because the CME is nonlinear in $X_{i}$, the linear interaction model is biased. Recall that the kernel estimator is specified as follows:
$Y_{i}=f(X_{i})+h(X_{i})D_{i}+Z_{i}^{\top}\gamma(X_{i})+\epsilon_{i}$
With a properly selected bandwidth, it can approximate the DGP effectively with $f(X_{i})=1+X_{i}$, $h(X_{i})=X_{i}^{2}$, and $\gamma(X_{i})=0.5$. Bandwidth selection via cross-validation, which targets out-of-sample predictive accuracy, helps to find an optimal balance between bias and variance.
The AIPW-Lasso and DML estimators both model the outcome and the propensity score as functions of $V$:
$\eta=\Bigl\{\,\underbrace{\mathbb{E}(Y_{i}\mid D_{i}=1,V)}_{\mu(1,V_{i})},\,\underbrace{\mathbb{E}(Y_{i}\mid D_{i}=0,V)}_{\mu(0,V_{i})},\,\underbrace{\mathbb{E}(D_{i}\mid V_{i})}_{\pi(V_{i})}\Bigr\},$
and construct the orthogonal signal of the outcome as
$\Lambda_{i}(\hat{\eta})=\hat{\mu}(1,V_{i})\;-\;\hat{\mu}(0,V_{i})\;+\;\frac{D_{i}\,[Y_{i}-\hat{\mu}(1,V_{i})]}{\hat{\pi}(V_{i})}\;-\;\frac{(1-D_{i})\,[Y_{i}-\hat{\mu}(0,V_{i})]}{1-\hat{\pi}(V_{i})}.$
With this DGP, the true nuisance functions are $\mu_{i}(1,V_{i})=X_{i}^{2}+1+X_{i}+0.5Z_{i}$, $\mu(0,V_{i})=1+X_{i}+0.5Z_{i}$, and $\pi(V_{i})=\textrm{logit}^{-1}(0.5X_{i}+0.5Z_{i})$. Although these functions are not complex, small samples can still challenge machine learning model fitting, leading to underfitting or convergence issues. Moreover, the cross-fitting procedure in DML—which reduces the training data in each fold—may further amplify these issues when data are limited.
In Figure 17(f), we compare the estimated CME to the true CME for a single simulation from the DGP. The first row presents model-based approaches, including the linear estimator (based on the linear interaction model) and the kernel estimator (with bandwidth selected via cross-validation). The second row shows results from the AIPW-Lasso estimator and the DML-NN estimator for $n=1,000$. The third row presents the DML method using HGB, showing results for both $n=1,000$ and $n=10,000$. At the bottom of each panel, we display the distribution of treated units (in red) and control units (in gray) along the moderator $X$ using histograms.
Figure 17: Comparison of Estimators with DGP1
(a) Linear (n=1,000n=1{,}000)
(b) Kernel (n=1,000n=1{,}000)
(c) AIPW-Lasso (n=1,000n=1{,}000)
(d) DML-NN (n=1,000n=1{,}000)
(e) DML-HGB (n=1,000n=1{,}000)
(f) DML-HGB (n=10,000n=10{,}000)
Note: In each figure: the red dashed line represents the true CME; the black solid line represents the estimated CME; the shaded area represents the pointwise confidence intervals, while the dashed gray lines represent the uniform confidence intervals.
Figure 17(f) shows that the linear estimator exhibits substantial bias, while the kernel estimator closely approximates the true CME. AIPW-Lasso performs comparably to the kernel method, offering more stable estimation and narrower confidence intervals. DML-NN yields similar point estimates but with slightly wider confidence intervals. At n=1,000n=1{,}000, DML-HGB performs poorly, with uniform confidence intervals that nearly always cover zero. However, at n=10,000n=10{,}000, its performance improves markedly, and the estimated CME closely aligns with the truth. DML-RF, whose results are omitted here for brevity, performs similarly to DML-HGB. These results suggest that both kernel and AIPW-Lasso estimators are well suited for estimating nonlinear CMEs when covariates enter the outcome model linearly and the sample size is moderate.
To evaluate the performance of each method in terms of accuracy and computational efficiency as the sample size increases, we compare two metrics in Figure 18: the root mean squared error (RMSE) and execution time. RMSE measures the discrepancy between the true CME, $\theta(X_{i})$, and its estimate, $\hat{\theta}(X_{i})$, weighted by the probability density of $X_{i}$, $f_{X}(X_{i})$. Since $X_{i}$ is uniformly distributed in this simulation, the weights are constant over $X$. Execution time serves as a practical metric for computational efficiency.
Figure 18 presents results for the linear estimator, the kernel estimator, the AIPW-Lasso estimator, and DML methods using three different learners (NN, RF, and HGB). As expected, the linear estimator has the highest RMSE across all sample sizes due to substantial bias. The kernel estimator achieves relatively low RMSE but incurs higher computational costs because of bandwidth selection via cross-validation. The AIPW-Lasso estimator performs slightly better in RMSE and is generally less computationally demanding than the kernel method in small samples, although its runtime increases rapidly as the sample size grows. Among the DML estimators, DML-NN performs comparably to AIPW-Lasso and becomes more efficient in runtime when n>5,000n>5{,}000. HGB achieves low RMSE once $n$ exceeds 10,000, but its execution time grows substantially for larger $n$. Random forest performs worse than both the NN and HGB in this setting. The kernel estimator is computationally more expensive than AIPW-Lasso and untuned DML methods, even in small samples, due to cross-validation.
Figure 18: Performance of Different Methods: DGP1
(a) RMSE
(b) Execution time
Notes: The above figures show how root mean squared error (left) and execution time (right) change as the sample size increases. Both axes in each figure are in log scale.
With this DGP, the only source of nonlinearity is in the CME, while the covariate $Z_{i}$ enters the outcome model linearly. Under these conditions, the kernel, AIPW-Lasso, and DML estimators can effectively capture the quadratic CME when the sample size is sufficiently large. However, with smaller samples, DML methods may suffer from instability and high variance, leading to worse performance. Therefore, when the sample size is small or moderate and computational efficiency is a key concern, AIPW-Lasso offers the best trade-off between accuracy and runtime. As the dataset grows, both AIPW-Lasso and DML methods become increasingly competitive in accuracy, albeit with greater computational cost.
5.2 Simulation Study 2: Nonlinear Covariate Effects
We use the second simulation study to illustrate the advantages of the AIPW-Lasso and DML methods when covariates enter the outcome model nonlinearly. Based on the first simulation study, we modify only the outcome model as follows:
$g(V_{i})=1+X_{i}+e^{2Z_{i}+2}.$
The CME, $\theta(X_{i})$, the propensity score, $\pi(V_{i})$, and the distributions of $X$ and $Z_{i}$ remain unchanged. While the kernel estimator can effectively capture nonlinearity in the moderator itself, it struggles to accommodate nonlinearities in other covariates. Specifically, the kernel estimator restricts $Z_{i}$ to enter linearly via $\gamma(X_{i})$, making it incapable of modeling terms such as $e^{2Z_{i}+2}$. By contrast, the AIPW-Lasso and DML methods can flexibly capture more complex effects of $Z_{i}$ on $Y_{i}$ through basis expansion or machine learning algorithms.
Figure 19: Comparison of Estimators with DGP2
(a) Kernel (n=1,000n=1{,}000)
(b) Kernel (n=10,000n=10{,}000)
(c) AIPW-Lasso (n=1,000n=1{,}000)
(d) AIPW-Lasso (n=10,000n=10{,}000)
(e) DML-NN (n=1,000n=1{,}000)
(f) DML-NN (n=10,000n=10{,}000)
Note: In each figure: the red dashed line represents the true CME; the black solid line represents the estimated CME; the shaded area represents the pointwise confidence intervals, while the dashed gray lines represent the uniform confidence intervals. At the bottom of each figure, we display histograms showing the distribution of treated units (in red) and control units (in gray) along the moderator $X$.
In Figure 19, we compare the estimated CME with the true CME in a single simulated sample based on this DGP. The first row presents the kernel estimator (with bandwidth selected via cross-validation) for n=1,000n=1{,}000 and n=10,000n=10{,}000. The second row shows results from the AIPW-Lasso estimator for n=1,000n=1{,}000 and n=10,000n=10{,}000, using basis expansion and post-selection Lasso for variable selection. The third row presents results from DML-NN, with both the outcome and propensity score models fitted using NN (DML-RF and DML-HGB behave similarly).
Unlike the first scenario, the kernel estimator fails to approximate the CME accurately at either sample size. For n=1,000n=1{,}000, cross-validation selects a small bandwidth, resulting in highly noisy estimates. At n=10,000n=10{,}000, the selected bandwidth is much larger, causing the kernel estimator to resemble the biased linear estimator and still fail to capture the nonlinearity introduced by $e^{2Z_{i}+2}$. In contrast, the AIPW-Lasso estimator performs remarkably well, even at n=1,000n=1{,}000, likely because the exponential term is well approximated by the B-spline expansion of $Z_{i}$. The DML method with the NN learner exhibits high variance at smaller $n$, but its accuracy becomes comparable to that of AIPW-Lasso as $n$ increases.
Figure 20: Performance of Different Methods: DGP2
(a) RMSE
(b) Execution time
Notes: The above figures show how root mean squared error (left) and execution time (right) change as the sample size increases. Both axes in each figure are in log scale.
Figure 20 summarizes the RMSE and execution time of each estimator based on 200 simulation runs. We evaluate the linear estimator, the kernel estimator with cross-validation, the AIPW-Lasso estimator with basis expansion and post-selection Lasso, and DML methods using NN, RF, or HGB. Under this DGP, both the linear and kernel estimators exhibit substantial bias, even at large $n$. By contrast, DML estimators improve significantly after approximately 3,000 observations, with NN and HGB achieving particularly low RMSE, albeit with higher computational cost than RF. Notably, the AIPW-Lasso estimator consistently attains low RMSE (as seen in Figure 19) with moderate computational overhead when the sample size is not large, although its runtime increases rapidly as $n$ grows.
In short, compared with the first study, the introduction of nonlinearity in $g(V_{i})$ through $e^{2Z+2}$ renders the kernel estimator ineffective at capturing this structure, resulting in biased CME estimates. In contrast, AIPW-Lasso and DML methods can flexibly accommodate nonlinearities in additional covariates. Therefore, when nonlinearities extend beyond the moderator, kernel methods become less suitable, and AIPW-Lasso or DML are more likely to provide greater accuracy and, in some cases, more efficient computation. AIPW-Lasso offers a strong advantage in accuracy at smaller sample sizes and remains computationally competitive at moderate sample sizes.
5.3 Simulation Study 3: Fine-tuning DML
For many machine learning methods, model tuning is crucial to ensure strong performance. In this third simulation study, we compare how different machine learners—NN, RF, and HGB—perform under default hyperparameters versus hyperparameters fine-tuned via cross-validation, using a more complex DGP. Although default settings are convenient and sometimes adequate, they often fail to capture the underlying complexity of the data, resulting in underfitting or overfitting depending on the context. In DML, where both the outcome and propensity score models must be estimated, it is especially unlikely that a single set of hyperparameters will perform well for both. Cross-validation allows for more targeted tuning by optimizing out-of-sample predictive performance separately for each model component. By comparing default and cross-validated configurations side by side, we highlight the trade-off between computational cost—since cross-validation can be time-intensive—and the potential for improved accuracy through more careful tuning.
In this setting, we consider a nonlinear outcome model and a nonlinear propensity score model, each depending on a single moderator $X$ and four additional covariates $Z_{1i}$, $Z_{2i}$, $Z_{3i}$, and $Z_{4i}$. Specifically, the outcome variable $Y_{i}$ is then given by
$Y_{i}=\theta(X_{i})\,D_{i}+g(V_{i})+\epsilon_{i},$
in which $D_{i}\sim\operatorname{Bernoulli}(\pi(V_{i}))$ and $\epsilon_{i}\sim\mathcal{N}(0,1)$ and
$g(V_{i})=$ $1+X_{i}+\exp(2X_{i}+2)+3\,\sin(X_{i}+Z_{i1})+2X_{i}Z_{i2}$
[…]
The propensity score model is given by:
$\pi(V_{i})=\text{logit}^{-1}\{0.25+X_{i}-X_{i}^{2}-Z_{1i}^{2}+2X_{i}Z_{1i}Z_{2i}-\text{logit}^{-1}(Z_{1i}+Z_{3i})+2Z_{2i}^{2}\cos(Z_{4i})\}$
This highly nonlinear specification poses a challenge for all estimation methods and underscores the importance of hyperparameter tuning in flexible learners. As in previous studies, the true CME is quadratic: $\theta(X_{i})=X_{i}^{2}$.
Tuning NN.
Our NN model begins with the default parameter settings in scikit-learn’s MLPRegressor and MLPClassifier, which include a single hidden layer of 100 neurons, ReLU activation, the Adam optimizer, an L2 regularization term (alpha) of 0.0001, and a maximum of 1,000 training iterations. To allow for greater flexibility, we focus our hyperparameter search on four key components: layer architecture, activation function, solver, and regularization strength. Specifically, we vary the hidden layer configurations among (50,), (100,), and (50, 50), allowing for both simpler single-layer networks and deeper architectures capable of capturing more complex relationships. We compare the tanh and relu activation functions to account for different gradient behaviors and consider both stochastic gradient descent (sgd) and Adam (adam) as solvers, given their distinct convergence properties. Finally, we tune the L2 penalty term alpha over $\{10^{-4},10^{-3},10^{-2}\}$, balancing the trade-off between bias and variance. Although a broader grid could yield further refinements, we limit the search space to ensure adequate coverage of key parameters while controlling the computational burden of cross-validation in repeated simulation runs.
The code snippet below demonstrates how to implement this hyperparameter tuning using the interflex package:
1param_grid_nn <- list(
2 hidden_layer_sizes = list(
3 list(50L),
4 list(100L),
5 list(50L, 50L)
6 ),
7 activation = c("tanh", "relu"),
8 solver = c("sgd", "adam"),
9 alpha = c(1e-4, 1e-3, 1e-2)
10)
11
12sol.nn.cv <- interflex(data = sim_data, Y="Y", D="D", X="X", Z=Z.list, estimator="DML", CV = TRUE, model.t="nn", model.y="nn",param.grid.y = param_grid_nn, param.grid.t = param_grid_nn)
Tuning RF.
Our random forest model uses the default parameter settings in scikit-learn’s RandomForestRegressor and RandomForestClassifier, which include 100 estimators (n_estimators = 100) and unrestricted tree depth. In our grid search, we broaden the search space by allowing n_estimators to be 100 or 300, striking a balance between computational overhead and variance reduction through ensemble averaging. We also vary max_depth among {NULL, 5, 10} so that the model can adapt its complexity based on data size and noise levels. To control the formation of splits, we choose min_samples_split in {2, 10} and min_samples_leaf in {1, 5}, thereby mitigating the risk of overfitting by requiring more observations in node splits and leaves. Furthermore, we compare max_features set to 1.0 (all features considered) versus 0.8 (a subsample of features at each split) to foster diversity among the trees. Finally, we explore both bootstrap = TRUE (standard bootstrap sampling) and FALSE (sampling without replacement) to gauge their effect on variance reduction.
Tuning HGB.
Our HGB model starts from the default parameter settings in scikit-learn’s HistGradientBoostingRegressor and HistGradientBoostingClassifier, which use a learning rate of 0.1 and a maximum of 100 boosting iterations. We tune several other key hyperparameters to enhance flexibility while avoiding overfitting. Specifically, we let learning_rate range over {0.01, 0.1}, spanning conservative to more aggressive updates, and allow up to either 100 or 200 boosting iterations (max_iter). We examine two distinct values for max_leaf_nodes (31 and 127) to regulate tree complexity, and vary max_depth among {NULL, 3, 5} to impose a bound on tree depth only when potentially beneficial. In addition, min_samples_leaf takes values {5, 20} to guard against overfitting by avoiding extremely small leaves, while l2_regularization in {0.0, 1.0} manages the trade-off between bias and variance. Finally, we tune max_features in {1.0, 0.8}, controlling the fraction of features available at each split.
In Figure 21(h), we compare the estimated CME to the true CME under two different hyperparameter settings (default vs. fine-tuned). Panels (b), (c), and (d) display estimates obtained using each learner’s default hyperparameters, while panels (f), (g), and (h) show estimates based on cross-validated hyperparameters. We also show the estimates given by the kernel estimator in (a), and the AIPW-Lasso estimator in (e), for comparison.
Figure 21: Estimated CME based on a single simulation following the Third DGP
(a) Kernel
(b) NN (default)
(c) HGB (default)
(d) RF (default)
(e) AIPW-Lasso
(f) NN (tuned)
(g) HGB (tuned)
(h) RF (tuned)
Note: In each subfigure, the black solid line represents the estimated CME, while the red dashed line denotes the true CME. The shaded gray areas illustrate the pointwise confidence intervals, and the gray dashed lines depict the uniform confidence intervals. At the bottom of each subplot, a histogram displays the distribution of treated and control units across varying values of the moderator $X$.
A key practical question is how to select the best learner and hyperparameter set without knowing the true CME. As shown at the top of each panel in Figure 21(h), the cross-fitted loss functions for the nuisance models, $\mu_{1}$, $\mu_{0}$, and $\pi$, provide a useful diagnostic. Lower loss—root mean squared error for $\mu$ or cross-entropy for $\pi$—indicates better out-of-sample fit for the corresponding nuisance component. [^9] For the AIPW-Lasso estimator, we use out-of-sample fitting errors from the linear outcome regression and the logistic treatment assignment model. Although not a guarantee, learners with lower average nuisance losses are generally expected to yield more accurate and reliable CME estimates. This pattern is evident in the figure: tuned models, which perform better visually, also report lower nuisance losses than their default counterparts (e.g., panel (b) versus (f)).
Notably, the NN learner highlights the importance of tuning hyperparameters to the data-generating process. For the outcome model, the selected hyperparameters are activation = tanh, alpha = 1e-4, hidden_layer_sizes = (50,50), and solver = adam; in contrast, the propensity score model performs best with activation = tanh, alpha = 0.01, solver = sgd, and hidden_layer_sizes = (50,50). This divergence in optimal specifications underscores the value of cross-validation, as the outcome and propensity score functions may differ in complexity and thus require different degrees of flexibility. Moreover, comparing subfigure (b) (default hyperparameters) with subfigure (f) (cross-validated hyperparameters) shows that DML with NN learners yields more accurate CME estimates when hyperparameters are properly tuned rather than set to defaults.
In contrast to the NN results, the HGB learner exhibits only modest improvement when moving from default hyperparameters to those selected via cross-validation. For the outcome model, the chosen hyperparameters are $\texttt{l2\_regularization}=1$, $\texttt{learning\_rate}=0.1$, $\texttt{max\_depth}=5$, $\texttt{max\_features}=1$, $\texttt{max\_iter}=200$, $\texttt{max\_leaf\_nodes}=31$, and $\texttt{min\_samples\_leaf}=20$. The propensity score model retains the same values for the penalty term, learning rate, and minimum samples per leaf but uses $\texttt{max\_leaf\_nodes}=127$ and $\texttt{max\_iter}=100$. Although these configurations differ meaningfully—particularly in tree size—between the outcome and propensity models, the resulting CME estimates show only a slight numerical improvement over the default settings, as illustrated in Figure 21(h),(c) and (g). This pattern suggests that the default HGB parameters already capture the underlying signal reasonably well, with cross-validation providing incremental rather than substantial gains.
Using a similar procedure for the RF learner, cross-validation selects bootstrap = TRUE, max_depth = 10, max_features = 0.8, min_samples_leaf = 1, min_samples_split = 2, and n_estimators = 300 for the outcome model, while the propensity score model uses the same values except for min_samples_leaf = 5. These settings imply slightly deeper trees, a broader feature subset, and more trees than the defaults, which may improve variance reduction. However, comparing Figure 21(h),(d) and (h), these adjustments do not yield noticeable improvements in the estimated CME, and overall performance remains inferior to both the NN and HGB learners. This result suggests that the random forest may be less well suited to capturing the nonlinear structure in this DGP, or that more specialized tuning, or larger sample sizes, are required to achieve performance comparable to the other methods.
In Figure 22, we compare RMSE, loss functions, and execution time across the kernel estimator, AIPW-Lasso, and DML with different ML learners. The comparison includes both default hyperparameters and those selected via cross-validation, based on 100 simulations of the previously described DGP.
Figure 22: Performance of Different Methods: DGP3
(a) RMSE
(b) Execution time
(c) Loss Function (Outcome with $D=1$)
(d) Loss Function (Treatment)
Notes: The above figures show how RMSE (top-left), execution time (top-right), loss of the outcome model for treated units (bottom-left), and loss of the treatment model (bottom-right) change as the sample size increases. Both axes in each figure are in log scale.
We observe the following patterns. First, all DML methods outperform the kernel estimator once the sample size $n$ is sufficiently large. Second, for NN and HGB, cross-validating hyperparameters consistently yields lower RMSE, especially when the sample size is modest. This likely reflects that fine-tuning mitigates overfitting in moderate samples. By contrast, RF shows little improvement from cross-validation, suggesting that gains in out-of-sample predictive performance do not necessarily translate into better CME estimates. Third, the AIPW-Lasso estimator performs competitively: it outperforms the kernel estimator even at moderate sample sizes and only slightly underperforms the tuned NN and HGB DML methods when $n$ is very large. Finally, the loss function plots (bottom row) tell a similar story. For all methods, outcome and treatment losses decline as the sample size increases, and fine-tuning hyperparameters (e.g., comparing ‘nn” to ‘nn-cv”) reduces loss in line with the RMSE improvements.
These plots also show that AIPW-Lasso often achieves lower loss than DML methods at small to moderate sample sizes. Finally, cross-validation substantially increases execution time, highlighting a trade-off between computational cost and improved model fit.
5.4 Simulation Study 4: High-Order Interactions and Discontinuities
In our final simulation study, we evaluate the estimators under a substantially more complex data-generating process. This DGP is specifically designed to challenge the limits of each method by incorporating two particularly difficult features: (1) sharp discontinuities (piecewise-defined regions) in the outcome model and (2) high-order interactions in both the outcome and propensity score models.
These features are known to be difficult for estimators that rely on smooth approximation functions, such as the B-spline basis expansion used in our AIPW-Lasso implementation. In theory, tree-based DML learners like RF and HGB should be far more adept at capturing these structures, as their core algorithm is based on partitioning the covariate space. This simulation puts that theory to the test.
As in the third study, we simulate data with $p=5$ covariates ( $X$ and four $Z$ variables), but we set $n=10,000$ and draw all covariates from a $\operatorname{Unif}(-\sqrt{3},\sqrt{3})$ distribution. The true CME is kept simple, following the quadratic form $\theta(x)=x^{2}$.
The complexity is concentrated entirely in the nuisance functions. The baseline outcome model, $g(V_{i})$, is constructed as a “rule-based” piecewise function, which creates sharp discontinuities. The model starts with a highly nonlinear and discontinuous base function, $g_{base}(V_{i})=1+X_{i}+f_{2}(Z_{i1})$, where $f_{2}$ is a piecewise function with three different cutoffs. The model then uses “if-then” logic to replace this base function in different regions of the covariate space. For example, one functional form (involving $X_{i}^{2}$, $Z_{i1}Z_{i2}$, and $\sin(X_{i}+Z_{i1})$) is used if $Z_{1}<1\text{ and }Z_{2}<0$, while a completely different function (involving $\text{abs}(X_{i}Z_{i1})$ and $Z_{i2}^{3}$) is used if $Z_{1}<1\text{ and }Z_{2}\geq 0$. This process creates a complex patchwork of different nonlinear surfaces with hard breaks at the region boundaries. The propensity score model, $\pi(V_{i})$, is similarly built by starting with a nonlinear baseline function (e.g., $X_{i}-X_{i}^{2}+X_{i}Z_{i1}-2\sin(X_{i}+Z_{i1})$) and then adding other high-order interaction terms (like $\cos(2X_{i})$ or $X_{i}Z_{i1}Z_{i2}$) only if observations fall into specific regions defined by $Z_{1}$ and $Z_{2}$.
In Figure 23(d), we compare the performance of the AIPW-Lasso and the three DML methods with fine-tuned hyperparameters. We find that AIPW-Lasso, despite its strong performance in earlier, smoother settings, is unable to approximate the complex nuisance functions in this DGP and produces heavily biased CME estimates.
Interestingly, DML-NN also performs poorly. This suggests that standard neural network architectures, including the $(50,50)$ specification in the candidate pool, are too shallow to capture the sharp discontinuities of this DGP. This failure is reminiscent of the classic XOR problem, where even simple nonlinear boundaries require sufficient network depth. Our DGP poses a substantially more demanding challenge, with multiple “cliffs” and high-order interactions. Neural networks are biased toward learning smooth functions and would likely require much deeper or more specialized architectures to approximate such highly piecewise structures.
In contrast, the tree-based DML learners—RF and HGB—perform markedly better. Their ability to partition the covariate space aligns well with the piecewise nature of the true functions, yielding substantially more accurate and reliable CME estimates.
Figure 23: Estimated CME based on a single simulation following the Fourth DGP
(a) NN (tuned)
(b) AIPW-Lasso
(c) HGB (tuned)
(d) RF (tuned)
Note: In each subfigure, the black solid line represents the estimated CME, while the red dashed line denotes the true CME. The shaded gray areas illustrate the pointwise confidence intervals, and the gray dashed lines depict the uniform confidence intervals. At the bottom of each subplot, a histogram displays the distribution of treated and control units across varying values of the moderator $X$.
This pattern is confirmed in Figure 24, which compares RMSE and nuisance loss functions across methods. In terms of RMSE, RF and HGB outperform AIPW-Lasso even at moderate sample sizes. The loss function plots clarify the source of this difference: the outcome model loss for AIPW-Lasso is substantially larger than that of RF and HGB and remains nearly flat as the sample size increases. This indicates that the B-spline basis expansion fails to capture the sharp, piecewise complexity of this DGP, regardless of sample size.
Figure 24: Performance of Different Methods: DGP4
(a) RMSE
(b) Loss Function (Outcome with $D=1$)
Notes: The above figures show how RMSE (left) and loss of the outcome model for treated units (right) change as the sample size increases. Both axes in each figure are in log scale.
5.5 Summary
Several key patterns emerge across the four simulation studies and provide practical guidance for researchers.
First, the semiparametric kernel estimator, while effective at capturing nonlinearities in the moderator (DGP1), is not robust. It fails once other covariates enter the outcome model nonlinearly (DGP2). By contrast, the AIPW-Lasso estimator, using a B-spline basis expansion, serves as a much stronger general-purpose benchmark. It performs well in both simple (DGP1) and smoothly complex (DGP2) settings, combining good accuracy with moderate computational cost, especially in small to medium samples.
Second, the limits of smooth approximation methods are clear. In the most complex scenario (DGP4), characterized by sharp discontinuities and piecewise structure, the B-spline basis in AIPW-Lasso fails entirely and yields heavily biased estimates. This stress test shows that when the data-generating process contains “cliffs” or rule-based logic in the covariate space, DML methods are not merely optional but necessary.
Third, the choice of DML learner matters and should align with the form of complexity. In the discontinuous setting (DGP4), DML-NN also performs poorly, as its architecture, even after tuning, struggles to capture sharp boundaries. In contrast, tree-based learners (RF and HGB), whose partitioning mechanisms are naturally suited to piecewise functions, perform substantially better.
Finally, the simulations yield two practical takeaways. First, fine-tuning is essential in complex settings. While default hyperparameters may suffice for simple problems, DGP3 shows that cross-validated tuning is critical for maximizing DML performance. Second, nuisance loss functions are an important diagnostic. Because the true CME is unknown in practice, we find that out-of-sample losses (e.g., RMSE or cross-entropy) for the outcome and treatment models strongly predict the accuracy of CME estimates. Researchers should use these losses to guide the choice between estimators (e.g., AIPW-Lasso versus DML, or among DML estimators) and to assess whether hyperparameter tuning has meaningfully improved model fit.
Chapter 6. Conclusion
Understanding conditional treatment effects—how the effect of a treatment $D$ on an outcome $Y$ varies with a moderating variable XX—is essential in social science research. Despite significant advances, current methods face notable challenges, including unclear estimands, misinterpretation of statistical results, lack of sufficient overlap in empirical data, rigid assumptions on functional forms, and difficulty with complex settings like discrete outcomes. This Element introduces a robust methodological framework that overcomes these limitations and ensures valid statistical inference for the key estimand, the conditional marginal effects (CME).
The linear interaction model is still widely used in empirical research. Despite its popularity, this model often suffers from inadequate common support, misspecifications, and multiple testing issues. To mitigate these issues, in Chapter 2, we advocate for diagnostic tools such as inspecting raw data and the binning estimator. Moreover, we discuss methods such as uniform confidence intervals via bootstrapping to address multiple comparisons and improve inferential validity. The semiparametric kernel estimator proposed by hainmueller2019much relaxes functional form assumptions. We further improve its robustness by incorporating adaptive kernels and fully moderated specifications.
In Chapter 3, we advance beyond classical approaches by introducing augmented inverse propensity score weighting (AIPW). The AIPW estimator for the CME combines outcome modeling and propensity score weighting, providing double robustness—maintaining consistency if either the propensity score model or outcome model is correctly specified. It takes three steps: (i) estimating the outcome and propensity score model, (ii) constructing the signals for treatment effects; (iii) smoothing over the moderator $X$. Compared with inverse propensity score weighting (IPW) alone, AIPW improves estimation stability and reduces variance. To accommodate high-dimensional covariates and relax functional form assumptions, we further introduce basis expansions and Lasso regularization techniques, which significantly stabilize estimation and improve precision. Finally, we extend AIPW to continuous treatments through partialing-out Lasso (PO-Lasso).
In Chapter 4, we present double/debiased machine learning (DML) as a powerful extension suitable for high-dimensional and complex nuisance parameters. DML achieves robust estimation through Neyman orthogonality, which ensures that small estimation errors in nuisance functions do not bias CME estimates to first order. This framework extends the double robustness of AIPW-Lasso and PO-Lasso by leveraging flexible machine learning methods, such as neural networks, random forests, and gradient boosting, to model complex relationships between covariates and the outcome and treatment. The combination of orthogonalization, which insulates estimators from regularization bias, and cross-fitting, which prevents overfitting by splitting the data, ensures valid inference at a $\sqrt{n}$-convergence rate. We illustrate DML for binary and continuous treatments with empirical examples from political science, including orthogonal signal construction, residualization strategies, and application of the interflex package. We also show that DML can accommodate empirical applications with discrete outcomes.
Moreover, in Chapter 5, we provide Monte Carlo evidence comparing kernel-based methods, AIPW-Lasso, and DML estimators. Simulation studies indicate that while kernel estimators effectively handle simple nonlinearities, they falter with complex covariate dependencies. In contrast, AIPW-Lasso and DML successfully model intricate nonlinearities, particularly in higher-dimensional settings, though their performance crucially depends on sample size and effective hyperparameter tuning. Tuning, while computationally costly, often significantly improves accuracy of the DML estimators, especially with complex learners like neural networks. However, simpler learners such as random forests show only marginal gains or even slight declines after tuning.
Finally, we offer the following practical recommendations for empirical researchers interested in estimating and interpreting conditional relationships:
  • Clearly define the quantity of interest, e.g., the CME, and state key assumptions—such as unconfoundedness, overlap, and functional form—before proceeding with the analysis.
  • Assess data quality by checking for missing values and outliers, and evaluate the overlap assumption. If overlap is violated, consider trimming and/or clipping the data to improve balance.
  • If the linear interaction model must be used for estimating the CME—for transparency, ease of computation and interpretation, or other reasons—supplement it with diagnostic tests for the rigid functional form, such as the binning estimator.
  • Use flexible modeling strategies to estimate the CME.
    • For experimental data, the kernel estimator is typically sufficient. Efficiency may be improved by incorporating basis expansions and post-Lasso selection.
    • For observational data with small to medium samples, we recommend Lasso-based methods: AIPW-Lasso for binary treatments and PO-Lasso for continuous treatments.
    • For observational data with large samples and rich covariates, DML estimators are more appropriate. When using complex machine learners such as neural networks, random forest, and histogram gradient boosting, apply cross-validation to tune hyperparameters and reduce overfitting.
  • Visualize estimated CME to facilitate intuitive understanding of effect heterogeneity.
  • Interpret results with care. Specifically, (i) avoid overusing causal language with respect to the moderator—CME captures the effect of $D$ along levels of the moderator $X$, but it does not represent the causal effect of $X$ itself. (ii) When making inferences about the sign or magnitude of the CME over an interval, use uniform confidence intervals to ensure proper coverage.
All the estimation strategies and diagnostic tools discussed in this Element can be implemented using the interflex package in R.
Appendix A Appendix
A.1 Chapter 1 Proofs
We show that the Stratified Difference-in-Means (SDIM) estimator and the IPW estimator of the CME are numerically equivalent when the covariate vector $V_{i}$ is discrete and the propensity score is estimated by within-cell frequencies: $\hat{\theta}_{\mathrm{SDIM}}(x)=\hat{\theta}_{\mathrm{IPW}}(x)$.
Proof. Fix the moderator value $x$. Let $V_{i}=(X_{i},Z_{i})$. Since $X_{i}$ is fixed to $x$, the variation in $V_{i}$ within this subpopulation is determined entirely by $Z_{i}$. Let $\mathcal{Z}_{x}$ denote the set of observed values of $Z$ for units with $X_{i}=x$. For each $z\in\mathcal{Z}_{x}$, define the cell $v=(x,z)$.
Define cell counts: $n_{1v}:=\#\{i:X_{i}=x,Z_{i}=z,D_{i}=1\}$, $n_{0v}:=\#\{i:X_{i}=x,Z_{i}=z,D_{i}=0\}$, and $n_{v}:=n_{1v}+n_{0v}$. Let $N_{x}:=\#\{i:X_{i}=x\}=\sum_{z\in\mathcal{Z}_{x}}n_{v}$. We assume strict overlap in the sample: for every $v$ such that $n_{v}>0$, we have $n_{1v}>0$ and $n_{0v}>0$.
Define within-cell sample means:
$\hat{\mu}^{1}(v)=\frac{1}{n_{1v}}\sum_{i\in I_{1v}}Y_{i},\qquad\hat{\mu}^{0}(v)=\frac{1}{n_{0v}}\sum_{i\in I_{0v}}Y_{i},$
where $I_{dv}=\{i:X_{i}=x,V_{i}=v,D_{i}=d\}$. The SDIM estimator is:
$\hat{\theta}_{\mathrm{SDIM}}(x)=\sum_{z\in\mathcal{Z}_{x}}\frac{n_{v}}{N_{x}}\bigl(\hat{\mu}^{1}(v)-\hat{\mu}^{0}(v)\bigr).$
The IPW estimator for the subpopulation $X=x$ is:
$\hat{\theta}_{\mathrm{IPW}}(x)=\frac{1}{N_{x}}\sum_{i:X_{i}=x}\left[\frac{D_{i}Y_{i}}{\hat{\pi}(V_{i})}-\frac{(1-D_{i})Y_{i}}{1-\hat{\pi}(V_{i})}\right].$
With discrete covariates, the nonparametric maximum likelihood estimator for the propensity score is the sample frequency: $\hat{\pi}(v)=n_{1v}/n_{v}$. Substituting this into the IPW formula and grouping terms by cell $v$:
$\hat{\theta}_{\mathrm{IPW}}(x)=\frac{1}{N_{x}}\sum_{z\in\mathcal{Z}_{x}}\left[\sum_{i\in I_{1v}}\frac{Y_{i}}{n_{1v}/n_{v}}-\sum_{i\in I_{0v}}\frac{Y_{i}}{1-n_{1v}/n_{v}}\right].$
Simplifying the fractions:
$\frac{1}{n_{1v}/n_{v}}=\frac{n_{v}}{n_{1v}}\quad\text{and}\quad\frac{1}{1-n_{1v}/n_{v}}=\frac{1}{n_{0v}/n_{v}}=\frac{n_{v}}{n_{0v}}.$
Thus, the inner sums become:
$\sum_{i\in I_{1v}}Y_{i}\left(\frac{n_{v}}{n_{1v}}\right)=n_{v}\left(\frac{1}{n_{1v}}\sum_{i\in I_{1v}}Y_{i}\right)=n_{v}\hat{\mu}^{1}(v),$
$\sum_{i\in I_{0v}}Y_{i}\left(\frac{n_{v}}{n_{0v}}\right)=n_{v}\left(\frac{1}{n_{0v}}\sum_{i\in I_{0v}}Y_{i}\right)=n_{v}\hat{\mu}^{0}(v).$
Substituting these back into the expression for $\hat{\theta}_{\mathrm{IPW}}(x)$:
$\hat{\theta}_{\mathrm{IPW}}(x)=\frac{1}{N_{x}}\sum_{z\in\mathcal{Z}_{x}}\bigl(n_{v}\hat{\mu}^{1}(v)-n_{v}\hat{\mu}^{0}(v)\bigr)=\sum_{z\in\mathcal{Z}_{x}}\frac{n_{v}}{N_{x}}\bigl(\hat{\mu}^{1}(v)-\hat{\mu}^{0}(v)\bigr)=\hat{\theta}_{\mathrm{SDIM}}(x).$
A.2 Chapter 3 Proofs
Identification based on IPW
Remark 1
Consider a binary treatment $D_{i}\in\{0,1\}$ and covariates $V_{i}=(X_{i},Z_{i})$. Let $\pi(v)=\Pr(D_{i}=1\mid V_{i}=v)$. Assume (i) consistency/SUTVA: $Y_{i}=D_{i}Y_{i}(1)+(1-D_{i})Y_{i}(0)$, (ii) unconfoundedness: $(Y_{i}(0),Y_{i}(1))\perp D_{i}\mid V_{i}$, and (iii) overlap: $0<\pi(V_{i})<1$ a.s. Then for any $v$,
$\mathbb{E}\!\Bigl[\tfrac{D_{i}}{\pi(V_{i})}\,Y_{i}\,\big|\,V_{i}=v\Bigr]=\mathbb{E}[Y_{i}(1)\mid V_{i}=v],\qquad\mathbb{E}\!\Bigl[\tfrac{1-D_{i}}{1-\pi(V_{i})}\,Y_{i}\,\big|\,V_{i}=v\Bigr]=\mathbb{E}[Y_{i}(0)\mid V_{i}=v].$
Proof. Fix $v$. By overlap, $\pi(v)\in(0,1)$, and since $\pi(V_{i})$ is a function of $V_{i}$,
$\mathbb{E}\!\left[\frac{D_{i}}{\pi(V_{i})}Y_{i}\,\big|\,V_{i}=v\right]=\frac{1}{\pi(v)}\,\mathbb{E}[D_{i}Y_{i}\mid V_{i}=v].$
Moreover,
$\mathbb{E}[D_{i}Y_{i}\mid V_{i}=v]=\mathbb{E}[D_{i}Y_{i}\mid V_{i}=v,D_{i}=1]\Pr(D_{i}=1\mid V_{i}=v),$
because $D_{i}Y_{i}=0$ when $D_{i}=0$. When $D_{i}=1$, consistency implies $Y_{i}=Y_{i}(1)$, so
$\mathbb{E}[D_{i}Y_{i}\mid V_{i}=v,D_{i}=1]=\mathbb{E}[Y_{i}(1)\mid V_{i}=v,D_{i}=1].$
By unconfoundedness, $Y_{i}(1)\perp D_{i}\mid V_{i}$, hence $\mathbb{E}[Y_{i}(1)\mid V_{i}=v,D_{i}=1]=\mathbb{E}[Y_{i}(1)\mid V_{i}=v]$. Therefore,
$\mathbb{E}[D_{i}Y_{i}\mid V_{i}=v]=\mathbb{E}[Y_{i}(1)\mid V_{i}=v]\;\pi(v),$
and substituting back yields
$\mathbb{E}\!\left[\frac{D_{i}}{\pi(V_{i})}Y_{i}\,\big|\,V_{i}=v\right]=\mathbb{E}[Y_{i}(1)\mid V_{i}=v].$
The control case is analogous:
$\mathbb{E}\!\left[\frac{1-D_{i}}{1-\pi(V_{i})}Y_{i}\,\big|\,V_{i}=v\right]=\frac{1}{1-\pi(v)}\,\mathbb{E}[(1-D_{i})Y_{i}\mid V_{i}=v],$
and $(1-D_{i})Y_{i}=0$ when $D_{i}=1$. When $D_{i}=0$, consistency gives $Y_{i}=Y_{i}(0)$, so
$\mathbb{E}[(1-D_{i})Y_{i}\mid V_{i}=v]=\mathbb{E}[Y_{i}(0)\mid V_{i}=v,D_{i}=0]\,(1-\pi(v))=\mathbb{E}[Y_{i}(0)\mid V_{i}=v]\,(1-\pi(v)),$
where the last equality uses unconfoundedness. Dividing by $1-\pi(v)$ gives the result.
Identification based on AIPW
Remark 2
Define the augmented inverse-propensity weighting (AIPW) signal
$\Lambda(V_{i}):=\mu_{1}(V_{i})+\frac{D_{i}}{\pi(V_{i})}\bigl(Y_{i}-\mu_{1}(V_{i})\bigr)-\mu_{0}(V_{i})-\frac{1-D_{i}}{1-\pi(V_{i})}\bigl(Y_{i}-\mu_{0}(V_{i})\bigr).$
Assume consistency/SUTVA, unconfoundedness $(Y_{i}(0),Y_{i}(1))\perp D_{i}\mid V_{i}$, and overlap $0<\pi(V_{i})<1$ a.s. If either (i) the propensity score is correct, $\pi(V_{i})=\Pr(D_{i}=1\mid V_{i})$, or (ii) the outcome regressions are correct, $\mu_{d}(V_{i})=\mathbb{E}[Y_{i}\mid D_{i}=d,V_{i}]$ for $d\in\{0,1\}$, then
$\mathbb{E}\bigl[\Lambda(V_{i})\mid V_{i}=v\bigr]=\mathbb{E}\bigl[Y_{i}(1)-Y_{i}(0)\mid V_{i}=v\bigr].$
Proof.
Case 1: The outcome regressions are correct. Suppose $\mu_{d}(V_{i})=\mathbb{E}[Y_{i}\mid D_{i}=d,V_{i}]$ for $d\in\{0,1\}$. Then
$\mathbb{E}\bigl[Y_{i}-\mu_{d}(V_{i})\mid V_{i}=v,D_{i}=d\bigr]=0.$
Taking conditional expectations given $V_{i}=v$,
$\mathbb{E}[\Lambda(V_{i})\mid V_{i}=v]$ $=\mu_{1}(v)-\mu_{0}(v)+\frac{1}{\pi(v)}\mathbb{E}\!\left[D_{i}\bigl(Y_{i}-\mu_{1}(v)\bigr)\mid V_{i}=v\right]-\frac{1}{1-\pi(v)}\mathbb{E}\!\left[(1-D_{i})\bigl(Y_{i}-\mu_{0}(v)\bigr)\mid V_{i}=v\right].$
Moreover,
$\mathbb{E}\!\left[D_{i}\bigl(Y_{i}-\mu_{1}(v)\bigr)\mid V_{i}=v\right]=\Pr(D_{i}=1\mid V_{i}=v)\,\mathbb{E}[Y_{i}-\mu_{1}(v)\mid V_{i}=v,D_{i}=1]=\pi(v)\cdot 0=0,$
and similarly the control residual term is zero. Hence
$\mathbb{E}[\Lambda(V_{i})\mid V_{i}=v]=\mu_{1}(v)-\mu_{0}(v).$
By consistency, when $D_{i}=1$, $Y_{i}=Y_{i}(1)$, so $\mu_{1}(v)=\mathbb{E}[Y_{i}(1)\mid V_{i}=v,D_{i}=1]$, and by unconfoundedness $\mathbb{E}[Y_{i}(1)\mid V_{i}=v,D_{i}=1]=\mathbb{E}[Y_{i}(1)\mid V_{i}=v]$. Thus $\mu_{1}(v)=\mathbb{E}[Y_{i}(1)\mid V_{i}=v]$. Analogously, $\mu_{0}(v)=\mathbb{E}[Y_{i}(0)\mid V_{i}=v]$. Therefore,
$\mathbb{E}[\Lambda(V_{i})\mid V_{i}=v]=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid V_{i}=v].$
Case 2: The propensity score is correct. Assume $\pi(V_{i})=\Pr(D_{i}=1\mid V_{i})$. Condition on $V_{i}=v$ and use $\mathbb{E}[D_{i}\mid V_{i}=v]=\pi(v)$ and $\mathbb{E}[1-D_{i}\mid V_{i}=v]=1-\pi(v)$. First, the $\mu_{1}$ terms satisfy
$\mathbb{E}\!\left[\mu_{1}(v)-\frac{D_{i}}{\pi(v)}\mu_{1}(v)\,\Big|\,V_{i}=v\right]=\mu_{1}(v)\left(1-\frac{\mathbb{E}[D_{i}\mid V_{i}=v]}{\pi(v)}\right)=0,$
and similarly the $\mu_{0}$ terms satisfy
$\mathbb{E}\!\left[-\mu_{0}(v)+\frac{1-D_{i}}{1-\pi(v)}\mu_{0}(v)\,\Big|\,V_{i}=v\right]=0.$
Hence
$\mathbb{E}[\Lambda(V_{i})\mid V_{i}=v]=\mathbb{E}\!\left[\frac{D_{i}}{\pi(v)}Y_{i}-\frac{1-D_{i}}{1-\pi(v)}Y_{i}\,\Big|\,V_{i}=v\right].$
By the IPW identification result (which follows from consistency, unconfoundedness, and overlap),
$\mathbb{E}\!\left[\frac{D_{i}}{\pi(v)}Y_{i}\,\Big|\,V_{i}=v\right]=\mathbb{E}[Y_{i}(1)\mid V_{i}=v],\qquad\mathbb{E}\!\left[\frac{1-D_{i}}{1-\pi(v)}Y_{i}\,\Big|\,V_{i}=v\right]=\mathbb{E}[Y_{i}(0)\mid V_{i}=v].$
Therefore,
$\mathbb{E}[\Lambda(V_{i})\mid V_{i}=v]=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid V_{i}=v].$
Combining the two cases proves double robustness.
Identification based on Outcome Modeling (On the Treated)
Remark 3
Consider a binary treatment $D_{i}\in\{0,1\}$ and covariates $V_{i}=(X_{i},Z_{i})$. Assume (i) consistency/SUTVA: $Y_{i}=D_{i}Y_{i}(1)+(1-D_{i})Y_{i}(0)$, (ii) unconfoundedness: $(Y_{i}(0),Y_{i}(1))\perp D_{i}\mid V_{i}$, and (iii) overlap in $X$: $0<\Pr(D_{i}=1\mid X_{i})<1$ a.s. Define $\mu_{0}(V_{i}):=\mathbb{E}[Y_{i}\mid D_{i}=0,V_{i}]$. Then
$\mathbb{E}\!\Biggl[\frac{(Y_{i}-\mu_{0}(V_{i}))\,D_{i}}{\Pr(D_{i}=1\mid X_{i})}\,\Bigg|\,X_{i}\Biggr]=\mathbb{E}\!\Bigl[Y_{i}(1)-Y_{i}(0)\mid D_{i}=1,X_{i}\Bigr].$
Proof. Fix $x$. Since $\Pr(D_{i}=1\mid X_{i}=x)$ depends only on $x$,
$\mathbb{E}\!\Biggl[\frac{(Y_{i}-\mu_{0}(V_{i}))\,D_{i}}{\Pr(D_{i}=1\mid X_{i})}\,\Bigg|\,X_{i}=x\Biggr]=\frac{1}{\Pr(D_{i}=1\mid X_{i}=x)}\,\mathbb{E}\!\Bigl[(Y_{i}-\mu_{0}(V_{i}))\,D_{i}\mid X_{i}=x\Bigr].$
Because $D_{i}$ is binary,
$\mathbb{E}\!\Bigl[(Y_{i}-\mu_{0}(V_{i}))\,D_{i}\mid X_{i}=x\Bigr]=\mathbb{E}\!\bigl[Y_{i}-\mu_{0}(V_{i})\mid D_{i}=1,X_{i}=x\bigr]\Pr(D_{i}=1\mid X_{i}=x).$
By consistency, conditional on $D_{i}=1$ we have $Y_{i}=Y_{i}(1)$, hence
$\mathbb{E}\!\bigl[Y_{i}-\mu_{0}(V_{i})\mid D_{i}=1,X_{i}=x\bigr]=\mathbb{E}\!\bigl[Y_{i}(1)-\mu_{0}(V_{i})\mid D_{i}=1,X_{i}=x\bigr].$
Next, for any $v$,
$\mu_{0}(v)=\mathbb{E}[Y_{i}\mid D_{i}=0,V_{i}=v]=\mathbb{E}[Y_{i}(0)\mid D_{i}=0,V_{i}=v]=\mathbb{E}[Y_{i}(0)\mid V_{i}=v]=\mathbb{E}[Y_{i}(0)\mid D_{i}=1,V_{i}=v],$
where the second equality uses consistency (when $D_{i}=0$) and the third uses unconfoundedness. Therefore,
$\mathbb{E}[Y_{i}(1)-\mu_{0}(V_{i})\mid D_{i}=1,V_{i}]=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid D_{i}=1,V_{i}].$
Taking conditional expectations given $D_{i}=1,X_{i}=x$ and using iterated expectations,
$\mathbb{E}[Y_{i}(1)-\mu_{0}(V_{i})\mid D_{i}=1,X_{i}=x]=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid D_{i}=1,X_{i}=x].$
Substituting back,
$\mathbb{E}\!\Biggl[\frac{(Y_{i}-\mu_{0}(V_{i}))\,D_{i}}{\Pr(D_{i}=1\mid X_{i})}\,\Bigg|\,X_{i}=x\Biggr]=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid D_{i}=1,X_{i}=x],$
as claimed.
Identification based on IPW (On the Treated)
Remark 4
Consider a binary treatment $D_{i}\in\{0,1\}$ and covariates $V_{i}=(X_{i},Z_{i})$. Assume (i) consistency/SUTVA: $Y_{i}=D_{i}Y_{i}(1)+(1-D_{i})Y_{i}(0)$, (ii) unconfoundedness: $(Y_{i}(0),Y_{i}(1))\perp D_{i}\mid V_{i}$, and (iii) overlap: $0<\pi(V_{i})<1$ a.s., and $0<\Pr(D_{i}=1\mid X_{i})<1$ a.s., where $\pi(V_{i}):=\Pr(D_{i}=1\mid V_{i})$. Then
$\mathbb{E}\!\Biggl[\frac{Y_{i}\bigl(D_{i}-\pi(V_{i})\bigr)}{\Pr(D_{i}=1\mid X_{i})\bigl(1-\pi(V_{i})\bigr)}\,\Bigg|\,X_{i}\Biggr]=\mathbb{E}\!\Bigl[Y_{i}(1)-Y_{i}(0)\mid D_{i}=1,X_{i}\Bigr].$
Proof. Define $\tau(v):=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid V_{i}=v]$. By unconfoundedness, $\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid D_{i}=1,V_{i}=v]=\tau(v)$. Hence for any $x$,
$\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid D_{i}=1,X_{i}=x]=\mathbb{E}[\tau(V_{i})\mid D_{i}=1,X_{i}=x].$
By iterated expectations and Bayes’ rule,
$\mathbb{E}[\tau(V_{i})\mid D_{i}=1,X_{i}=x]$ $=\int\tau(x,z)\,f(z\mid D_{i}=1,X_{i}=x)\,dz$
$=\int\tau(x,z)\,\frac{f(z\mid X_{i}=x)\pi(x,z)}{\Pr(D_{i}=1\mid X_{i}=x)}\,dz$
$=\frac{1}{\Pr(D_{i}=1\mid X_{i}=x)}\,\mathbb{E}\!\left[\tau(V_{i})\,\pi(V_{i})\ \big|\ X_{i}=x\right].$
Next, by the IPW identity (from consistency, unconfoundedness, and overlap),
$\tau(V_{i})=\mathbb{E}\!\Biggl[\frac{Y_{i}D_{i}}{\pi(V_{i})}-\frac{Y_{i}(1-D_{i})}{1-\pi(V_{i})}\ \Bigg|\ V_{i}\Biggr].$
Therefore,
$\mathbb{E}\!\left[\tau(V_{i})\pi(V_{i})\mid X_{i}=x\right]$ $=\mathbb{E}\!\left[\mathbb{E}\!\left[\left(\frac{Y_{i}D_{i}}{\pi(V_{i})}-\frac{Y_{i}(1-D_{i})}{1-\pi(V_{i})}\right)\pi(V_{i})\ \Bigg|\ V_{i}\right]\Bigg|\ X_{i}=x\right]$
$=\mathbb{E}\!\left[Y_{i}D_{i}-\frac{Y_{i}(1-D_{i})\pi(V_{i})}{1-\pi(V_{i})}\ \Bigg|\ X_{i}=x\right].$
Combining the two steps,
$\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid D_{i}=1,X_{i}=x]=\mathbb{E}\!\left[\frac{Y_{i}D_{i}}{\Pr(D_{i}=1\mid X_{i}=x)}-\frac{Y_{i}(1-D_{i})\pi(V_{i})}{\Pr(D_{i}=1\mid X_{i}=x)(1-\pi(V_{i}))}\ \Bigg|\ X_{i}=x\right].$
Finally, note
$Y_{i}D_{i}-\frac{Y_{i}(1-D_{i})\pi(V_{i})}{1-\pi(V_{i})}=\frac{Y_{i}\bigl(D_{i}-\pi(V_{i})\bigr)}{1-\pi(V_{i})},$
so
$\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid D_{i}=1,X_{i}=x]=\mathbb{E}\!\left[\frac{Y_{i}\bigl(D_{i}-\pi(V_{i})\bigr)}{\Pr(D_{i}=1\mid X_{i}=x)(1-\pi(V_{i}))}\ \Bigg|\ X_{i}=x\right],$
which is the desired result.
Identification based on AIPW (On the Treated)
Remark 5
Assume (i) consistency/SUTVA: $Y_{i}=D_{i}Y_{i}(1)+(1-D_{i})Y_{i}(0)$, (ii) unconfoundedness: $(Y_{i}(0),Y_{i}(1))\perp D_{i}\mid V_{i}$, and (iii) overlap: $0<\pi(V_{i})<1$ a.s. and $0<\Pr(D_{i}=1\mid X_{i})<1$ a.s., where $\pi(V_{i}):=\Pr(D_{i}=1\mid V_{i})$. Define the signal
$\Xi(V_{i}):=\frac{1}{\Pr(D_{i}=1\mid X_{i})}\Bigl(Y_{i}-\mu_{0}(V_{i})\Bigr)\left(D_{i}-\frac{\pi(V_{i})(1-D_{i})}{1-\pi(V_{i})}\right).$
If either $\pi(V_{i})$ is the true propensity score or $\mu_{0}(V_{i})=\mathbb{E}[Y_{i}\mid D_{i}=0,V_{i}]$ is the true control regression, then for any $x$,
$\mathbb{E}\bigl[\Xi(V_{i})\mid X_{i}=x\bigr]=\mathbb{E}\bigl[Y_{i}(1)-Y_{i}(0)\mid D_{i}=1,X_{i}=x\bigr].$
Proof. Fix $x$ and let $p(x):=\Pr(D_{i}=1\mid X_{i}=x)$. Let $\tau(v):=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid V_{i}=v]$. By unconfoundedness, $\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid D_{i}=1,V_{i}=v]=\tau(v)$. Thus:
$\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid D_{i}=1,X_{i}=x]=\mathbb{E}[\tau(V_{i})\mid D_{i}=1,X_{i}=x].$
Using Bayes’ rule (or the definition of conditional expectation), we can rewrite the target as:
$\mathbb{E}[\tau(V_{i})\mid D_{i}=1,X_{i}=x]=\frac{1}{p(x)}\,\mathbb{E}\bigl[\tau(V_{i})\Pr(D_{i}=1|V_{i})\mid X_{i}=x\bigr].$
It therefore suffices to show that the numerator of $\Xi(V_{i})$ satisfies:
$\mathbb{E}\left[\Bigl(Y_{i}-\mu_{0}(V_{i})\Bigr)\left(D_{i}-\frac{\pi(V_{i})(1-D_{i})}{1-\pi(V_{i})}\right)\ \Bigg|\ X_{i}=x\right]=\mathbb{E}\bigl[\tau(V_{i})\Pr(D_{i}=1|V_{i})\mid X_{i}=x\bigr].$
Case 1: The propensity score is correct. Assume $\pi(V_{i})=\Pr(D_{i}=1\mid V_{i})$. Conditioning on $V_{i}$, the expectation of the term inside the square brackets is:
$\mathbb{E}\left[\Bigl(Y_{i}-\mu_{0}(V_{i})\Bigr)D_{i}\mid V_{i}\right]-\frac{\pi(V_{i})}{1-\pi(V_{i})}\mathbb{E}\left[\Bigl(Y_{i}-\mu_{0}(V_{i})\Bigr)(1-D_{i})\mid V_{i}\right].$
Note that $\mu_{0}(V_{i})$ is arbitrary here (potentially misspecified). Consider the second term. By consistency, if $D_{i}=0$, then $Y_{i}=Y_{i}(0)$. Thus:
$\mathbb{E}[(Y_{i}-\mu_{0}(V_{i}))(1-D_{i})\mid V_{i}]=(1-\pi(V_{i}))\bigl(\mathbb{E}[Y_{i}(0)\mid V_{i}]-\mu_{0}(V_{i})\bigr).$
Consider the first term. If $D_{i}=1$, $Y_{i}=Y_{i}(1)$. Thus:
$\mathbb{E}[(Y_{i}-\mu_{0}(V_{i}))D_{i}\mid V_{i}]=\pi(V_{i})\bigl(\mathbb{E}[Y_{i}(1)\mid V_{i}]-\mu_{0}(V_{i})\bigr).$
Subtracting the second term (scaled by $\frac{\pi}{1-\pi}$) from the first:
$\pi(V_{i})\bigl(\mathbb{E}[Y_{i}(1)|V_{i}]-\mu_{0}(V_{i})\bigr)-\frac{\pi(V_{i})}{1-\pi(V_{i})}(1-\pi(V_{i}))\bigl(\mathbb{E}[Y_{i}(0)|V_{i}]-\mu_{0}(V_{i})\bigr)$
$=\pi(V_{i})\Bigl(\mathbb{E}[Y_{i}(1)|V_{i}]-\mu_{0}(V_{i})-\mathbb{E}[Y_{i}(0)|V_{i}]+\mu_{0}(V_{i})\Bigr)$
$=\pi(V_{i})\tau(V_{i}).$
The misspecified $\mu_{0}$ cancels out perfectly. Taking the expectation over $V_{i}|X_{i}=x$ matches the target.
Case 2: The outcome regression is correct. Assume $\mu_{0}(V_{i})=\mathbb{E}[Y_{i}\mid D_{i}=0,V_{i}]$, but allow $\pi(V_{i})$ to be misspecified. The term of interest is:
$(Y_{i}-\mu_{0}(V_{i}))D_{i}\;-\;\frac{\pi(V_{i})}{1-\pi(V_{i})}(Y_{i}-\mu_{0}(V_{i}))(1-D_{i}).$
Conditional on $V_{i}$, the expected value of the second part is:
$\frac{\pi(V_{i})}{1-\pi(V_{i})}\Pr(D_{i}=0\mid V_{i})\,\mathbb{E}[Y_{i}-\mu_{0}(V_{i})\mid V_{i},D_{i}=0].$
Since $\mu_{0}$ is the true conditional mean for controls, $\mathbb{E}[Y_{i}-\mu_{0}(V_{i})\mid V_{i},D_{i}=0]=0$. Thus, the entire second term vanishes, regardless of the value of $\pi(V_{i})$. The remaining term is:
$\mathbb{E}[(Y_{i}-\mu_{0}(V_{i}))D_{i}\mid V_{i}].$
Using consistency and unconfoundedness:
$\mathbb{E}[(Y_{i}(1)-\mu_{0}(V_{i}))\cdot 1\mid V_{i},D_{i}=1]\Pr(D_{i}=1|V_{i}).$
Since $\mu_{0}(V_{i})=\mathbb{E}[Y_{i}(0)|V_{i}]$, this becomes:
$\Pr(D_{i}=1\mid V_{i})\bigl(\mathbb{E}[Y_{i}(1)|V_{i}]-\mathbb{E}[Y_{i}(0)|V_{i}]\bigr)=\Pr(D_{i}=1|V_{i})\tau(V_{i}).$
This matches the target moment exactly. Thus, the estimator is consistent if either nuisance parameter is correctly specified.
A.3 Chapter 4 Proofs
Signals and CME (Continuous Treatment)
To prove: $\mathbb{E}[\Lambda(\eta)\mid V=v]=\mathbb{E}[\partial_{d}Y(d)\mid V=v]$.
Proof
By the tower rule, $\mathbb{E}[\Lambda(\eta)\mid X=x]=\mathbb{E}\Bigl[\mathbb{E}[\Lambda(\eta)\mid V]\,\Big|\;X=x\Bigr],$ so it suffices to show that $\mathbb{E}[\Lambda(\eta)\mid V=v]=\mathbb{E}[\partial_{d}Y(d)\mid V=v].$ We want to show:
$\mathbb{E}[\Lambda(\eta)\mid V=v]=\mathbb{E}[\partial_{d}Y(D)\mid V=v].$
Expanding on the Neyman orthogonal score:
$\mathbb{E}\bigl[\Lambda(\eta)\mid V=v\bigr]$ $=\mathbb{E}\Bigl[\partial_{d}\mu(D,v)\mid V=v\Bigr]-\mathbb{E}\Bigl[\partial_{d}\log f(D|v)\,\bigl[Y-\mu(D,v)\bigr]\,\Big|\;V=v\Bigr]$
Note that we are integrating over the joint distribution of $Y$ and $D$ given $V=v$. We verify the second term is zero.
$\mathbb{E}\Bigl[-\,\partial_{d}\log f(D|v)\,\bigl(Y-\mu(D,v)\bigr)\;\Big|\;V=v\Bigr]=$
$\int_{d}\!\!\int_{y}\Bigl(-\,\partial_{d}\log f(d|v)\,\bigl[y-\mu(d,v)\bigr]\Bigr)\;f_{Y\mid D,V}(y\mid d,v)\,f_{D\mid V}(d\mid v)\;dy\,dd.$
By definition of the conditional mean, $\int y\,f_{Y\mid D,V}(y\mid d,v)\,dy=\mu(d,v).$ It follows that $\int[y-\mu(d,v)]\,f_{Y\mid D,V}(y\mid d,v)\,dy=0.$ Hence:
$\mathbb{E}\Bigl[-\,\partial_{d}\log f(D|v)\,\bigl(Y-\mu(D,v)\bigr)\;\Big|\;V=v\Bigr]=0.$
Therefore,
$\mathbb{E}[\Lambda(\eta)\mid V=v]=\mathbb{E}[\partial_{d}\mu(D,v)\mid V=v]=\mathbb{E}[\partial_{d}Y(D)\mid V=v].$
Score Definition
To verify the Neyman orthogonality for the AIPW score, we show that the Gâteaux derivative of the expected score with respect to the nuisance functions is zero at the true nuisance parameters.
For each unit $i$, define the AIPW score:
$\Lambda_{i}(\eta)=\mu_{1}(V_{i})-\mu_{0}(V_{i})+\frac{D_{i}}{\pi(V_{i})}\bigl[Y_{i}-\mu_{1}(V_{i})\bigr]-\frac{1-D_{i}}{1-\pi(V_{i})}\bigl[Y_{i}-\mu_{0}(V_{i})\bigr].$
where $\eta$ collectively denotes the nuisance functions $\eta=(\mu_{1},\mu_{0},\pi)$. At the true nuisance parameters $\eta_{0}=(\mu_{1}^{0},\mu_{0}^{0},\pi^{0})$, this score identifies the pointwise CATE.
To show Neyman orthogonality, we must verify:
$\left.\frac{\partial}{\partial\eta}\mathbb{E}\bigl[\Lambda_{i}(\eta)\bigr]\right|_{\eta=\eta_{0}}=0.$
We will look at the partial derivatives of $\mathbb{E}[\Lambda_{i}(\eta)]$ with respect to each nuisance function $\mu_{1},\mu_{0},\text{and }\pi$. In each case, we show the derivative is zero when evaluated at $\eta_{0}$.
Derivative with Respect to $\mu_{1}$
Consider a small directional perturbation $h_{\mu_{1}}(v)$ of $\mu_{1}$. The Gâteaux derivative in that direction is:
$\left.\frac{d}{dt}\right|_{t=0}\mathbb{E}\Bigl[\Lambda_{i}\bigl(\mu_{1}+t\,h_{\mu_{1}},\,\mu_{0},\pi\bigr)\Bigr].$
Within $\Lambda_{i}$, $\mu_{1}$ appears in two places: $\mu_{1}(V_{i})$ and inside the term $\frac{D_{i}}{\pi(V_{i})}[Y_{i}-\mu_{1}(V_{i})]$. Taking the derivative yields:
$\text{Partial w.r.t.\ }\mu_{1}:\quad h_{\mu_{1}}(V_{i})\;-\;\frac{D_{i}}{\pi(V_{i})}\bigl[h_{\mu_{1}}(V_{i})\bigr]=h_{\mu_{1}}(V_{i})\Bigl(1-\frac{D_{i}}{\pi(V_{i})}\Bigr).$
Since $h_{\mu_{1}}(V_{i})$ is a function of $V_{i}$:
$\mathbb{E}\Bigl[h_{\mu_{1}}(V_{i})\Bigl(1-\frac{D_{i}}{\pi(V_{i})}\Bigr)\Bigr]=\mathbb{E}\Bigl[h_{\mu_{1}}(V_{i})\,\mathbb{E}\bigl[1-\frac{D_{i}}{\pi(V_{i})}\bigm|\,V_{i}\bigr]\Bigr].$
At the true $\eta_{0}$, $\pi^{0}(v)=\Pr(D_{i}=1\mid V_{i}=v)$. Therefore:
$\mathbb{E}\Bigl[\,1-\frac{D_{i}}{\pi^{0}(V_{i})}\Bigm|\,V_{i}=v\Bigr]=1-\frac{\mathbb{E}[D_{i}\mid V_{i}=v]}{\,\pi^{0}(v)\,}=1-\frac{\pi^{0}(v)}{\,\pi^{0}(v)\,}=0.$
By an identical argument, the derivative w.r.t. $\mu_{0}$ also vanishes.
Derivative with Respect to $\pi$
Now consider a small perturbation $h_{\pi}(v)$ of $\pi$. Within $\Lambda_{i}$, $\pi$ appears in the denominators:
$\frac{D_{i}}{\,\pi(V_{i})\,}\quad\text{and}\quad\frac{1-D_{i}}{\,1-\pi(V_{i})\,}.$
[…]
$h_{\pi}(V_{i})\,\left(-\,\frac{D_{i}}{\,\pi(V_{i})^{2}\,}\bigl[Y_{i}-\mu_{1}(V_{i})\bigr]\;+\;\frac{1-D_{i}}{\,[1-\pi(V_{i})]^{2}}\bigl[Y_{i}-\mu_{0}(V_{i})\bigr]\right).$
Factor out $h_{\pi}(V_{i})$ and condition on $V_{i}=v$:
$\mathbb{E}\Bigl[h_{\pi}(V_{i})\,\Bigl(-\,\frac{D_{i}}{\,\pi(V_{i})^{2}\,}(Y_{i}-\mu_{1}(V_{i}))\;+\;\frac{1-D_{i}}{\,[1-\pi(V_{i})]^{2}}\,(Y_{i}-\mu_{0}(V_{i}))\Bigr)\Bigr]=$
$\mathbb{E}\Bigl[h_{\pi}(V_{i})\,\mathbb{E}\Bigl[-\,\frac{D_{i}}{\,\pi^{0}(v)^{2}\,}(Y_{i}-\mu_{1}^{0}(v))\;+\;\frac{1-D_{i}}{\,[1-\pi^{0}(v)]^{2}\,}(Y_{i}-\mu_{0}^{0}(v))\Bigm|\,V_{i}=v\Bigr]\Bigr].$
Define $\Gamma(v)$ as the sum of two expectations:
$\Gamma(v)=\underbrace{\mathbb{E}\Bigl[-\frac{D_{i}}{\pi^{0}(v)^{2}}(Y_{i}-\mu_{1}^{0}(v))\Bigm|V_{i}=v\Bigr]}_{T_{1}(v)}\quad+\quad\underbrace{\mathbb{E}\Bigl[\frac{1-D_{i}}{[1-\pi^{0}(v)]^{2}}(Y_{i}-\mu_{0}^{0}(v))\Bigm|V_{i}=v\Bigr]}_{T_{2}(v)}.$
We will show each of $T_{1}(v)$ and $T_{2}(v)$ is zero under the true nuisance parameters.
$T_{1}(v)$ Vanishes
$T_{1}(v)=-\frac{1}{\pi^{0}(v)^{2}}\mathbb{E}\Bigl[D_{i}(Y_{i}-\mu_{1}^{0}(v))\Bigm|V_{i}=v\Bigr].$
Expand the inner expectation:
$\mathbb{E}[D_{i}(Y_{i}-\mu_{1}^{0}(v))\mid V_{i}=v]=\mathbb{E}[D_{i}Y_{i}\mid V_{i}=v]-\mu_{1}^{0}(v)\mathbb{E}[D_{i}\mid V_{i}=v].$
Evaluate at the true parameters:
$\mathbb{E}[D_{i}Y_{i}\mid V_{i}=v]$ $=\Pr(D_{i}=1\mid V_{i}=v)\cdot\mathbb{E}[Y_{i}\mid D_{i}=1,V_{i}=v]=\pi^{0}(v)\mu_{1}^{0}(v),$
$\mathbb{E}[D_{i}\mid V_{i}=v]$ $=\pi^{0}(v).$
Thus:
$\mathbb{E}[D_{i}Y_{i}\mid V_{i}=v]$ $=\pi^{0}(v)\mu_{1}^{0}(v)$
$\mathbb{E}[D_{i}(Y_{i}-\mu_{1}^{0}(v))\mid V_{i}=v]$ $=\pi^{0}(v)\mu_{1}^{0}(v)-\mu_{1}^{0}(v)\pi^{0}(v)=0.$
Therefore $T_{1}(v)=-\frac{1}{\pi^{0}(v)^{2}}\cdot 0=0$.
$T_{2}(v)$ Vanishes
$T_{2}(v)=\frac{1}{[1-\pi^{0}(v)]^{2}}\mathbb{E}\Bigl[(1-D_{i})(Y_{i}-\mu_{0}^{0}(v))\Bigm|V_{i}=v\Bigr].$
Similarly, expand the inner expectation:
$\mathbb{E}[(1-D_{i})(Y_{i}-\mu_{0}^{0}(v))\mid V_{i}=v]=\mathbb{E}[(1-D_{i})Y_{i}\mid V_{i}=v]-\mu_{0}^{0}(v)\mathbb{E}[(1-D_{i})\mid V_{i}=v].$
Evaluate at the true parameter:
$\mathbb{E}[(1-D_{i})Y_{i}\mid V_{i}=v]$ $=(1-\pi^{0}(v))\mu_{0}^{0}(v)$
$\mathbb{E}[(1-D_{i})\mid V_{i}=v]$ $=1-\pi^{0}(v).$
Thus,
$\mathbb{E}[(1-D_{i})(Y_{i}-\mu_{0}^{0}(v))\mid V_{i}=v]=(1-\pi^{0}(v))\mu_{0}^{0}(v)-\mu_{0}^{0}(v)(1-\pi^{0}(v))=0.$
Thus, $T_{2}(v)=\frac{1}{[1-\pi^{0}(v)]^{2}}\cdot 0=0$.
Putting the two parts together: $\Gamma(v)=T_{1}(v)+T_{2}(v)=0+0=0$ which means for every $v$:
$\mathbb{E}\Bigl[-\frac{D_{i}}{\pi^{0}(v)^{2}}(Y_{i}-\mu_{1}^{0}(V_{i}))+\frac{1-D_{i}}{[1-\pi^{0}(v)]^{2}}(Y_{i}-\mu_{0}^{0}(v))\Bigm|V_{i}=v\Bigr]=0.$
Consequently,
$\mathbb{E}\Bigl[h_{\pi}(V_{i})\Bigl(-\frac{D_{i}}{\pi^{0}(V_{i})^{2}}(Y_{i}-\mu_{1}^{0}(V_{i}))+\frac{1-D_{i}}{[1-\pi^{0}(V_{i})]^{2}}(Y_{i}-\mu_{0}^{0}(V_{i}))\Bigr)\Bigr]=0,$
for any directional perturbation $h_{\pi}(\cdot)$. This completes the proof that the derivative of $\mathbb{E}[\Lambda_{i}(\eta)]$ w.r.t. $\pi$ is zero at $\eta_{0}$.
By combining this argument with the analogous derivatives w.r.t. $\mu_{1}$ and $\mu_{0}$, we conclude that
$\left.\frac{\partial}{\partial\eta}\,\mathbb{E}[\Lambda_{i}(\eta)]\right|_{\eta=\eta_{0}}=0$
for all directions of perturbation. Hence, the AIPW score is Neyman-orthogonal to first-order errors in $\eta$, which underlies its double robustness property and the extension to DML settings.